### v2

# Statistical Inference for Pairwise Comparison Models

Abstract
Pairwise comparison models have been widely used for utility evaluation and ranking across various fields. The increasing scale of problems today underscores the need to understand statistical inference in these models when the number of subjects diverges, a topic currently lacking in the literature except in a few special instances. To partially address this gap, this paper establishes a near-optimal asymptotic normality result for the maximum likelihood estimator in a broad class of pairwise comparison models, as well as a non-asymptotic convergence rate for each individual subject under comparison. The key idea lies in identifying the Fisher information matrix as a weighted graph Laplacian, which can be studied via a meticulous spectral analysis. Our findings provide a unified theory for performing statistical inference in a wide range of pairwise comparison models beyond the Bradley–Terry model, benefiting practitioners with theoretical guarantees for their use. Simulations utilizing synthetic data are conducted to validate the asymptotic normality result, followed by a hypothesis test using a tennis competition dataset.
Keywords: Imbalanced data, Maximum likelihood estimation, Ordinal data, Pairwise comparison models, Sparsity
1 Introduction
Pairwise comparison involves assessing subjects in pairs to establish their relative preferences, a practice pertinent to numerous applications such as sports analytics (Glickman and Jones, 1999; Ley et al., 2019; Tsokos et al., 2019), econometrics (McFadden, 1973), video coding (Lee et al., 2011), and social science (Loewen et al., 2012). A prevalent approach for pairwise comparison modeling employs a latent score framework. Originated from the ideas in Thurstone (Thurstone, 1927) and Zermelo (Zermelo, 1929), a mathematical model for pairwise comparison data analysis was proposed by Bradley and Terry (Bradley and Terry, 1952). Since then, multiple generalizations have been developed, including ordinal models such as the Rao–Kupper model (Rao and Kupper, 1967) and the Davidson model (Davidson, 1970), which account for ties, the cumulative link model (Agresti, 1992) that considers more refined ordinal scales, and cardinal models such as the paired cardinal model (Shah et al., 2016). We recommend Cattelan (2012) for a review of pairwise comparison modeling from a practical perspective.
In light of the fast-growing scale of problems in the big-data era, a recent research area of pairwise comparison models focuses on understanding the asymptotics of estimating the latent score vector as the number of compared subjects approaches infinity. Along this line of work, considerable attention has been devoted to the study of the Bradley–Terry (BT) model, which is popular due to its simple parametrization form and connections to other problems such as matrix balancing (Qu et al., 2023) and completion (Chen et al., 2023). Notably, an efficient Minorize-Maximization (MM) algorithm for implementing the maximum likelihood estimator (MLE) was proposed in Hunter (2004) and accelerated in Newman (2023) in the BT model, and its uniform consistency was established when the comparison graph is dense (Simons and Yao, 1999; Yan et al., 2012). Subsequent works have established analogous consistency results for alternative estimators such as the spectral estimator and regularized MLE in the sparse settings under various error metrics (Chen and Suh, 2015; Shah et al., 2016; Negahban et al., 2017; Chen et al., 2019). The first result that generalized the uniform consistency of the MLE to the sparse setting appeared in Han et al. (2020), paralleled by a few other works with a slightly different focus (e.g. asymptotic efficiency and minimax rates); see Hendrickx et al. (2020); Bong and Rinaldo (2022); Chen et al. (2022). It is noteworthy that Han et al. (2020) also contains an asymptotic normality result for dense graphs, which was further explored in the sparse setting (Liu et al., 2023; Gao et al., 2023).
Despite extensive research on the asymptotic properties of the BT model, a significant gap persists between theory and practice, as many comparison models used in applications either have multiple scales or a continuous outcome. This suggests a practical need to extend this understanding to other widely used yet less explored models. The recent work Han et al. (2023) partially addresses this by introducing a general pairwise comparison model framework and proving a uniform consistency result for the MLE therein. However, an asymptotic normality result has not yet been established. Nonetheless, such a result is often crucial for practical statistical inference and cannot be deduced from the existing approaches to analyzing the BT model.
To take one step further in filling the gap, this paper aims to derive an asymptotic normality result for the MLE in the general pairwise comparison models. In particular, our contributions can be summarized as follows:
  • We establish an asymptotic normality result for the MLE in the general pairwise comparison models under a near-optimal sparsity condition on the comparison graphs. Additionally, we derive a non-asymptotic convergence rate for each individual subject under comparison. The crux of our analysis involves recognizing the Fisher information matrix as a weighted graph Laplacian, enabling a meticulous analysis based on spectral expansion. Aside from enjoying the sufficient flexibility of model parametrization, our results additionally allow the incorporation of imbalanced data, which is new compared to many existing results in the field and may be of independent interest; see Table 1 for a detailed comparison between our result and some existing results in the literature.
  • We verify that the MLE in many pairwise comparison models in the literature, including the BT model, the Thurstone–Mosteller model (Thurstone, 1927; Mosteller, 1951), the Davidson model, the Rao–Kupper model, the cumulative link model, and the paired cardinal model, is asymptotically normal under near-optimal sparsity conditions. We also provide explicit and computable formulas for their asymptotic variance that may be of interest to practitioners using these models.
Besides the theoretical findings, we conduct numerical experiments to illustrate our results through concrete examples, including building confidence intervals in a simulation study and hypothesis testing on an Association of Tennis Professionals (ATP) dataset.
Results Models Comparison Graphs Sparsity
Simons and Yao (1999) BT homogeneous $\Omega(1)$
Yan et al. (2012) BT heterogeneous $\Omega(1)$
Han et al. (2020) BT homogeneous $\widetilde{\Omega}(n^{-1/10})$
Gao et al. (2023) BT homogeneous $\widetilde{\Omega}(n^{-1})$
This work general heterogeneous $\widetilde{\Omega}(n^{-1})$
Table 1: Comparison of several asymptotic normality results in the pairwise comparison model literature. Homogeneous comparison graphs indicate the balanced data while heterogeneous graphs correspond to the imbalanced data. The notation $\Omega(\cdot)$ is the standard Bachmann–Landau notation and $\widetilde{\Omega}(\cdot)$ means that the asymptotic relation holds up to polynomial terms of $\log n$.
The rest of the paper is organized as follows. In Section 2, we introduce the mathematical framework for general pairwise comparison models. In Section 3, we establish the main results concerning asymptotic normality and non-asymptotic convergence rates (Theorems 8, 11, 12) for the MLE and provide the relevant interpretations. In Section 4, we show that the MLE in many pairwise comparison models in the literature is asymptotically normal under minimal sparsity assumptions. Furthermore, we offer explicit formulas for their asymptotic variance. In Section 5, we validate the theoretical findings using both synthetic and real data. In Section 6, we outline the key ideas for the proof of the main result, deferring the technical details to the appendix section. In Section 7, we point out some future directions.
2 General pairwise comparison models
We introduce the general pairwise comparison models proposed in Han et al. (2023). Consider an undirected graph ${\mathcal{G}}=({\mathcal{V}},{\mathcal{E}})$ with $n$ vertices, where ${\mathcal{V}}=[n]:=\{1,\ldots,n\}$ represents the subjects being modeled, and ${\mathcal{E}}$ is the edge set of the comparison data, that is, for any $i,j\in{\mathcal{V}}$, $(i,j)\in{\mathcal{E}}$ if there exists a comparison between subjects $i$ and $j$. For ease of illustration, we assume that ${\mathcal{G}}$ is simple so that there exists at most one edge between two subjects; the general case consisting of multiple edges can be considered similarly but requires heavier notation. We use $\delta\{i\}=\{j\in{\mathcal{V}}:(i,j)\in{\mathcal{E}}\}$ to denote the graph neighbourhood of $i$, which refers to the set of vertices adjacent to $i$. For the moment, we assume that ${\mathcal{G}}$ is known and describe the statistical model for the comparison outcomes.
The comparison outcomes are modeled using latent scores. Let ${\bm{u}}^{*}=(u^{*}_{1},\ldots,u^{*}_{n})\in{\mathbb{R}}^{n}$ denote the latent score vector of the subjects under comparison. For instance, in a sports competition, one may think of $u^{*}_{i}$ as the internal strength of the $i$th player when competing with others. For $i,j\in[n]$, assuming there is a comparison between them, the outcome is a random variable $X_{ij}$ taking values in some symmetric subset ${\mathbb{A}}\subseteq{\mathbb{R}}$. The probability mass/density function of $X_{ij}$ follows a single-parameter family $f(x;u^{*}_{i}-u^{*}_{j})$, depending on whether $X_{ij}$ is discrete or continuous. In the general pairwise comparison models, the function $f$ is assumed to be valid in the sense of the following definition.
Definition 1 (Valid parameterization)
A function $f:{\mathbb{A}}\times\mathbb{R}\to\mathbb{R}^{+}$, where ${\mathbb{A}}=-{\mathbb{A}}\subseteq{\mathbb{R}}$ denotes the possible comparison outcomes, is said to be valid if it satisfies the following assumptions:
(A1). (normalization) For $y\in{\mathbb{R}}$, $\int_{\mathbb{A}}f(x;y)\operatorname{d}\!{x}=1$, where the integral is interpreted as summation if ${\mathbb{A}}$ is discrete.
(A2). (symmetry) $f$ is even with respect to $(x;y)$, that is, $f(x;y)=f(-x;-y)$, for $(x,y)\in{\mathbb{A}}\times{\mathbb{R}}.$
(A3). (monotonicity) For $x<0$, $f(x;y)$ is decreasing in $y$, and $f(x;y)\to 0$ as $y\to\infty$.
(A4). (boundedness) $\sup_{y\in{\mathbb{R}}}f(x;y)<+\infty$ for every $x\in{\mathbb{A}}$.
(A5). (log-concavity) $f(x;y)$ is strictly log-concave with respect to $y$.
All pairwise comparison models mentioned in the introduction are special cases of general pairwise comparison models with suitable choices of valid $f$. In the rest of the article, we shall always assume $f$ is valid.
Having specified $f$, we can write down the likelihood function given the observed data and compute the MLE $\widehat{{\bm{u}}}$ for ${\bm{u}}^{*}$ under identifiability constraints; see Section 3. It was shown in Han et al. (2023) that $\widehat{{\bm{u}}}$ is a uniformly consistent estimator for ${\bm{u}}^{*}$ under appropriate conditions on both $f$ and ${\mathcal{G}}$. To identify graph configurations satisfying these conditions, it is often convenient to assume that ${\mathcal{G}}$ is sampled from certain random graph ensembles. A commonly used one is the random graph model ${\mathsf{G}}\left(n,p_{n},q_{n}\right)$ that generalizes the Erdős–Rényi model (Erdős and Rényi, 1960).
Definition 2 (Random graph models)
${\mathsf{G}}\left(n,p_{n},q_{n}\right)$ is a random graph with vertices set ${\mathcal{V}}=[n]$ where each edge $(i,j)\in{\mathcal{V}}\times{\mathcal{V}},i\neq j$ is formed independently with deterministic probability $p_{ij,n}\in\left[p_{n},q_{n}\right]$.
Remark 3
By definition, for fixed $p_{n}$ and $q_{n}$, different specifications of $\{p_{ij,n}\}_{1\leq i<j\leq n}$ may belong to the same ${\mathsf{G}}\left(n,p_{n},q_{n}\right)$. Such ambiguity is irrelevant to our discussion as only the upper and lower bounds are used in the analysis.
In the following section, we assume that the comparison graph ${\mathcal{G}}({\mathcal{V}},{\mathcal{E}})$ is sampled from ${\mathsf{G}}\left(n,p_{n},q_{n}\right)$. Meanwhile, for every $(i,j)\in{\mathcal{E}}$, we assume that a single comparison outcome $X_{ij}\sim f(x;u^{*}_{i}-u^{*}_{j})$ is observed. Note that $X_{ij}=-X_{ji}$, and we assume that $\{X_{ij}\}_{i<j,(i,j)\in{\mathcal{E}}}$ are independent. It is worth noting that we allow $\lim\sup_{n\to\infty}q_{n}/p_{n}=+\infty$, which leads to heterogeneity of the observed data. In particular, one subject may have $nq_{n}$ comparisons while another has only $np_{n}$ comparisons. Such an imbalance of data volume will bring additional difficulty in theoretical data analysis.
3 Main results
This section establishes the asymptotic normality of the MLE $\widehat{{\bm{u}}}$ in the general pairwise comparison model introduced in Section 2. In the following, we use $O(\cdot)$, $O_{p}(\cdot)$, $o(\cdot)$, and $o_{p}(\cdot)$ as a standard notation to represent the respective asymptotic order between two sequences.
The log-likelihood function based on the observed outcomes $X_{ij}$ (conditional on ${\mathcal{G}}({\mathcal{V}},{\mathcal{E}})$) is
$l({\bm{u}})=\sum_{(i,j)\in{\mathcal{E}}}\log f(X_{ij};u_{i}-u_{j}).$
Note that $l({\bm{u}})$ is invariant if all components of ${\bm{u}}$ are translated by a constant. To ensure model identifiability, we let ${\bm{1}}_{n}^{\top}{\bm{u}}=0$, where ${\bm{1}}_{n}$ is the all-ones vector with the same dimension as ${\bm{u}}$. The MLE ${\widehat{\bm{u}}}$ for ${\bm{u}}^{*}$ satisfies
${\widehat{\bm{u}}}=\operatorname*{arg\,max}_{{\bm{u}}\in{\mathbb{R}}^{n}:{\bm{1}}_{n}^{\top}{\bm{u}}=0}l({\bm{u}}).$ (1)
To state the asymptotic normality result for $\widehat{{\bm{u}}}$, we need the following assumptions. Let $M_{n}=\max_{i,j\in[n]}|u^{*}_{i}-u^{*}_{j}|$ denote the dynamic range of ${\bm{u}}^{*}$. The global discrepancy of the model, which measures the probability of a subject with the highest score winning over another with the lowest score, is defined as
$c_{n,1}=\int_{{\mathbb{A}}\cap[0,+\infty)}f(x;M_{n})\operatorname{d}\!{x}.$
It can be verified using Definition 1 that $c_{n,1}\in[1/2,1)$.
Let $g(x;y)=\partial_{2}(\log f(x;y))=\partial_{2}f(x;y)/f(x;y)$ be the Fisher score function, where $\partial_{2}$ denotes the partial derivative operator for the second argument $y$. The first assumption, which ensures the unique existence of the MLE, states that $c_{n,1}$ if converging to one, must have a controlled rate.
Assumption 4
The global discrepancy $c_{n,1}$ satisfies
$\frac{\log n}{np_{n}|\log c_{n,1}|}\to 0\qquad n\to\infty.$ (2)
The next assumption is concerned with the tail decay of $g(X;y)$ with $X\sim f(x;y)$ for all $|y|\leq M_{n}$, where $g(X;y)$ is centered, that is, ${\mathbb{E}}[g(X;y)]=0$.
Assumption 5
The sequence of random variables $\{g(X_{y};y)\}_{|y|\leq M_{n}}$ is uniformly subgaussian, where $X_{y}\sim f(x;y)$. That is,
$c_{n,2}=\max_{|y|\leq M_{n}}\|g(X_{y};y)\|_{\psi_{2}}<\infty,$
where $\|\cdot\|_{\psi_{2}}$ is the subgaussian norm (Vershynin, 2018, Definition 2.5.6).
In addition, we further requires a boundedness condition on the partial derivative of $g(x;y)$ in $y$.
Assumption 6
The partial derivative of the score function $g$ for $y$ is uniformly bounded from both above and below:
$0<c_{n,3}=\inf_{x\in{\mathbb{A}},|y|\leq M_{n}+1}|\partial_{2}g(x;y)|\leq\sup_{x\in{\mathbb{A}},|y|\leq M_{n}+1}|\partial_{2}g(x;y)|=c_{n,4}<\infty.$
Assumptions 4-6 are sufficient to guarantee the uniform consistency of the maximum likelihood estimator (Han et al., 2023). The common pathway from consistency to asymptotic normality is via Taylor’s expansion. As a result, one often require more stringent regularity conditions on higher-order derivatives.
Assumption 7
The second-order partial derivative of the score function $g(x,y)$ for $y$ is uniformly bounded for all $|y|\leq M_{n}+1$:
$c_{n,5}=\sup_{x\in{\mathbb{A}},|y|\leq M_{n}+1}|\partial_{22}g(x;y)|<\infty.$
Theorem 8 (Asymptotic normality)
Let ${\mathcal{G}}\sim{\mathsf{G}}(n,p_{n},q_{n})$. Under Assumptions 47, if
$\beta_{n}=\max\left\{\frac{c_{n,2}^{2}c_{n,4}^{5/2}c_{n,5}}{c_{n,3}^{5}},\frac{c_{n,2}c_{n,4}^{11/2}}{c_{n,3}^{6}}\right\}\left\{\frac{q_{n}^{10}(\log n)^{8}}{np_{n}^{11}}\right\}^{1/2}\to 0\qquad n\to\infty,$ (3)
then for each $i\in[n]$, $\{\rho_{i}({\bm{u}}^{*})\}^{-1/2}(\widehat{u}_{i}-u_{i}^{*})\to N(0,1)$, where $N(0,1)$ is the standard normal distribution and the asymptotic variance $\rho_{i}({\bm{u}}^{*})$ is given by
$\rho_{i}({\bm{u}}^{*})$ $=\left[\sum_{j\in\delta\{i\}}\int_{{\mathbb{A}}}\frac{\{\partial_{2}f(x;u^{*}_{i}-u^{*}_{j})\}^{2}}{f(x;u^{*}_{i}-u^{*}_{j})}\operatorname{d}\!{x}\right]^{-1}.$ (4)
Moreover, for any finite $S=\{i_{1},\ldots,i_{s}\}\subset{\mathbb{N}}$ where $S$ is independent of $n$, $\{\rho_{i_{1}}({\bm{u}}^{*})\}^{-1/2}(\widehat{u}_{i_{1}}-u_{i_{1}}^{*}),\ldots,\{\rho_{i_{s}}({\bm{u}}^{*})\}^{-1/2}(\widehat{u}_{i_{s}}-u_{i_{s}}^{*})$ are asymptotically independent.
Remark 9
Subject to additional conditions, such as a uniform upper bound on the number of edges between two subjects that is independent of ${\bm{u}}^{*}$, Theorem 8 still holds when two subjects are compared multiple times. Under such circumstances, the asymptotic variance formula (4) remains valid if one counts multiple edges for each adjacent vertex.
A proof sketch for Theorem 8 is provided in Section 6. To better understand Theorem 8, note that although the multiplicative factor involving $c_{n,i}$ may seem daunting, it is $O(1)$ under additional regularity and boundedness conditions.
Corollary 10
Assume that $M^{*}=\sup_{n}M_{n}<\infty$ and ${\mathbb{A}}$ is finite. If $f(x;y)$ is a vaild function, and $g(x;y),\partial_{2}g(x;y),\partial_{22}g(x;y)$ are continuous in $y$, then the asymptotic normality result in Theorem 8 holds if $q^{10}_{n}(\log n)^{8}/(np^{11}_{n})\to 0$ as $n\to\infty$.
Corollary 10 can be further simplified under balanced data conditions. If $\sup_{n}q_{n}/p_{n}<\infty$, then the condition in Corollary 10 reduces to $(\log n)^{8}/(np_{n})\to 0$ as $n\to\infty$. This bound coincides with the well-known Erdős–Rényi connectivity threshold up to logarithmic factors and is thus close to optimal. In general, Corollary 10 allows $p_{n}$ and $q_{n}$ to have different asymptotic orders. For instance, if $p_{n}=n^{-b}$ for $10\leq b<$ 1, then choosing $q_{n}=n^{-b^{\prime}}$ for any $b^{\prime}\geq 0$ with $b-(1-b)/10<b^{\prime}\leq b$ satisfies the required conditions. To the best of our knowledge, this is the first result that works for general comparison models with imbalanced data.
Theorem 8 provides the limit distribution of the MLE, which lays the foundation for further statistical inference, such as the construction of interval-based estimates, testing hypothesis on the difference of merits, and testing the hypothesis on ranking position (Fan et al., 2024). Additionally, we can derive a non-asymptotic error rate for each parameter as well as a practical construction of interval-valued estimators, which are summarized in Theorem 11 and Theorem 12, respectively.
Theorem 11 (Individual estimation error)
Let ${\mathcal{G}}\sim{\mathsf{G}}(n,p_{n},q_{n})$. If both Assumptions 47 and (3) hold, then there exists an absolute constant $C>0$ such that for each $i\in[n]$, the following holds for all sufficiently large $n$ with probability tending to one:
$|\widehat{u}_{i}-u_{i}^{*}|\leq C\cdot\frac{c_{n,2}}{c_{n,3}}\left(\frac{\log n}{|\delta\{i\}|}\right)^{1/2}.$ (5)
We call the result in (5) the individual estimation error bound. Compared with the existing works (Han et al., 2020; Chen et al., 2022), this result presents new findings on the individual estimation error. Specifically, previous works considered the overall estimation error $\max_{i\in[n]}|\widehat{u}_{i}-u_{i}^{*}|$, while our result in (5) further illustrates the error for each individual. As expected, this error depends on the degree $|\delta\{i\}|$, which serves as a local normalization factor for each individual. When the comparison graph is homogeneous, individual estimation errors converge at the same rate, matching the results in Han et al. (2020); Chen et al. (2022). For heterogeneous graphs, however, individuals will have different convergence rates, which cannot be reflected in the worst-case error $\max_{i\in[n]}|\widehat{u}_{i}-u_{i}^{*}|$.
Next, we construct the confidence interval estimators. Since the asymptotic variance $\rho_{i}({\bm{u}}^{*})$ is unknown, it is natural to consider the following plug-in variance estimator:
$\rho_{i}({\widehat{\bm{u}}})$ $=\left[\sum_{j\in\delta\{i\}}\int_{{\mathbb{A}}}\frac{\{\partial_{2}f(x;\widehat{u}_{i}-\widehat{u}_{i})\}^{2}}{f(x;\widehat{u}_{i}-\widehat{u}_{i})}\operatorname{d}\!{x}\right]^{-1}.$ (6)
Then for any fixed $i\in[n]$ and $\alpha\in(0,1)$, the $(1-\alpha)$-confidence interval of $u_{i}^{*}$ is
$\mathcal{CI}_{i}=\big{(}\widehat{u}_{i}-z_{\alpha/2}\{\rho_{i}({\widehat{\bm{u}}})\}^{-1/2},\widehat{u}_{i}+z_{\alpha/2}\{\rho_{i}({\widehat{\bm{u}}})\}^{-1/2}\big{)},$
where $z_{\alpha/2}$ is the $(1-\alpha/2)$-quantile of standard normal distribution. Note that confidence intervals constructed using consistently estimated plug-in estimators are often automatically valid due to Slutsky’s theorem if the parameters to be estimated are independent of $n$. In our case, the increasing dimension of the parameter vector makes such a result less obvious, though it remains true, as justified by the following theorem.
Theorem 12 (Confidence interval)
Let ${\mathcal{G}}\sim{\mathsf{G}}(n,p_{n},q_{n})$. Under Assumptions 47, if (3) holds, then for any fixed $i\in[n]$ and $\alpha\in(0,1)$,
$\mathbb{P}\left(u^{*}_{i}\in\mathcal{CI}_{i}\right)\to 1-\alpha\ \text{ as }n\to\infty.$
The computation of $\mathcal{CI}_{i}$ is straightforward after obtaining ${\widehat{\bm{u}}}$. For example, if one wants to construct a $95\%$-confidence interval for the $i$th subject, one can take $\mathcal{CI}_{i}=\big{(}\widehat{u}_{i}-1.96\{\rho_{i}({\widehat{\bm{u}}})\}^{-1/2},\widehat{u}_{i}+1.96\{\rho_{i}({\widehat{\bm{u}}})\}^{-1/2}\big{)}$.
4 Examples
We establish the asymptotic normality of the MLEs for a variety of pairwise comparison models found in the existing literature under conditions characterized by $p_{n},q_{n}$, and $M_{n}$. These models encompass the BT model, the Thurstone–Mosteller model, the Rao–Kupper model, the Davidson model, the cumulative link model (with four outcomes), and the paired cardinal model. In particular, explicit formulas are derived for their asymptotic variance. For conveniences, we let $\Delta_{ij}=u^{*}_{i}-u^{*}_{j}$ for $i,j\in[n]$.
Example 1 (BT model)
The BT model assumes a binary comparison outcome between two subjects and parameterizes using a logistic link function:
$f(1;y)=\frac{e^{y}}{1+e^{y}};\quad f(-1;y)=\frac{1}{1+e^{y}}.$
It can be verified that both conditions in (2) and (3) are satisfied if
$e^{6M_{n}}\left\{\frac{q_{n}^{10}(\log n)^{8}}{np_{n}^{11}}\right\}^{1/2}\to 0\qquad n\to\infty.$
The asymptotic variance $\rho_{i}({\bm{u}}^{*})$ for the subject $i\in[n]$ is given by
$\rho_{i}({\bm{u}}^{*})=\left\{\sum_{j\in\delta\{i\}}\frac{e^{\Delta_{ij}}}{\left(1+e^{\Delta_{ij}}\right)^{2}}\right\}^{-1}.$
That is, $\{{\rho_{i}({\bm{u}}^{*})}\}^{-1/2}(\widehat{u}_{i}-u^{*}_{i})\to N(0,1)$ as $n\to\infty$.
Example 2 (Thurstone–Mosteller model)
The Thurstone–Mosteller model considers an alternative link function than the BT model:
$f(1;y)=\Phi(y);\quad f(-1;y)=1-\Phi(y),$
where $\Phi(y)=\int_{-\infty}^{y}\varphi(z)\operatorname{d}\!{z}$ with $\varphi(z)=(2\pi)^{-1/2}e^{-z^{2}/2}$. It can be verified that both conditions in (2) and (3) are satisfied if
$M_{n}^{6}e^{3M^{2}_{n}}\left\{\frac{q_{n}^{10}(\log n)^{8}}{np_{n}^{11}}\right\}^{1/2}\to 0\qquad n\to\infty.$
The asymptotic variance $\rho_{i}({\bm{u}}^{*})$ for the subject $i\in[n]$ is given by
$\rho_{i}({\bm{u}}^{*})=\left\{\sum_{j\in\delta\{i\}}\frac{\varphi^{2}(\Delta_{ij})}{\Phi(\Delta_{ij})(1-\Phi(\Delta_{ij}))}\right\}^{-1}.$
That is, $\{{\rho_{i}({\bm{u}}^{*})}\}^{-1/2}(\widehat{u}_{i}-u^{*}_{i})\to N(0,1)$ as $n\to\infty$.
Example 3 (Rao–Kupper model)
The Rao–Kupper model extends the BT model by incorporating ties. The link function is given by
$f(1;y)$ $=\frac{e^{y}}{e^{y}+\theta};\quad f(0;y)=\frac{\left(\theta^{2}-1\right)e^{y}}{\left(e^{y}+\theta\right)\left(\theta e^{y}+1\right)};\quad f(-1;y)=\frac{1}{\theta e^{y}+1},$
where $\theta>1$ is the threshold parameter which is predetermined. It can be verified that both conditions in (2) and (3) are satisfied if
$e^{6M_{n}}\left\{\frac{q_{n}^{10}(\log n)^{8}}{np_{n}^{11}}\right\}^{1/2}\to 0\qquad n\to\infty.$
The asymptotic variance $\rho_{i}({\bm{u}}^{*})$ for the subject $i\in[n]$ can be computed as
$\rho_{i}({\bm{u}}^{*})=\left\{\sum_{j\in\delta\{i\}}\left[\frac{\theta^{2}e^{\Delta_{ij}}}{(\theta+e^{\Delta_{ij}})^{3}}+\frac{\theta^{2}(\theta^{2}-1)e^{\Delta_{ij}}(1-e^{2\Delta_{ij}})^{2}}{(e^{\Delta_{ij}}+\theta)^{3}(\theta e^{\Delta_{ij}}+1)^{3}}+\frac{\theta^{2}e^{2\Delta_{ij}}}{(\theta e^{\Delta_{ij}}+1)^{3}}\right]\right\}^{-1}.$
That is, $\{{\rho_{i}({\bm{u}}^{*})}\}^{-1/2}(\widehat{u}_{i}-u^{*}_{i})\to N(0,1)$ as $n\to\infty$.
Example 4 (Davidson model)
As opposed to the Rao–Kupper model, the Davidson model considers an alternative parameterization of being tied:
$f(1;y)=\frac{e^{y}}{e^{y}+\theta e^{\frac{y}{2}}+1};\quad f(0;y)=\frac{\theta e^{\frac{y}{2}}}{e^{y}+\theta e^{\frac{y}{2}}+1};\quad f(-1;y)$ $=\frac{1}{e^{y}+\theta e^{\frac{y}{2}}+1},$
where $\theta>0$ is assumed to be prefixed. It can be verified that both conditions in (2) and (3) are satisfied if
$e^{{3M_{n}}}\left\{\frac{q_{n}^{10}(\log n)^{8}}{np_{n}^{11}}\right\}^{1/2}\to 0\qquad n\to\infty.$
The asymptotic variance $\rho_{i}({\bm{u}}^{*})$ for the subject $i\in[n]$ is given by
$\rho_{i}({\bm{u}}^{*})=\left\{\sum_{j\in\delta\{i\}}\frac{e^{\Delta_{ij}}(\theta e^{\Delta_{ij}/2}+2)^{2}+\theta e^{\Delta_{ij}/2}(1-e^{\Delta_{ij}})^{2}+(2e^{\Delta_{ij}}+\theta e^{\Delta_{ij}/2})^{2}}{4(e^{\Delta_{ij}}+\theta e^{\Delta_{ij}/2}+1)^{3}}\right\}^{-1}.$
That is, $\{{\rho_{i}({\bm{u}}^{*})}\}^{-1/2}(\widehat{u}_{i}-u^{*}_{i})\to N(0,1)$ as $n\to\infty$.
Example 5 (Cumulative link model with four outcomes)
The cumulative link model is defined via the ordinalization of some distribution function $F$ (Agresti, 1992). The BT model, the Rao–Kupper model, and the Thurstone–Mosteller model can all be considered as specific cases of the cumulative link model with appropriate choices of $F$. In this example, we focus on the situation when $F$ is logistic with four outcomes, which will be used later in Section 5. In this case, $\mathbb{A}=\{-2,-1,1,2\}$ and the corresponding link function takes the form of
$f(1;y)=\frac{(\theta-1)e^{y}}{\left(\theta+e^{y}\right)\left(1+e^{y}\right)},\ \ \ f(2;y)=\frac{e^{y}}{\theta+e^{y}},$ (7)
where $\theta>1$ is the threshold parameter that is predetermined. It can be verified that both conditions in (2) and (3) are satisfied if
$e^{6M_{n}}\left\{\frac{q_{n}^{10}(\log n)^{8}}{np_{n}^{11}}\right\}^{1/2}\to 0\qquad n\to\infty.$
The asymptotic variance $\rho_{i}({\bm{u}}^{*})$ for the subject $i\in[n]$ is given by
${\rho_{i}({\bm{u}}^{*})=\left\{\sum_{j\in\delta\{i\}}\left[\frac{\theta^{2}e^{\Delta_{ij}}}{(\theta+e^{\Delta_{ij}})^{3}}+\frac{\theta^{2}e^{2\Delta_{ij}}}{(\theta e^{\Delta_{ij}}+1)^{3}}+\frac{(\theta-1)e^{\Delta_{ij}}}{(1+e^{\Delta_{ij}})^{3}}\left(\frac{(\theta-e^{\Delta_{ij}})^{2}}{(\theta+e^{\Delta_{ij}})^{3}}+\frac{1-\theta e^{2\Delta_{ij}}}{(\theta e^{\Delta_{ij}}+1)^{3}}\right)\right]\right\}^{-1}.}$
That is, $\{{\rho_{i}({\bm{u}}^{*})}\}^{-1/2}(\widehat{u}_{i}-u^{*}_{i})\to N(0,1)$ as $n\to\infty$. The result above can be extended to the cumulative link model with the logistic link function and finite outcomes but involve more complicated formulas for asymptotic variances.
Example 6 (Paired cardinal model)
The paired cardinal model can be viewed as a fully observed version of the Thurstone–Mosteller model, where the comparison outcome takes values in $\mathbb{A}={\mathbb{R}}$, and the corresponding link function is parameterized using the normal density:
$f(x;y)=\left(2\pi\sigma^{2}\right)^{-\frac{1}{2}}e^{-\frac{(x-y)^{2}}{2\sigma^{2}}},$
where $\sigma>0$ is some fixed parameter. It can be verified that both conditions in (2) and (3) are satisfied if
$\max\left\{M_{n}e^{\frac{M_{n}^{2}}{2\sigma^{2}}}\frac{\log n}{np_{n}},\left({\frac{q_{n}^{10}(\log n)^{8}}{np_{n}^{11}}}\right)^{1/2}\right\}\to 0\qquad n\to\infty.$
The asymptotic variance $\rho_{i}({\bm{u}}^{*})$ for the subject $i\in[n]$ is given by
$\rho_{i}({\bm{u}}^{*})=\frac{\sigma^{2}}{|\delta\{i\}|},$
which is independent of ${\bm{u}}^{*}$. That is, $\{{\rho_{i}({\bm{u}}^{*})}\}^{-1/2}(\widehat{u}_{i}-u^{*}_{i})\to N(0,1)$ as $n\to\infty$.
5 Numerical studies
5.1 Synthetic data
We conduct a simulation study to verify the asymptotic normality of the MLE. We focus on three comparison models, the Davidson model, the Rao–Kupper model, and the paired cardinal model, as discussed in the preceding section.
We utilize the general random graph model introduced in Definition 2 to generate comparison graphs, where we set $p_{n}=n^{-1/2}$ and $q_{n}=p_{n}(\log n)$, with the individual edge probabilities chosen uniformly at random between them. The total number of subjects $n$ in all the tested models is chosen from the set $\{500,1000,2000\}$. To generate the utility vector ${\bm{u}}^{*}$, we uniformly select values from the range $[-M_{n},M_{n}]$, where the dynamic range $M_{n}$ is chosen from $\{1,\log(\log n)\}$. For global parameters in each model, we set $\theta=1$ in the Davidson model, $\theta=2$ in the Rao–Kupper model, and $\sigma=2$ in the paired cardinal model. Each simulation scenario is characterized by the pair $(n,M_{n})$ and repeated $300$ times. We report the average standard deviation of the estimated parameters and the coverage probability of the $95\%$ confidence interval, as shown in Table 2. We also examine the $z$-scores of the first coordinate of ${{\widehat{\bm{u}}}}$ over 300 experiments and plot the corresponding quantile-quantile plots against normal quantiles in Figure 1.
In Table 2, we observe a decreasing trend in the average standard deviation as $n$ increases, which is consistent with our theoretical findings on asymptotic variance in Equation (4). Additionally, we note variations in standard deviation relative to $M_{n}$ across different models. In particular, the standard deviations are increasing in $M_{n}$ in the Rao–Kupper and Davidson models, while remaining unchanged in the paired cardinal model. This observation aligns with their respective asymptotic variance formulas $\rho_{i}({\bm{u}}^{*})$ computed in Section 4; in particular, $\rho_{i}({\bm{u}}^{*})$ in the paired cardinal model is independent of $M_{n}$.
The empirical coverage probabilities of the constructed confidence intervals for all models under comparison closely approximate the desired coverage of $95\%$. This, combined with the diagonal alignment of the data in Figure 1, suggests that the expected asymptotic normality holds.
Table 2: Summarized simulation results over 300 replications.
$M_{n}=1$ $M_{n}=\log(\log n)$
$n$ 500 1000 2000 500 1000 2000
Rao–Kupper Standard deviation $0.157$ $0.126$ $0.101$ $0.172$ $0.139$ $0.113$
Coverage probability $0.949$ $0.949$ $0.949$ $0.948$ $0.949$ $0.949$
Davidson Standard deviation $0.209$ $0.167$ $0.134$ $0.225$ $0.181$ $0.147$
Coverage probability $0.949$ $0.950$ $0.950$ $0.949$ $0.951$ $0.950$
Paired cardinal Standard deviation $0.164$ $0.131$ $0.105$ $0.163$ $0.131$ $0.105$
Coverage probability $0.950$ $0.950$ $0.950$ $0.950$ $0.950$ $0.949$
Figure 1: Quantile-quantile plots comparing the theoretical and sample quantiles of $\widehat{u}_{1}$ in the Rao–Kupper model, the Davidson model, and the paired cardinal model with fixed $M_{n}=1$. Simulation parameters: For each $n$, the empirical quantile curves are based on 300 experiments.
5.2 Real data analysis
[…]
$f(X_{ij}=1,u_{i}-u_{j})=\frac{(\theta-1)e^{u_{i}-u_{j}}}{\left(\theta+e^{u_{i}-u_{j}}\right)\left(1+e^{u_{i}-u_{j}}\right)},$
$f(X_{ij}=2,u_{i}-u_{j})=\frac{e^{u_{i}-u_{j}}}{\theta+e^{u_{i}-u_{j}}},\ \ \ \theta>1.$
To identify whether a significant distinction exists between the Big Three in tennis and Andy Murray, we employed a multiple testing procedure on the difference in their latent scores. Specifically, let $u_{1},u_{2},u_{3}$ and $u_{4}$ denote the latent score of Djokovic, Nadal, Federer, and Murray respectively. We consider the hypothesis testing with the null hypothesis $H_{0}:u_{i}=u_{4},i\in[3]$ and alternative hypothesis $H_{A}:$ there exists $i\in[3]$ such that $u_{i}\neq u_{4}$. The significance level $\alpha=0.05.$ We first obtained the maximum likelihood estimates of the latent scores $\widehat{{\bm{u}}}$, and the threshold parameter $\widehat{\theta}=2.32$. The estimator $\widehat{u}_{i}$ and estimated standard deviation $(\widehat{\rho}_{i})^{1/2}$ of the 4 candidates are presented in Table 3. We construct the $z$-statistics $(\widehat{u}_{i}-\widehat{u}_{4})/{{(\widehat{\rho}_{i}+\widehat{\rho}_{4})^{1/2}}},\ i\in[3]$, which yields $p$-values $0.229$, $0.226$ and $0.334$ accordingly. For multiple testing, we adopt the Benjamini–Hochberg procedure (Benjamini and Hochberg, 1995). We order three $p$-values in ascending order and compare them with the corresponding critical values, that is $(\alpha)/3$, $(2\alpha)/3$, and $\alpha$. It turns out that all $p$-values exceed the corresponding critical value, implying that the difference in latent scores between Big Three and Murray is not statistically significant. Consequently, there is no sufficient statistical evidence under the selected threshold to reject the null hypothesis which supports the Big Four.
Table 3: The estimated scores and their standard deviation of four players by using CLM4.
Players Estimated scores $\widehat{{\bm{u}}}_{i}$ Estimated standard deviation $(\widehat{\rho}_{i})^{1/2}$
Novak Djokovic 3.235 $0.229$
Rafael Nadal 3.214 $0.203$
Roger Federer 3.129 $0.179$
Andy Murray 2.872 $0.196$
6 Proof of Theorem 8
In this section, we provide a sketch of the proof for Theorem 8. Following a common route for establishing asymptotic normality, we apply Taylor’s expansion to the log-likelihood function and try to find the inverse of the corresponding Hessian matrix. However, the randomness of the Hessian matrix introduces additional challenges. The novelty of our approach lies in identifying the Hessian matrix as the weighted Laplacian matrix of appropriate random graphs, which can be carefully analyzed using truncated spectral methods.
We start by computing the first- and second-order derivatives of the log-likelihood function $l$ for ${\bm{u}}$. For $i,j\in[n]$ with $i\neq j$,
$\partial_{i}l({\bm{u}})$ $=\sum_{j\in\delta\{i\}}\frac{\partial_{2}f(X_{ij},u_{i}-u_{j})}{f(X_{ij},u_{i}-u_{j})}=\sum_{j\in\delta\{i\}}g(X_{ij},u_{i}-u_{j}),$
$\partial_{ij}l({\bm{u}})$ $=\frac{-\partial_{22}f(X_{ij},u_{i}-u_{j})\cdot f(X_{ij},u_{i}-u_{j})+\{\partial_{2}f(X_{ij},u_{i}-u_{j})\}^{2}}{f^{2}(X_{ij},u_{i}-u_{j})}\cdot\mathbb{I}_{\{(i,j)\in{\mathcal{E}}\}}$
$=-\partial_{2}g(X_{ij},u_{i}-u_{j})\cdot\mathbb{I}_{\{(i,j)\in{\mathcal{E}}\}}\geq 0$
$\partial_{ii}l({\bm{u}})$ $=-\sum_{j\in\delta\{i\}}\partial_{ij}l({\bm{u}})\leq 0,$
where the first inequality follows from the log-concavity of $f$ that is assumed under the validity assumption. Then the Hession of $l({\bm{u}})$, denoted by $H({\bm{u}})$ is given as
$\{H({\bm{u}})\}_{ij}$ =$=\left\{\begin{array}[]{ll}\partial_{ij}l({\bm{u}})&\quad i\neq j\\ -\sum_{k\in\delta\{i\}}\partial_{ik}l({\bm{u}})&\quad i=j.\end{array}\right.$
A crucial observation is that $H({\bm{u}})$ is the negation of an unnormalized graph Laplacian associated with a weighted graph on $[n]$ with edge weights $\partial_{ij}l({\bm{u}})$. Conditional on ${\mathcal{E}}$, $H({\bm{u}})$ is a random quantity due to the comparison outcomes $X_{ij}$. As such, we let $H^{*}({\bm{u}})={\mathbb{E}}[H({\bm{u}})\mid{\mathcal{E}}]$. It is easy to check that $-H^{*}({\bm{u}})$ remains an unnormalized weighted graph Laplacian that depends only on ${\mathcal{G}}({\mathcal{V}},{\mathcal{E}})$ and ${\bm{u}}$. In the following, we normalize $-H^{*}({\bm{u}}^{*})$ as
$-H^{*}({\bm{u}}^{*})={\mathcal{D}}^{1/2}{\mathcal{L}}_{sym}{\mathcal{D}}^{1/2}={\mathcal{D}}^{1/2}(I-{\mathcal{A}}){\mathcal{D}}^{1/2},$ (8)
where ${\mathcal{D}}=\text{diag}\{-H^{*}({\bm{u}}^{*})\}$ is the diagonal of $-H^{*}({\bm{u}}^{*})$ (degree matrix), ${\mathcal{L}}_{sym}$ is the normalized Laplacian of the weighted graph associated with $-H^{*}({\bm{u}}^{*})$, ${\mathcal{A}}=I-{\mathcal{L}}_{sym}$ is the normalized adjacency matrix, and $I$ is the identity matrix. By the mean value theorem,
$\nabla l({\bm{u}}^{*})$ $=\nabla l({\bm{u}}^{*})-\nabla l({\widehat{\bm{u}}})=-\bar{H}({\bm{W}})({\widehat{\bm{u}}}-{\bm{u}}^{*})$
$=-H^{*}({\bm{u}}^{*})({\widehat{\bm{u}}}-{\bm{u}}^{*})+\{H^{*}({\bm{u}}^{*})-H({\bm{u}}^{*})\}({\widehat{\bm{u}}}-{\bm{u}}^{*})+\{H({\bm{u}}^{*})-\bar{H}({\bm{W}})\}({\widehat{\bm{u}}}-{\bm{u}}^{*}),$
where
$\bar{H}({\bm{W}})=\begin{pmatrix}\partial_{11}l({\bm{w}}_{1}),\ldots,\partial_{1n}l({\bm{w}}_{1})\\ \\ \partial_{n1}l({\bm{w}}_{n}),\ldots,\partial_{nn}l({\bm{w}}_{n})\end{pmatrix}\in{\mathbb{R}}^{n\times n},$
with ${\bm{W}}=({\bm{w}}_{1},\ldots,{\bm{w}}_{n}),$ and ${\bm{w}}_{1},\ldots,{\bm{w}}_{n}$ are $n$ (possibly distinct) points on the line segment between ${\widehat{\bm{u}}}$ and ${\bm{u}}^{*}$. It is worth noting that $\bar{H}({\bm{W}})$ is not a Hessian matrix (it is not even symmetric). Heuristically, thanks to the matrix concentration and uniform consistency of ${\widehat{\bm{u}}}$, $H^{*}({\bm{u}}^{*})-H({\bm{u}}^{*})=o(H({\bm{u}}^{*}))$ and $H({\bm{u}}^{*})-\bar{H}({\bm{W}})=o(H({\bm{u}}^{*}))$, where the $o(\cdot)$ is interpreted in an appropriate sense that will be made clear in Lemma 13.
With some rearrangement and substituting (8) into $H^{*}({\bm{u}}^{*})$, we have
${\mathcal{D}}^{1/2}({\widehat{\bm{u}}}-{\bm{u}}^{*})$ $=-{\mathcal{L}}^{\dagger}_{sym}{\mathcal{D}}^{-1/2}\nabla l({\bm{u}}^{*})$ (9)
$\ \ \ \ \ +{\mathcal{L}}^{\dagger}_{sym}{\mathcal{D}}^{-1/2}\{H^{*}({\bm{u}}^{*})-H({\bm{u}}^{*})\}({\widehat{\bm{u}}}-{\bm{u}}^{*})$
$\ \ \ \ \ +{\mathcal{L}}^{\dagger}_{sym}{\mathcal{D}}^{-1/2}\{H({\bm{u}}^{*})-\bar{H}({\bm{W}})\}({\widehat{\bm{u}}}-{\bm{u}}^{*}),$
where ${\mathcal{L}}_{sym}^{\dagger}$ is the pseudoinverse of ${\mathcal{L}}_{sym}$ and can be explicitly expressed using spectral expansion:
${\mathcal{L}}_{sym}^{\dagger}=\sum_{t=0}^{\infty}({\mathcal{A}}-\mathcal{P}_{1})^{t}-\mathcal{P}_{1},$ (10)
where $\mathcal{P}_{1}$ is the orthogonal projector to the null space of ${\mathcal{L}}$. Particularly, letting $H_{ij}=\{H^{*}({\bm{u}}^{*})\}_{ij}$, the entry of ${\mathcal{A}}$ and $\mathcal{P}_{1}$ can be written as
${\mathcal{D}}_{ii}=\sum_{j\in\delta\{i\}}H_{ij},\quad{\mathcal{A}}_{ij}=\frac{H_{ij}}{{{\mathcal{D}}_{ii}^{1/2}{\mathcal{D}}_{jj}^{1/2}}},\quad\{\mathcal{P}_{1}\}_{ij}=\frac{{{\mathcal{D}}_{ii}^{1/2}{\mathcal{D}}_{jj}^{1/2}}}{\sum_{k\in[n]}{\mathcal{D}}_{kk}}.$ (11)
Based on the decomposition in (C.2), we will proceed to prove the desired asymptotic normality result as follows. We will show that the first term on the right-hand side of (C.2) appropriately normalizes $({\widehat{\bm{u}}}-{\bm{u}}^{*})$ and gives the asymptotic normality. Meanwhile, through a meticulous analysis (Lemma 13), it can be established that the remaining terms are of order $o_{p}(1)$. Consequently, the desired follows by applying Slutsky’s theorem.
Lemma 13
Under the assumptions in Theorem 8, the following are true:
  1. (1)
    $\|{\mathcal{L}}^{\dagger}_{sym}{\mathcal{D}}^{-1/2}\{H^{*}({\bm{u}}^{*})-H({\bm{u}}^{*})\}({\widehat{\bm{u}}}-{\bm{u}}^{*})\|_{\infty}=o_{p}(1)$.
  2. (2)
    $\|{\mathcal{L}}^{\dagger}_{sym}{\mathcal{D}}^{-1/2}\{H({\bm{u}}^{*})-\bar{H}({\bm{W}})\}({\widehat{\bm{u}}}-{\bm{u}}^{*})\|_{\infty}=o_{p}(1)$.
  3. (3)
    $\max_{i\in[n]}\sum_{t=2}^{\infty}\{({\mathcal{A}}-\mathcal{P}_{1})^{t}\}_{ii}=o_{p}(1)$.
The proof of Lemma 13 is rather technical and deferred to the appendix. We now use the results in Lemma 13 to establish the asymptotic normality of ${\widehat{\bm{u}}}$. By statements (1) and (2) in Lemma 13 and Slutsky’s theorem, it remains to verify for every fixed $S\subset{\mathbb{N}}$ with $|S|=s$ ($S$ independent of $n$), $\{{\mathcal{L}}^{\dagger}_{sym}{\mathcal{D}}^{-1/2}\nabla l({\bm{u}}^{*})\}_{S}\to N(0,I_{s})$. Since
${\mathcal{L}}^{\dagger}_{sym}{\mathcal{D}}^{-1/2}\nabla l({\bm{u}}^{*})={\mathcal{D}}^{-1/2}\nabla l({\bm{u}}^{*})+({\mathcal{L}}^{\dagger}_{sym}-I){\mathcal{D}}^{-1/2}\nabla l({\bm{u}}^{*})$ (12)
and $\{{\mathcal{D}}^{-1/2}\nabla l({\bm{u}}^{*})\}_{S}\to N(0,I_{s})$ by the central limit theorem, the proof is finished if $\{({\mathcal{L}}^{\dagger}_{sym}-I){\mathcal{D}}^{-1/2}\nabla l({\bm{u}}^{*})\}_{S}=o_{p}(1)$ componentwise. To verify this, we apply Chebyshev’s inequality. Note ${\mathbb{E}}[({\mathcal{L}}^{\dagger}_{sym}-I){\mathcal{D}}^{-1/2}\nabla l({\bm{u}}^{*})]=0$ and the second moment calculation yields that for each $k\in[n]$ and the $k$th canonical basis vector $e_{k}\in{\mathbb{R}}^{n}$,
${\mathbb{E}}\left[e_{k}^{\top}({\mathcal{L}}^{\dagger}_{sym}-I){\mathcal{D}}^{-1/2}\nabla l({\bm{u}}^{*})\{\nabla l({\bm{u}}^{*})\}^{\top}{\mathcal{D}}^{-1/2}({\mathcal{L}}^{\dagger}_{sym}-I)e_{k}\right]$
$=$ $\ e_{k}^{\top}({\mathcal{L}}^{\dagger}_{sym}-I){\mathcal{D}}^{-1/2}{\mathbb{E}}[\nabla l({\bm{u}}^{*})\{\nabla l({\bm{u}}^{*})\}^{\top}]{\mathcal{D}}^{-1/2}({\mathcal{L}}^{\dagger}_{sym}-I)e_{k}$
$=$ $\ -e_{k}^{\top}({\mathcal{L}}^{\dagger}_{sym}-I){\mathcal{D}}^{-1/2}H^{*}({\bm{u}}^{*}){\mathcal{D}}^{-1/2}({\mathcal{L}}^{\dagger}_{sym}-I)e_{k}$
$=$ $\ e_{k}^{\top}({\mathcal{L}}^{\dagger}_{sym}-I){\mathcal{L}}_{sym}({\mathcal{L}}^{\dagger}_{sym}-I)e_{k}$
$=$ $\ ({\mathcal{L}}^{\dagger}_{sym}+\mathcal{P}_{1}-I)_{kk}+({\mathcal{L}}_{sym}+\mathcal{P}_{1}-I)_{kk}$
${{=}}$ $\ \sum_{t=2}^{\infty}\{({\mathcal{A}}-\mathcal{P}_{1})^{t}\}_{kk}=o_{p}(1),$
where the second equality uses the fact that ${\mathbb{E}}[\nabla l({\bm{u}}^{*})\{\nabla l({\bm{u}}^{*})\}^{\top}]=-H^{*}({\bm{u}}^{*})$ and the last equality follows from statement (3) in Lemma 13. The proof is finished by applying Chebyshev’s inequality.
7 Discussion
Despite the broad applicability of the established result in terms of model parameterization, there are practical considerations that motivate further exploration of the problem. For example, in scenarios where multiple data sources are available for the same subjects of interest, different pairwise comparison models might be employed. A promising avenue for future investigation involves integrating the estimated information derived from different yet related models. The second aspect pertains to temporal dynamics. As can be seen in the ATP example, the data span nearly fifty years. In such contexts, it is more appropriate to also account for the temporal effects on the statistical models employed to gauge comparison outcomes and on the sampling of the graph. Addressing these considerations may involve incorporating additional covariate information. We intend to pursue these directions in our future work.
Appendices
Appendix A Proof of Lemma 13
We start by making two observations. The first states that the spectral gap of ${\mathcal{L}}_{sym}$ can be bounded using the degree of heterogeneity of the graph. The second provides an upper bound for ${\mathcal{D}}^{-1/2}\{H^{*}({\bm{u}}^{*})-H({\bm{u}}^{*})\}({\widehat{\bm{u}}}-{\bm{u}}^{*})$ and ${\mathcal{D}}^{-1/2}\{H({\bm{u}}^{*})-\bar{H}({\bm{W}})\}({\widehat{\bm{u}}}-{\bm{u}}^{*})$ in the $\ell_{\infty}$ norm. These two observations are summarized as the following two lemmas respectively. Let $\kappa_{n}=c_{n,4}/c_{n,3}\in[1,\infty)$.
Lemma 14
If $\lim_{n\to\infty}(\kappa^{7}_{n}q^{6}_{n}\log n)/(np^{7}_{n})=0$, then there exists an absolute constant $C$ such that $\|{\mathcal{A}}-\mathcal{P}_{1}\|_{2}\leq 1-C\{p_{n}/(\kappa_{n}q_{n})\}^{2}$.
Lemma 15
If Assumptions 47 hold, then
$\|{\mathcal{D}}^{-1/2}\{H^{*}({\bm{u}}^{*})-H({\bm{u}}^{*})\}({\widehat{\bm{u}}}-{\bm{u}}^{*})\|_{\infty}=O_{p}\left(\frac{c_{n,2}c_{n,4}^{3}}{c_{n,3}^{7/2}}\left\{\frac{q_{n}^{5}(\log n)^{4}}{np_{n}^{6}}\right\}^{1/2}\right)$
$\|{\mathcal{D}}^{-1/2}\{H({\bm{u}}^{*})-\bar{H}({\bm{W}})\}({\widehat{\bm{u}}}-{\bm{u}}^{*})\|_{\infty}=O_{p}\left(\frac{c^{2}_{n,2}c_{n,5}}{c_{n,3}^{5/2}}\left\{\frac{q_{n}^{5}(\log n)^{6}}{np_{n}^{6}}\right\}^{1/2}\right)$
The proofs of Lemmas 1415 are deferred to the end. We now prove Lemma 13 assuming both Lemmas 1415 hold. We begin with the first two statements. Let
$\bm{\zeta}\in\left\{{\mathcal{D}}^{-1/2}\{H^{*}({\bm{u}}^{*})-H({\bm{u}}^{*})\}({\widehat{\bm{u}}}-{\bm{u}}^{*}),{\mathcal{D}}^{-1/2}\{H({\bm{u}}^{*})-\bar{H}({\bm{W}})\}({\widehat{\bm{u}}}-{\bm{u}}^{*})\right\}.$
Recall that
${\mathcal{D}}_{ii}=\sum_{j\in\delta\{i\}}H_{ij},\quad{\mathcal{A}}_{ij}=\frac{H_{ij}}{{{\mathcal{D}}_{ii}^{1/2}{\mathcal{D}}_{jj}^{1/2}}},\quad\{\mathcal{P}_{1}\}_{ij}=\frac{{{\mathcal{D}}_{ii}^{1/2}{\mathcal{D}}_{jj}^{1/2}}}{\sum_{k\in[n]}{\mathcal{D}}_{kk}}.$ (13)
Note the all-ones vector $\bm{1}_{n}$ is in the zero-eigenspace of $H({\bm{u}}^{*})$ or $H^{*}({\bm{u}}^{*})$ (in both the left- and right-eigenspace since $H({\bm{u}}^{*})$ and $H^{*}({\bm{u}}^{*})$ are symmetric). As a result,
$\mathcal{P}_{1}{\mathcal{D}}^{-1/2}\{H^{*}({\bm{u}}^{*})-H({\bm{u}}^{*})\}({\widehat{\bm{u}}}-{\bm{u}}^{*})$
$=$ $\ \frac{1}{\sum_{i\in[n]}{\mathcal{D}}_{ii}}({\mathcal{D}}_{11}^{1/2},\ldots,{\mathcal{D}}_{nn}^{1/2})^{\top}\bm{1}^{\top}_{n}\{H^{*}({\bm{u}}^{*})-H({\bm{u}}^{*})\}({\widehat{\bm{u}}}-{\bm{u}}^{*})=0.$
On the other hand, since $\nabla l({\bm{u}}^{*})-\nabla l({\widehat{\bm{u}}})=-\bar{H}({\bm{W}})({\widehat{\bm{u}}}-{\bm{u}}^{*})$,
$-\mathcal{P}_{1}{\mathcal{D}}^{-1/2}\bar{H}({\bm{W}})({\widehat{\bm{u}}}-{\bm{u}}^{*})=\frac{1}{\sum_{i\in[n]}{\mathcal{D}}_{ii}}({\mathcal{D}}_{11}^{1/2},\ldots,{\mathcal{D}}_{nn}^{1/2})^{\top}\bm{1}^{\top}_{n}(\nabla l({\bm{u}}^{*})-\nabla l({\widehat{\bm{u}}}))$
$=$ $\ \frac{1}{\sum_{i\in[n]}{\mathcal{D}}_{ii}}\left\{\sum_{i\in[n]}\sum_{j\in\delta\{i\}}g(X_{ij},\widehat{u}_{i}-\widehat{u}_{j})-g(X_{ij},u_{i}^{*}-u_{j}^{*})\right\}({\mathcal{D}}_{11}^{1/2},\ldots,{\mathcal{D}}_{nn}^{1/2})^{\top}=0,$
which follows as a result of the validity assumption on $f$ (that implies $g(x;y)+g(-x;-y)=0$). Combining both pieces, we obtain $\mathcal{P}_{1}\bm{\zeta}=0$. Meanwhile, since ${\mathcal{L}}_{sym}^{\dagger}=\sum_{t=0}^{\infty}({\mathcal{A}}-\mathcal{P}_{1})^{t}-\mathcal{P}_{1}$, it follows from the Cauchy–Schwarz inequality that
$\|{\mathcal{L}}^{\dagger}_{sym}\bm{\zeta}\|_{\infty}=$ $\left\|\sum_{t=0}^{\infty}({\mathcal{A}}-\mathcal{P}_{1})^{t}\bm{\zeta}\right\|_{\infty}$ (14)
[…]
If choosing $L=C(\kappa_{n}q_{n}/p_{n})^{2}\log n$ for some sufficiently large constant $C>0$, then the second term in (15) is $o_{p}(1)$ as a result of Lemma 14. For the first term in (15), it follows from the direct computation using (13) that ${\mathcal{A}}^{t}:{\mathbb{R}}^{n}\to{\mathbb{R}}^{n}$, viewed as an operator in the $\ell_{\infty}$ norm, can be bounded as
$\|{\mathcal{A}}^{t}\|_{\ell_{\infty}\to\ell_{\infty}}$ $=\max_{i\in[n]}\sum_{j\in[n]}|\{{\mathcal{A}}^{t}\}_{ij}|$ (16)
(13)${{\leq}}\max_{i\in[n]}\sum_{j_{1}\in[n]}\cdots\sum_{j_{t-1}\in[n]}\sum_{j\in[n]}\frac{H_{ij_{1}}\cdots H_{j_{t-1}j}}{{{\mathcal{D}}_{ii}^{1/2}}{\mathcal{D}}_{j_{1}j_{1}}\cdots{\mathcal{D}}_{j_{t-1}j_{t-1}}{{\mathcal{D}}_{jj}^{1/2}}}$
$\leq\max_{i\in[n]}\frac{1}{\min_{k\in[n]}{{\mathcal{D}}_{kk}^{1/2}}}\sum_{j_{1}\in[n]}\cdots\sum_{j_{t-1}\in[n]}\frac{H_{ij_{1}}\cdots H_{j_{t-1}j}}{{{\mathcal{D}}_{ii}^{1/2}}{\mathcal{D}}_{j_{1}j_{1}}\cdots{\mathcal{D}}_{j_{t-1}j_{t-1}}}$
$=\frac{\max_{i\in[n]}{{\mathcal{D}}_{ii}^{1/2}}}{\min_{k\in[n]}{{\mathcal{D}}_{kk}^{1/2}}}\lesssim\left(\frac{\kappa_{n}q_{n}}{p_{n}}\right)^{1/2}.$
Consequently,
$\left\|\sum_{t=0}^{L}{\mathcal{A}}^{t}\bm{\zeta}\right\|_{\infty}\lesssim L\left({\frac{\kappa_{n}q_{n}}{p_{n}}}\right)^{1/2}\|\bm{\zeta}\|_{\infty}\lesssim(\log n)\cdot\left(\frac{\kappa_{n}q_{n}}{p_{n}}\right)^{5/2}\|\bm{\zeta}\|_{\infty}=o_{p}(1),$
where the last step follows from Lemma 15 and the assumptions in the lemma.
To prove the last statement in Lemma 13, note that since $({\mathcal{A}}-\mathcal{P}_{1})^{t}$ is semi-positive definite, $\{({\mathcal{A}}-\mathcal{P}_{1})^{t}\}_{ii}\leq\|({\mathcal{A}}-\mathcal{P}_{1})^{t}\|_{2}$ for every $t$. By a similar truncation argument as above, we only need to prove for summation of $t$ from $2$ to $L$, and this can be verified by a direct computation as in (16):
$\max_{i\in[n]}\sum_{t=2}^{L}\{({\mathcal{A}}-\mathcal{P}_{1})^{t}\}_{ii}\lesssim\frac{L}{({c_{n,3}np_{n}})^{1/2}}\lesssim\frac{c_{n,4}^{2}}{c_{n,3}^{5/2}}\left\{\frac{q_{n}^{4}(\log n)^{2}}{np_{n}^{5}}\right\}^{1/2}=o_{p}(1),$
where the details are omitted.
Appendix B Proofs of Lemmas 1415
Before showing Lemmas 1415, we first cite the uniform consistency result in Han et al. (2023):
Lemma 16
Let ${\mathcal{G}}\sim{\mathsf{G}}(n,p_{n},q_{n})$. Under Assumptions 4-6, if
$\alpha_{n}=\frac{c_{n,2}}{c_{n,3}}\left\{\frac{q_{n}^{2}(\log n)^{3}}{np_{n}^{3}}\right\}^{1/2}\to 0\qquad n\to\infty,$
then there exists $C>0$ (independent of $n$) such that, with probability one, the MLE $\widehat{{\bm{u}}}$ uniquely exists and $\|\widehat{{\bm{u}}}-{\bm{u}}^{*}\|_{\infty}\leq C\alpha_{n}$ for all sufficiently large $n$.
B.1 Proof of Lemma 14
Denote eigenvalues of $\mathcal{L}_{\text{sym }}=I-\mathcal{A}=I-\mathcal{D}^{-1/2}\mathcal{W}\mathcal{D}^{-1/2}$ in increasing order as $0=\lambda_{1}\leq\cdots\leq\lambda_{n}\leq 2$, where ${\mathcal{D}}$ and $\mathcal{W}$ are all random matrices (where randomness is from the comparison graph ${\mathcal{G}}$ only). In particular, $\mathcal{W}$ is an off-diagonal weight matrix with $\mathcal{W}_{ij}=\mathbb{I}_{\{(i,j)\in{\mathcal{E}}\}}z_{ij}$, where
$z_{ij}$ $={\mathbb{E}}_{X_{ij}}\left[-\partial_{2}g(X_{ij},u^{*}_{i}-u^{*}_{j})\right]\in[c_{n,3},c_{n,4}],\quad i\neq j$
due to Assumption 6, with ${\mathbb{E}}_{X_{ij}}[\cdot]$ denoting expectation for $X_{ij}$.
It is clear $\left\|\mathcal{A}-\mathcal{P}_{1}\right\|_{2}=\max\left\{1-\lambda_{2},\lambda_{n}-1\right\}$.
To analyze the spectrum of $\mathcal{L}_{\text{sym }}$, we consider its expected version first. Let $\overline{\mathcal{D}}$ and $\overline{\mathcal{W}}$ be the expectation of $\mathcal{D}$ and $\mathcal{W}$, respectively, and ¯$\bar{\mathcal{L}}_{\text{sym }}=I-\bar{\mathcal{D}}^{-1/2}\bar{\mathcal{W}}\bar{\mathcal{D}}^{-1/2}$. By definition, ¯$\bar{\mathcal{L}}_{\text{sym }}$ is also a normalized graph Laplacian matrix; we denote the eigenvalues of ¯$\overline{\mathcal{L}}_{\text{sym }}$ by $0=\bar{\lambda}_{1}\leq\cdots\leq\bar{\lambda}_{n}\leq 2$. A computation based on Cheeger’s inequalities yields a lower bound for the spectral gap of the expected Laplacian (Chung, 1997; Bauer and Jost, 2013; Han and Xu, 2023): There exists an absolute constant $c>0$ such that
$\max\left\{1-\bar{\lambda}_{2},\bar{\lambda}_{n}-1\right\}\leq 1-c\left(\frac{p_{n}}{\kappa_{n}q_{n}}\right)^{2}.$ (17)
We next apply a concentration argument to show that the spectra of $\mathcal{L}_{\text{sym }}$ and ¯$\bar{\mathcal{L}}_{\text{sym }}$ are close. To this end, we first apply Weyl’s inequality to obtain
$\max\left\{|(1-\lambda_{2})-(1-\bar{\lambda}_{2})|,|(\lambda_{n}-1)-(\bar{\lambda}_{n}-1)|\right\}$
$\leq$ $\left\|\mathcal{L}_{\text{sym }}-\overline{\mathcal{L}}_{\text{sym }}\right\|_{2}=\left\|\mathcal{D}^{-1}\mathcal{W}-\overline{\mathcal{D}}^{-1}\overline{\mathcal{W}}\right\|_{2}$
$\leq$ $\left\|\left(\mathcal{D}^{-1}-\overline{\mathcal{D}}^{-1}\right)\overline{\mathcal{W}}\right\|_{2}+\left\|\mathcal{D}^{-1}(\mathcal{W}-\overline{\mathcal{W}})\right\|_{2}$
$\leq$ $\left\|\mathcal{D}^{-1}-\overline{\mathcal{D}}^{-1}\right\|_{2}\|\overline{\mathcal{W}}\|_{2}+\left\|\mathcal{D}^{-1}\right\|_{2}\|\mathcal{W}-\overline{\mathcal{W}}\|_{2}.$
By the Chernoff bound, with probability at least $1-n^{-2}$, all $D_{ii}$ are concentrated around their means
$\left|\mathcal{D}_{ii}-\overline{\mathcal{D}}_{ii}\right|\lesssim\left({c_{n,4}\overline{\mathcal{D}}_{ii}\log n}\right)^{1/2}\quad i\in[n],$
so that
$\left\|\mathcal{D}^{-1}-\overline{\mathcal{D}}^{-1}\right\|_{2}$ $=\max_{i\in[n]}\left|\mathcal{D}_{ii}^{-1}-\overline{\mathcal{D}}_{ii}^{-1}\right|\lesssim\max_{i\in[n]}\left(\overline{\mathcal{D}}_{ii}^{-2}\left|\mathcal{D}_{ii}-\overline{\mathcal{D}}_{ii}\right|\right)\lesssim\left\{\frac{c_{n,4}\log n}{(c_{n,3}np_{n})^{3}}\right\}^{1/2}$
$\left\|\mathcal{D}^{-1}\right\|_{2}$ $\lesssim\left\|\overline{\mathcal{D}}^{-1}\right\|_{2}\lesssim\frac{1}{c_{n,3}np_{n}}.$
On the other hand,
$\|\overline{\mathcal{W}}\|_{2}=\left\|\overline{\mathcal{D}}\overline{\mathcal{D}}^{-1}\overline{\mathcal{W}}\right\|_{2}\leq\|\overline{\mathcal{D}}\|_{2}\left\|\overline{\mathcal{D}}^{-1}\overline{\mathcal{W}}\right\|_{2}\leq 2\|\overline{\mathcal{D}}\|_{2}=2\max_{i\in[n]}\overline{\mathcal{D}}_{ii}\lesssim c_{n,4}nq_{n}.$
By the matrix Bernstein inequality (Tropp, 2012), with probability at least $1-n^{-2}$,
$\|\mathcal{W}-\overline{\mathcal{W}}\|_{2}\lesssim c_{n,4}\left({nq_{n}\log n}\right)^{1/2}.$
In conclusion, the upper bound for $\left\|\mathcal{L}_{\text{sym}}-\overline{\mathcal{L}}_{\text{sym}}\right\|_{2}$ is
$\left\|\mathcal{L}_{\text{sym }}-\overline{\mathcal{L}}_{\text{sym}}\right\|_{2}\lesssim\left(\frac{\kappa_{n}^{3}q_{n}^{2}\log n}{np_{n}^{3}}\right)^{1/2}+\left(\frac{\kappa_{n}^{2}q_{n}\log n}{np_{n}^{2}}\right)^{1/2}.$ (18)
The desired result follows if the expected spectral gap in (17) dominates the concentration deviations (18):
$\frac{\left(\frac{\kappa_{n}^{3}q_{n}^{2}\log n}{np_{n}^{3}}\right)^{1/2}+\left(\frac{\kappa_{n}^{2}q_{n}\log n}{np_{n}^{2}}\right)^{1/2}}{\left(\frac{p_{n}}{\kappa_{n}q_{n}}\right)^{2}}\to 0\quad n\to\infty,$
which holds true if $\lim_{n\to\infty}(\kappa_{n}^{7}q_{n}^{6}\log n)/(np_{n}^{7})\to 0$.
B.2 Proof of Lemma 15
Recall the formula for the Hessian matrix $H({\bm{u}})$ and $\bar{H}({\bm{W}})$:
$\{H({\bm{u}})\}_{ij}$ =$=\left\{\begin{array}[]{ll}\partial_{ij}l({\bm{u}})=-\partial_{2}g(X_{ij},u_{i}-u_{j})\times\mathbb{I}_{\{(i,j)\in{\mathcal{E}}\}}&\quad i\neq j\\ -\sum_{k\in\delta\{i\}}\partial_{ik}l({\bm{u}})&\quad i=j\end{array}\right.$ (21)
$\{\bar{H}({\bm{W}})\}_{ij}$ =$=\left\{\begin{array}[]{ll}\partial_{ij}l(\bm{w}_{i})=-\partial_{2}g(X_{ij},w_{ii}-w_{ij})\times\mathbb{I}_{\{(i,j)\in{\mathcal{E}}\}}&\quad i\neq j\\ -\sum_{k\in\delta\{i\}}\partial_{ik}l(\bm{w}_{i})&\quad i=j\end{array}\right.$ (24)
where ${\bm{w}}_{1},\ldots,{\bm{w}}_{n}$ are $n$ points on the line segment between ${\widehat{\bm{u}}}$ and ${\bm{u}}^{*}$, and ${\bm{w}}_{i}=(w_{i1},\ldots,w_{in})\in{\mathbb{R}}^{n}$. Write the $i$th element of $\{\bar{H}({\bm{W}})-H\left({\bm{u}}^{*}\right)\}\left(\widehat{{\bm{u}}}-{\bm{u}}^{*}\right)$ as
|[$\left|\left[\{\bar{H}({\bm{W}})-H\left({\bm{u}}^{*}\right)\}\left(\widehat{{\bm{u}}}-{\bm{u}}^{*}\right)\right]_{i}\right|$
$=$ $\left|\sum_{j\neq i}\left\{\bar{H}({\bm{W}})-H\left({\bm{u}}^{*}\right)\right\}_{ij}(\widehat{u}_{j}-u^{*}_{j})+\left\{\bar{H}({\bm{W}})-H\left({\bm{u}}^{*}\right)\right\}_{ii}(\widehat{u}_{i}-u^{*}_{i})\right|$
$=$ $\left|\sum_{j\neq i}\left\{\bar{H}({\bm{W}})-H\left({\bm{u}}^{*}\right)\right\}_{ij}(\widehat{u}_{j}-u^{*}_{j})-\sum_{j\neq i}\left\{\bar{H}({\bm{W}})-H\left({\bm{u}}^{*}\right)\right\}_{ij}(\widehat{u}_{i}-u^{*}_{i})\right|$
$=$ $\left|\sum_{j\neq i}\left\{\bar{H}({\bm{W}})-H\left({\bm{u}}^{*}\right)\right\}_{ij}\left\{\left(\widehat{u}_{j}-u_{j}^{*}\right)-\left(\widehat{u}_{i}-u_{i}^{*}\right)\right\}\right|$
By the mean value theorem, for every $i,j$ and all sufficiently large $n$, there exists $\xi_{ij}\in[-M_{n}-1,M_{n}+1]$ depending on $i,j$ such that
|$\left|\left\{\bar{H}({\bm{W}})-H\left({\bm{u}}^{*}\right)\right\}_{ij}\right|$ =(21)|{${{=}}\left|\left\{\partial_{2}g(X_{ij},u^{*}_{i}-u^{*}_{j})-\partial_{2}g(X_{ij},w_{ii}-w_{ij})\right\}\times\mathbb{I}_{\{(i,j)\in{\mathcal{E}}\}}\right|$
$=\left|\partial_{22}g(X_{ij},\xi_{ij})\left\{\left(u_{i}^{*}-w_{ii}\right)-\left(u_{j}^{*}-w_{ij}\right)\right\}\times\mathbb{I}_{\{(i,j)\in{\mathcal{E}}\}}\right|$
$\leq 2c_{n,5}\|\bm{w}_{i}-{\bm{u}}^{*}\|_{\infty}\times\mathbb{I}_{\{(i,j)\in{\mathcal{E}}\}}$
$\leq 2c_{n,5}\|{\widehat{\bm{u}}}-{\bm{u}}^{*}\|_{\infty}\times\mathbb{I}_{\{(i,j)\in{\mathcal{E}}\}},$
where the penultimate inequality follows from Assumption 7. Summing over $j\in[n]$,
|[$\left|\left[\left\{\bar{H}({\bm{W}})-H\left({\bm{u}}^{*}\right)\right\}\left(\widehat{{\bm{u}}}-{\bm{u}}^{*}\right)\right]_{i}\right|\leq 4c_{n,5}\|\widehat{{\bm{u}}}-{\bm{u}}^{*}\|^{2}_{\infty}|\delta\{i\}|.$
Combining the results with Lemma 16, we can have
$\left|\left[D^{-1/2}\left\{\bar{H}({\bm{W}})-H\left({\bm{u}}^{*}\right)\right\}\left(\widehat{{\bm{u}}}-{\bm{u}}^{*}\right)\right]_{i}\right|\leq[D^{-1/2}]_{ii}\left|\left[\left\{\bar{H}({\bm{W}})-H\left({\bm{u}}^{*}\right)\right\}\left(\widehat{{\bm{u}}}-{\bm{u}}^{*}\right)\right]_{i}\right|$
$\leq 4c_{n,5}[D^{-1/2}]_{ii}\|\widehat{{\bm{u}}}-{\bm{u}}^{*}\|^{2}_{\infty}\left({|\delta\{i\}|}\right)^{1/2}\leq\frac{c^{2}_{n,2}c_{n,5}}{c^{5/2}_{n,3}}\frac{q_{n}^{5/2}(\log n)^{3}}{n^{1/2}p_{n}^{3}}.$
The proof of the first statement is completed. Next, we prove the second statement in Lemma 15. Similar to the previous proof, we can decompose |[$\left\lvert[\{H({\bm{u}}^{*})-H^{\ast}({\bm{u}}^{\ast})\}({\widehat{\bm{u}}}-{\bm{u}}^{*})]_{i}\right\lvert$ into two parts and obtain
|[$\left\lvert[\{H({\bm{u}}^{*})-H^{\ast}({\bm{u}}^{\ast})\}({\widehat{\bm{u}}}-{\bm{u}}^{*})]_{i}\right\lvert$
$\leq$ $\left|\sum_{j\neq i}\left\{H({\bm{u}}^{*})-H^{\ast}({\bm{u}}^{\ast})\right\}_{ij}(\widehat{u}_{j}-u^{*}_{j})\right|+\left|\sum_{j\neq i}\left\{H({\bm{u}}^{*})-H^{\ast}({\bm{u}}^{\ast})\right\}_{ij}(\widehat{u}_{i}-u^{*}_{i})\right|$
$\leq$ $\underbrace{\left|\sum_{j\neq i}\left\{H({\bm{u}}^{*})-H^{\ast}({\bm{u}}^{\ast})\right\}_{ij}(\widehat{u}_{j}-u^{*}_{j})\right|}_{(A.1)}+\underbrace{\left|\sum_{j\neq i}\left\{H({\bm{u}}^{*})-H^{\ast}({\bm{u}}^{\ast})\right\}_{ij}\right|\|{\widehat{\bm{u}}}-{\bm{u}}^{*}\|_{\infty}}_{(A.2)}.$
[…]
$\left|\sum_{j\neq i}\left\{H({\bm{u}}^{*})-H^{\ast}({\bm{u}}^{\ast})\right\}_{ij}\right|\lesssim c_{n,4}\left(|\delta\{i\}|\log n\right)^{1/2}$
Therefore,
$(A.2)\lesssim c_{n,4}\left(|\delta\{i\}|\log n\right)^{1/2}\times\|{\widehat{\bm{u}}}-{\bm{u}}^{\ast}\|_{\infty}\lesssim\left({|\delta\{i\}|}\right)^{1/2}\times\frac{c_{n,2}c_{n,4}q_{n}(\log n)^{2}}{c_{n,3}np_{n}^{3/2}}.$ (25)
[…]
$\left\{H({\bm{u}}^{*})-H^{\ast}({\bm{u}}^{\ast})\right\}_{ij}$ and $\widehat{u}_{j}-u^{*}_{j}$. To address this issue, we apply a leave-one-out approach based on a modification of Gao et al. (2023). For $i\in[n]$, let
$l^{(-i)}({\bm{u}})=\sum_{j\neq i,k\neq i,(j,k)\in{\mathcal{E}}}\log f(X_{jk},u_{j}-u_{k}).$ (26)
We define $\widehat{{\bm{u}}}^{(-i)}\in{\mathbb{R}}^{n-1}$ as leave-one-out estimator by maximizing the likelihood
${\widehat{\bm{u}}}^{(-i)}=\operatorname*{arg\,max}_{{\bm{u}}\in{\mathbb{R}}^{n-1}:\bm{1}_{n-1}^{\top}{\bm{u}}=-u^{*}_{i}}l^{(-i)}({\bm{u}}).$
We also denote ${\bm{u}}_{-i}\in{\mathbb{R}}^{n-1}$ as ${\bm{u}}$ deleting the $i$th component. Note that ${\widehat{\bm{u}}}^{(-i)}$ is also uniformly consistent, that is,
$\|{\widehat{\bm{u}}}^{(-i)}-{\bm{u}}_{-i}^{\ast}\|_{\infty}\lesssim\frac{c_{n,2}}{c_{n,3}}\left\{\frac{q_{n}^{2}(\log n)^{3}}{np_{n}^{3}}\right\}^{1/2}.$
[…]
$(J.2)\lesssim c_{n,4}\left(|\delta\{i\}|\log n\right)^{1/2}\times\|{\widehat{\bm{u}}}^{(-i)}-{\bm{u}}_{-i}^{\ast}\|_{\infty}.$ (28)
[…]
$\lesssim c_{n,4}\left({|\delta\{i\}|}\right)^{1/2}\times\|{\widehat{\bm{u}}}_{-i}-{\widehat{\bm{u}}}^{(-i)}\|_{2},$ (29)
where $\left\{H({\bm{u}}^{*})-H^{\ast}({\bm{u}}^{\ast})\right\}_{i,:}$ denotes the $i$th row of $H({\bm{u}}^{*})-H^{\ast}({\bm{u}}^{\ast})$. To further upper bound $\|{\widehat{\bm{u}}}_{-i}-{\widehat{\bm{u}}}^{(-i)}\|_{2}$, we use a Taylor expansion as follows:
$(\widehat{{\bm{u}}}^{(-i)}-\widehat{{\bm{u}}}_{-i})^{\top}\nabla l^{(-i)}(\widehat{{\bm{u}}}_{-i})$ $=-(\widehat{{\bm{u}}}^{(-i)}-\widehat{{\bm{u}}}_{-i})^{\top}\left\{\nabla l^{(-i)}(\widehat{{\bm{u}}}^{(-i)})-\nabla l^{(-i)}(\widehat{{\bm{u}}}_{-i})\right\}$
$=(\widehat{{\bm{u}}}^{(-i)}-\widehat{{\bm{u}}}_{-i})^{\top}\{-\nabla^{2}l^{(-i)}(\widetilde{{\bm{u}}})\}(\widehat{{\bm{u}}}^{(-i)}-\widehat{{\bm{u}}}_{-i}),$
where $\widetilde{{\bm{u}}}$ lies on the line segment between $\widehat{{\bm{u}}}^{(-i)}$ and $\widehat{{\bm{u}}}_{-i}$. Applying the Cauchy–Schwarz inequality,
$\|\widehat{{\bm{u}}}^{(-i)}-\widehat{{\bm{u}}}_{-i}\|_{2}\leq\frac{\lVert\nabla l^{(-i)}(\widehat{{\bm{u}}}_{-i})\lVert_{2}}{\lambda_{2}\{-\nabla^{2}l^{(-i)}(\widetilde{{\bm{u}}})\}},$ (30)
where $\lambda_{2}\{\cdot\}$ (with some abuse of notation) denotes the second smallest eigenvalue. Since $\nabla l(\widehat{{\bm{u}}})=0$, separating the terms involving $i$ and the rest apart, we have
$\|\nabla l^{(-i)}(\widehat{{\bm{u}}}_{-i})\|_{2}^{2}=\sum_{j\neq i,(i,j)\in{\mathcal{E}}}g(X_{ij},\widehat{u}_{i}-\widehat{u}_{j})^{2}$
$\lesssim\sum_{j\neq i,(i,j)\in{\mathcal{E}}}\left\{g(X_{ij},\widehat{u}_{i}-\widehat{u}_{j})-g(X_{ij},u^{*}_{i}-u^{*}_{j})\right\}^{2}+\sum_{j\neq i,(i,j)\in{\mathcal{E}}}g(X_{ij},u^{*}_{i}-u^{*}_{j})^{2}$
$\lesssim c_{n,4}nq_{n}\|{\widehat{\bm{u}}}-{\bm{u}}^{\ast}\|_{\infty}^{2}+\sum_{j\neq i,(i,j)\in{\mathcal{E}}}g(X_{ij},u^{*}_{i}-u^{*}_{j})^{2}.$
Under Assumption 5, $\{g(X_{ij},u^{*}_{i}-u^{*}_{j})\}_{j:(i,j)\in{\mathcal{E}}}$ are subgaussian random variables with subgaussian norms uniformly bounded by $c_{n,2}$. According to (Vershynin, 2018, Lemma 2.7.6), $\{g(X_{ij},u^{*}_{i}-u^{*}_{j})^{2}\}_{j:(i,j)\in{\mathcal{E}}}$ are sub-exponential random variables with sub-exponential norms uniformly bounded by $c_{n,2}^{2}$. It follows from (Vershynin, 2018, Proposition 2.7.1, Exercise 2.7.10) that
$\max_{j:(i,j)\in{\mathcal{E}}}{\mathbb{E}}[g(X_{ij},u^{*}_{i}-u^{*}_{j})^{2}]\lesssim c_{n,2}^{2}$
$\max_{j:(i,j)\in{\mathcal{E}}}\|g(X_{ij},u^{*}_{i}-u^{*}_{j})^{2}-{\mathbb{E}}[g(X_{ij},u^{*}_{i}-u^{*}_{j})^{2}]\|_{\psi_{1}}\lesssim c_{n,2}^{2},$
where $\|\cdot\|_{\psi_{1}}$ denotes the sub-exponential norm. By Bernstein’s inequality,
$\sum_{j\neq i,(i,j)\in{\mathcal{E}}}\{g(X_{ij},u^{*}_{i}-u^{*}_{j})\}^{2}$
$\lesssim\sum_{j\neq i,(i,j)\in{\mathcal{E}}}{\mathbb{E}}[g(X_{ij},u^{*}_{i}-u^{*}_{j})^{2}]$
$\qquad\qquad+\left|\sum_{j\neq i,(i,j)\in{\mathcal{E}}}\left[g(X_{ij},u^{*}_{i}-u^{*}_{j})^{2}-{\mathbb{E}}[g(X_{ij},u^{*}_{i}-u^{*}_{j})^{2}]\right]\right|$
$\leq 2\sum_{j\neq i,(i,j)\in{\mathcal{E}}}{\mathbb{E}}[g(X_{ij},u^{*}_{i}-u^{*}_{j})^{2}]\lesssim c_{n,2}^{2}nq_{n}$
holds with probability at least $1-n^{-3}$. Consequently, with probability at least $1-n^{-3}$,
$\|\nabla l^{(-i)}(\widehat{{\bm{u}}}_{-i})\|_{2}$$\lesssim\left\{(c_{n,4}\|{\widehat{\bm{u}}}-{\bm{u}}^{\ast}\|_{\infty}^{2}+c_{n,2}^{2})nq_{n}\right\}^{1/2}\lesssim c_{n,2}(nq_{n})^{1/2},$ (31)
where the last step follows from the observation
$c_{n,4}\|{\widehat{\bm{u}}}-{\bm{u}}^{\ast}\|_{\infty}^{2}\lesssim\frac{c_{n,2}^{2}c_{n,4}}{c_{n,3}^{2}}\frac{q_{n}^{2}\log n}{np_{n}^{3}}\to 0,$
as $n\to\infty.$
On the other hand, by a similar calculation as in the proof of Lemma 14 (for normalized Laplacian), with probability at least $1-n^{-2}$,
$\lambda_{2}\{\nabla^{2}l^{(-i)}(\widetilde{{\bm{u}}})\}\gtrsim c_{n,3}np_{n}\left(\frac{p_{n}}{\kappa_{n}q_{n}}\right)^{2},$ (32)
where the extra term $c_{n,3}np_{n}$ stems from a lower bound for the singular value of the normalization matrix. Combining (30)–(32), we conclude
$\|\widehat{{\bm{u}}}^{(-i)}-\widehat{{\bm{u}}}_{-i}\|_{2}\lesssim\frac{c_{n,2}c_{n,4}^{2}q_{n}^{5/2}}{c_{n,3}^{3}p_{n}^{3}n^{1/2}}.$ (33)
Putting (27), (28), (29), and (33) together,
[…]
$\lesssim\left({|\delta\{i\}|}\right)^{1/2}\times\frac{c_{n,2}c_{n,4}^{3}q_{n}^{5/2}(\log n)^{2}}{c_{n,3}^{3}p_{n}^{3}n^{1/2}}.$ (34)
Consequently, combining (25) and (34), we have
$|D^{-1/2}\{H({\bm{u}}^{*})-H^{\ast}({\bm{u}}^{\ast})\}({\widehat{\bm{u}}}-{\bm{u}}^{*})|_{i}$ $=D_{ii}^{-1/2}\left\lvert[\{H({\bm{u}}^{*})-H^{\ast}({\bm{u}}^{\ast})\}({\widehat{\bm{u}}}-{\bm{u}}^{*})]_{i}\right\lvert$
[…]
$\lesssim\frac{c_{n,2}c_{n,4}^{3}q_{n}^{5/2}(\log n)^{2}}{c_{n,3}^{7/2}p_{n}^{3}n^{1/2}}.$
The second part of Lemma 15 is proved.
Appendix C Proofs of Corollary 10 and Theorems 11-12
C.1 Proof of Corollary 10
Proof  If $M^{*}<\infty$, then $\sup_{n}c_{n,1}<1$ so that Assumption 4 holds under condition $(\log n)^{8}/(np_{n})\to 0$. Since ${\mathbb{A}}$ is finite and $g(x;y)$ is continuous in $y$, $g(x;y)$ is uniformly bounded over ${\mathbb{A}}\times[-M^{*},M^{*}]$, implying $\sup_{n}c_{n,2}<\infty$. Additionally, by the strict log-concavity of $f$, $\partial_{2}g(x;y)>0$, which combined with the continuity of $\partial_{2}g(x;y)$, the boundedness of $M^{*}$, and the finiteness of ${\mathbb{A}}$ implies $0<\inf_{n}c_{n,3}\leq\sup_{n}c_{n,4}<\infty$. A similar argument shows $\sup_{n}c_{n,5}<\infty$. The proof is finished by appealing to (3).
C.2 Proof of Theorem 11
According to (C.2) and (12), we have
${\mathcal{D}}^{1/2}({\widehat{\bm{u}}}-{\bm{u}}^{*})$ $=-{\mathcal{D}}^{-1/2}\nabla l({\bm{u}}^{*})-({\mathcal{L}}^{\dagger}_{sym}-I){\mathcal{D}}^{-1/2}\nabla l({\bm{u}}^{*})$
$+{\mathcal{L}}^{\dagger}_{sym}{\mathcal{D}}^{-1/2}\{H^{*}({\bm{u}}^{*})-H({\bm{u}}^{*})\}({\widehat{\bm{u}}}-{\bm{u}}^{*})$
$+{\mathcal{L}}^{\dagger}_{sym}{\mathcal{D}}^{-1/2}\{H({\bm{u}}^{*})-\bar{H}({\bm{W}})\}({\widehat{\bm{u}}}-{\bm{u}}^{*}).$
Then, for any $k\in[n]$, we have
$\{{\mathcal{D}}^{1/2}({\widehat{\bm{u}}}-{\bm{u}}^{*})\}_{k}$ =$=-\{{\mathcal{D}}^{-1/2}\nabla l({\bm{u}}^{*})\}_{k}+o_{p}(1),$
via Lemma 13 and Chebyshev’s inequality. Since ${\mathcal{D}}$ is diagonal matrix,
$\widehat{u}_{k}-u^{*}_{k}=-{\mathcal{D}}^{-1}_{kk}\{\nabla l({\bm{u}}^{*})\}_{k}+o_{p}({\mathcal{D}}^{-1/2}_{kk}).$
By Hoeffding’s inequality, we obtain
${\mathcal{D}}^{-1}_{kk}=O_{p}\left(\{c_{n,3}|\delta\{i\}|\}^{-1}\right),\ \{\nabla l({\bm{u}}^{*})\}_{k}=O_{p}(c_{n,2}\{|\delta\{i\}|\cdot\log n\}^{1/2}).$
Therefore,
$|\widehat{u}_{k}-u^{*}_{k}|=O_{p}\left(\frac{c_{n,2}}{c_{n,3}}\left\{\frac{\log n}{|\delta\{i\}|}\right\}^{1/2}\right).$
We finish the proof of Theorem 11.
C.3 Proof of Theorem 12
According to Slutsky’s Theorem and Thereom 8, it suffices to show that for any fixed $i\in[n]$ that
$\frac{\{\rho_{i}({\bm{u}}^{*})\}^{-1}-\{\rho_{i}({\widehat{\bm{u}}})\}^{-1}}{\{\rho_{i}({\bm{u}}^{*})\}^{-1}}=o_{p}(1).$
Note that $g(x,y)=\partial_{2}(\log f(x;y))$, we have
$\{\rho_{i}({\bm{u}}^{*})\}^{-1}=\sum_{j\in\delta\{i\}}\int_{{\mathbb{A}}}\{g(x;{u}^{*}_{i}-{u}^{*}_{j})\}^{2}f(x;{u}^{*}_{i}-{u}^{*}_{j})\operatorname{d}\!{x}$
and
$\{\rho_{i}({\bm{u}}^{*})\}^{-1}-\{\rho_{i}({\widehat{\bm{u}}})\}^{-1}$
$=$ $\sum_{j\in\delta\{i\}}\int_{{\mathbb{A}}}\{g(x;{u}^{*}_{i}-{u}^{*}_{j})\}^{2}f(x;{u}^{*}_{i}-{u}^{*}_{j})-\{g(x;\widehat{u}_{i}-\widehat{u}_{j})\}^{2}f(x;\widehat{u}_{i}-\widehat{u}_{j})\operatorname{d}\!{x}.$
For simplicity, let $g_{ij}^{*}(x):=g(x;{u}^{*}_{i}-{u}^{*}_{j})$, $f_{ij}^{*}(x):=f(x;{u}^{*}_{i}-{u}^{*}_{j})$, $\widehat{g}_{ij}(x):=g(x;\widehat{u}_{i}-\widehat{u}_{j})$ and $\widehat{f}_{ij}(x):=f(x;\widehat{u}_{i}-\widehat{u}_{j})$. It is enough to show that for any $j\in[n]$,
$\frac{\int_{{\mathbb{A}}}\{g_{ij}^{*}(x)\}^{2}f_{ij}^{*}(x)-\{\widehat{g}_{ij}(x)\}^{2}\widehat{f}_{ij}(x)\operatorname{d}\!{x}}{\int_{{\mathbb{A}}}\{g_{ij}^{*}(x)\}^{2}f_{ij}^{*}(x)\operatorname{d}\!{x}}=o_{p}(1).$
According to Hölder’s inequality,
$\left|\frac{\int_{{\mathbb{A}}}\{g_{ij}^{*}(x)\}^{2}f_{ij}^{*}(x)-\{\widehat{g}_{ij}(x)\}^{2}\widehat{f}_{ij}(x)\operatorname{d}\!{x}}{\int_{{\mathbb{A}}}\{g_{ij}^{*}(x)\}^{2}f_{ij}^{*}(x)\operatorname{d}\!{x}}\right|\leq\max_{x\in{\mathbb{A}}}\left|\frac{\{g_{ij}^{*}(x)\}^{2}f_{ij}^{*}(x)-\{\widehat{g}_{ij}(x)\}^{2}\widehat{f}_{ij}(x)}{\{g_{ij}^{*}(x)\}^{2}f_{ij}^{*}(x)}\right|.$
For the right-hand side,
$\frac{\{g_{ij}^{*}(x)\}^{2}f_{ij}^{*}(x)-\{\widehat{g}_{ij}(x)\}^{2}\widehat{f}_{ij}(x)}{\{g_{ij}^{*}(x)\}^{2}f_{ij}^{*}(x)}=\underbrace{\frac{\{g_{ij}^{*}(x)\}^{2}-\{\widehat{g}_{ij}(x)\}^{2}}{\{g_{ij}^{*}(x)\}^{2}}}_{\Gamma_{1}}+\underbrace{\frac{\{\widehat{g}_{ij}(x)\}^{2}}{\{g_{ij}^{*}(x)\}^{2}}\cdot\frac{f_{ij}^{*}(x)-\widehat{f}_{ij}(x)}{f_{ij}^{*}(x)}}_{\Gamma_{2}}.$
On the one hand, we decompose the term $\Gamma_{1}$ into two parts,
$\Gamma_{1}=\frac{g_{ij}^{*}(x)-\widehat{g}_{ij}(x)}{g_{ij}^{*}(x)}\cdot\frac{g_{ij}^{*}(x)+\widehat{g}_{ij}(x)}{g_{ij}^{*}(x)}.$
By the mean value theorem,
$\left|\frac{g_{ij}^{*}(x)-\widehat{g}_{ij}(x)}{g_{ij}^{*}(x)}\right|=\left|\frac{\partial_{2}g(x;w_{ij})}{g_{ij}^{*}(x)}\right|\cdot\left|({u}^{*}_{i}-{u}^{*}_{j})-(\widehat{u}_{i}-\widehat{u}_{j})\right|,$
where $w_{ij}$ is the intermediate value between $({u}^{*}_{i}-{u}^{*}_{j})$ and $(\widehat{u}_{i}-\widehat{u}_{j})$. Therefore, according to Lemma 16, we have with probability approaching one that
$\left|\frac{g_{ij}^{*}(x)-\widehat{g}_{ij}(x)}{g_{ij}^{*}(x)}\right|\lesssim\frac{c_{n,2}c_{n,4}}{c_{n,3}}\left\{\frac{q_{n}^{2}(\log n)^{3}}{np_{n}^{3}}\right\}^{1/2}=o(1),$
where the last step follows assuming (3) holds. Meanwhile,
$\left|\frac{g_{ij}^{*}(x)+\widehat{g}_{ij}(x)}{g_{ij}^{*}(x)}\right|=\left|2+\frac{\widehat{g}_{ij}(x)-g_{ij}^{*}(x)}{g_{ij}^{*}(x)}\right|=O_{p}(1).$
Therefore, $|\Gamma_{1}|=o_{p}(1).$ On the other hand, $\{\widehat{g}_{ij}(x)/g_{ij}^{*}(x)\}^{2}=O_{p}(1)$ following a similar approach as above. By the same reasoning,
$\log(f_{ij}^{*}(x))-\log(\widehat{f}_{ij}(x))=g(x;v_{ij})\{({u}^{*}_{i}-{u}^{*}_{j})-(\widehat{u}_{i}-\widehat{u}_{j})\}.$
Then, based on Assumption 5, Lemma 16, and Hoeffding’s inequality, the following inequlity
$|\log(f_{ij}^{*}(x))-\log(\widehat{f}_{ij}(x))|\leq\frac{c_{n,2}^{2}}{c_{n,3}}\left\{\frac{q_{n}^{2}(\log n)^{4}}{np_{n}^{3}}\right\}^{1/2}=o(1).$
holds with probability approaching one. As a result,
$\left|\frac{f_{ij}^{*}(x)-\widehat{f}_{ij}(x)}{f_{ij}^{*}(x)}\right|=o_{p}(1).$
Therefore, $|\Gamma_{2}|=o_{p}(1).$ The above proof is shown for fixed $x$. However, it is straightforward to extend it to the uniform case since we only rely on the event that Lemma 16 holds. Consequently,
$\max_{x\in{\mathbb{A}}}\left|\frac{\{g_{ij}^{*}(x)\}^{2}f_{ij}^{*}(x)-\{\widehat{g}_{ij}(x)\}^{2}\widehat{f}_{ij}(x)}{\{g_{ij}^{*}(x)\}^{2}f_{ij}^{*}(x)}\right|=o_{p}(1).$
We finish the proof of Theorem 12.

### v3

@@
 Abstract
-Pairwise comparison models have been widely used for utility evaluation and ranking across various fields.
-The increasing scale of problems today underscores the need to understand statistical inference in these models when the number of subjects diverges, a topic currently lacking in the literature except in a few special instances.
-To partially address this gap, this paper establishes a near-optimal asymptotic normality result for the maximum likelihood estimator in a broad class of pairwise comparison models, as well as a non-asymptotic convergence rate for each individual subject under comparison.
+Pairwise comparison models have been widely used for utility evaluation and rank aggregation across various fields.
+The increasing scale of modern problems underscores the need to understand statistical inference in these models when the number of subjects diverges, a topic that is currently underexplored in the literature.
+To address this gap, this paper establishes a near-optimal asymptotic normality result for the maximum likelihood estimator in a broad class of pairwise comparison models.
 The key idea lies in identifying the Fisher information matrix as a weighted graph Laplacian, which can be studied via a meticulous spectral analysis.
-Our findings provide a unified theory for performing statistical inference in a wide range of pairwise comparison models beyond the Bradley–Terry model, benefiting practitioners with theoretical guarantees for their use.
+Our findings provide theoretical foundations for performing statistical inference in a wide range of pairwise comparison models beyond the Bradley–Terry model.
 Simulations utilizing synthetic data are conducted to validate the asymptotic normality result, followed by a hypothesis test using a tennis competition dataset.
-Keywords: Imbalanced data, Maximum likelihood estimation, Ordinal data, Pairwise comparison models, Sparsity
+Keywords: Confidence interval; Maximum likelihood estimation; Ordinal data; Pairwise comparison; Sparsity.
 1 Introduction
-Pairwise comparison involves assessing subjects in pairs to establish their relative preferences, a practice pertinent to numerous applications such as sports analytics (Glickman and Jones, 1999; Ley et al., 2019; Tsokos et al., 2019), econometrics (McFadden, 1973), video coding (Lee et al., 2011), and social science (Loewen et al., 2012).
+Pairwise comparison involves assessing subjects in pairs to establish their relative preferences, a practice pertinent to various applications such as econometrics (eco1958; mcfadden1973conditional; 10.1257/pandp.20221063; lentz2023anatomy), sports analytics (baker2014dynamic; bozoki2016application; angelini2022weighted; collingwood2022evaluating), social science (loewen2012testing; 10.1111/rssa.12124; wapman2022quantifying), and, more recently, human preference studies aimed at improving models in artificial intelligence (Christiano2017; Rafailov2023; Sun2025).
 A prevalent approach for pairwise comparison modeling employs a latent score framework.
-Originated from the ideas in Thurstone (Thurstone, 1927) and Zermelo (Zermelo, 1929), a mathematical model for pairwise comparison data analysis was proposed by Bradley and Terry (Bradley and Terry, 1952).
-Since then, multiple generalizations have been developed, including ordinal models such as the Rao–Kupper model (Rao and Kupper, 1967) and the Davidson model (Davidson, 1970), which account for ties, the cumulative link model (Agresti, 1992) that considers more refined ordinal scales, and cardinal models such as the paired cardinal model (Shah et al., 2016).
-We recommend Cattelan (2012) for a review of pairwise comparison modeling from a practical perspective.
+Originated from the ideas in Thurstone (thurstone1927method) and Zermelo (MR1545015), a mathematical model for pairwise comparison data analysis was formulated by Bradley and Terry (MR0070925).
+Since then, multiple generalizations have been developed, including ordinal models such as the Rao–Kupper model (MR217963) and the Davidson model (davidson1970extending), which account for ties, the cumulative link model (agresti1992analysis) that considers more refined ordinal scales, and cardinal models such as the paired cardinal model (MR3504618).
+We recommend MR3012434 for a review of pairwise comparison modeling from a practical perspective.
 In light of the fast-growing scale of problems in the big-data era, a recent research area of pairwise comparison models focuses on understanding the asymptotics of estimating the latent score vector as the number of compared subjects approaches infinity.
-Along this line of work, considerable attention has been devoted to the study of the Bradley–Terry (BT) model, which is popular due to its simple parametrization form and connections to other problems such as matrix balancing (Qu et al., 2023) and completion (Chen et al., 2023).
-Notably, an efficient Minorize-Maximization (MM) algorithm for implementing the maximum likelihood estimator (MLE) was proposed in Hunter (2004) and accelerated in Newman (2023) in the BT model, and its uniform consistency was established when the comparison graph is dense (Simons and Yao, 1999; Yan et al., 2012).
-Subsequent works have established analogous consistency results for alternative estimators such as the spectral estimator and regularized MLE in the sparse settings under various error metrics (Chen and Suh, 2015; Shah et al., 2016; Negahban et al., 2017; Chen et al., 2019).
-The first result that generalized the uniform consistency of the MLE to the sparse setting appeared in Han et al.
-(2020), paralleled by a few other works with a slightly different focus (e.g.
-asymptotic efficiency and minimax rates); see Hendrickx et al.
-(2020); Bong and Rinaldo (2022); Chen et al.
-(2022).
-It is noteworthy that Han et al.
-(2020) also contains an asymptotic normality result for dense graphs, which was further explored in the sparse setting (Liu et al., 2023; Gao et al., 2023).
-Despite extensive research on the asymptotic properties of the BT model, a significant gap persists between theory and practice, as many comparison models used in applications either have multiple scales or a continuous outcome.
+Along this line of work, considerable attention has been devoted to the study of the Bradley–Terry (BT) model, which is popular due to its simple parametrization form and connections to other problems such as matrix balancing (qu2023sinkhorn) and completion (chen2023note).
+Notably, an efficient Minorize-Maximization (MM) algorithm for implementing the maximum likelihood estimator (MLE) was proposed in MR2051012 and accelerated in newman2023efficient in the BT model, and its uniform consistency was established when the comparison graph is dense (MR1724040; MR2987494).
+Subsequent works have established analogous consistency results for alternative estimators such as the spectral estimator and regularized MLE in the sparse settings under various error metrics (chen2015spectral; MR3504618; MR3613103; MR3953449).
+Recently, han2020asymptotic and chen2022optimal extended the uniform consistency of the MLE to the sparse setting, paralleled by a few other works with a slightly different focus (e.g., asymptotic efficiency and minimax rates); see hendrickx2020minimax; bong2022generalized.
+It is noteworthy that han2020asymptotic also contains an asymptotic normality result for dense graphs, which was further explored in the sparse setting (MR4560195; gao2023uncertainty).
+Despite extensive research on the asymptotic properties of the BT model, a gap persists between theory and practice as many comparison models used in applications either have multiple scales or a continuous outcome.
 This suggests a practical need to extend this understanding to other widely used yet less explored models.
-The recent work Han et al.
-(2023) partially addresses this by introducing a general pairwise comparison model framework and proving a uniform consistency result for the MLE therein.
+The recent work (han2022general) partially addresses this by introducing a general pairwise comparison model framework and proving a uniform consistency result for the MLE therein.
 However, an asymptotic normality result has not yet been established.
-Nonetheless, such a result is often crucial for practical statistical inference and cannot be deduced from the existing approaches to analyzing the BT model.
+Such a result is often crucial for practical inference and cannot be deduced from the existing approaches to analyzing the BT model.
 To take one step further in filling the gap, this paper aims to derive an asymptotic normality result for the MLE in the general pairwise comparison models.
@@
 Aside from enjoying the sufficient flexibility of model parametrization, our results additionally allow the incorporation of imbalanced data, which is new compared to many existing results in the field and may be of independent interest; see Table 1 for a detailed comparison between our result and some existing results in the literature.
-  • We verify that the MLE in many pairwise comparison models in the literature, including the BT model, the Thurstone–Mosteller model (Thurstone, 1927; Mosteller, 1951), the Davidson model, the Rao–Kupper model, the cumulative link model, and the paired cardinal model, is asymptotically normal under near-optimal sparsity conditions.
+  • We verify that the MLE in many pairwise comparison models in the literature, including the BT model, the Thurstone–Mosteller model (thurstone1927method; mosteller2006remarks), the Davidson model, the Rao–Kupper model, the cumulative link model, and the paired cardinal model, is asymptotically normal under near-optimal sparsity conditions.
 We also provide explicit and computable formulas for their asymptotic variance that may be of interest to practitioners using these models.
@@
 Results Models Comparison Graphs Sparsity
-Simons and Yao (1999) BT homogeneous $\Omega(1)$
-Yan et al. (2012) BT heterogeneous $\Omega(1)$
-Han et al. (2020) BT homogeneous $\widetilde{\Omega}(n^{-1/10})$
-Gao et al. (2023) BT homogeneous $\widetilde{\Omega}(n^{-1})$
+MR1724040 BT homogeneous $\Omega(1)$
+MR2987494 BT heterogeneous $\Omega(1)$
+han2020asymptotic BT homogeneous $\widetilde{\Omega}(n^{-1/10})$
+gao2023uncertainty BT homogeneous $\widetilde{\Omega}(n^{-1})$
 This work general heterogeneous $\widetilde{\Omega}(n^{-1})$
-Table 1: Comparison of several asymptotic normality results in the pairwise comparison model literature.
-Homogeneous comparison graphs indicate the balanced data while heterogeneous graphs correspond to the imbalanced data.
-The notation $\Omega(\cdot)$ is the standard Bachmann–Landau notation and $\widetilde{\Omega}(\cdot)$ means that the asymptotic relation holds up to polynomial terms of $\log n$.
+[…]
 The rest of the paper is organized as follows.
 In Section 2, we introduce the mathematical framework for general pairwise comparison models.
-In Section 3, we establish the main results concerning asymptotic normality and non-asymptotic convergence rates (Theorems 8, 11, 12) for the MLE and provide the relevant interpretations.
+In Section 3, we establish the main results concerning asymptotic normality and non-asymptotic convergence rates (Theorems 1, 2, 3) for the MLE and provide the relevant interpretations.
 In Section 4, we show that the MLE in many pairwise comparison models in the literature is asymptotically normal under minimal sparsity assumptions.
@@
 2 General pairwise comparison models
-We introduce the general pairwise comparison models proposed in Han et al.
-(2023).
+We introduce the general pairwise comparison models proposed in han2022general.
 Consider an undirected graph ${\mathcal{G}}=({\mathcal{V}},{\mathcal{E}})$ with $n$ vertices, where ${\mathcal{V}}=[n]:=\{1,\ldots,n\}$ represents the subjects being modeled, and $\mathcal{E}$} is the edge set of the comparison data, that is, for any $i,j\in{\mathcal{V}}$, $(i,j)\in{\mathcal{E}}$ if there exists a comparison between subjects $i$ and $j$.
@@
 In the general pairwise comparison models, the function $f$ is assumed to be valid in the sense of the following definition.
-Definition 1 (Valid parameterization)
+Definition 1 (Valid parameterization).
 A function $f:{\mathbb{A}}\times\mathbb{R}\to\mathbb{R}^{+}$, where ${\mathbb{A}}=-{\mathbb{A}}\subseteq{\mathbb{R}}$ denotes the possible comparison outcomes, is said to be valid if it satisfies the following assumptions:
 (A1). (normalization) For $y\in{\mathbb{R}}$, $\int_{\mathbb{A}}f(x;y)\operatorname{d}\!{x}=1$, where the integral is interpreted as summation if $\mathbb{A}$} is discrete.
-(A2). (symmetry) $f$ is even with respect to $(x;y)$, that is, $f(x;y)=f(-x;-y)$, for $(x,y)\in{\mathbb{A}}\times{\mathbb{R}}.$
+(A2). (symmetry) $f$ is even with respect to $(x;y)$, that is, f(x;y)=f(x;y)f(x;y)=f(-x;-y), for $(x,y)\in{\mathbb{A}}\times{\mathbb{R}}.$
 (A3). (monotonicity) For $x<0$, $f(x;y)$ is decreasing in $y$, and $f(x;y)\to 0$ as $y\to\infty$.
@@
 Having specified $f$, we can write down the likelihood function given the observed data and compute the MLE $\widehat{{\bm{u}}}$ for $\bm{u}$$^{*}$ under identifiability constraints; see Section 3.
-It was shown in Han et al.
-(2023) that $\widehat{{\bm{u}}}$ is a uniformly consistent estimator for ${\bm{u}}^{*}$ under appropriate conditions on both $f$ and ${\mathcal{G}}$.
+It was shown in han2022general that $\widehat{{\bm{u}}}$ is a uniformly consistent estimator for $\bm{u}$$^{*}$ under appropriate conditions on both $f$ and $\mathcal{G}$}.
 To identify graph configurations satisfying these conditions, it is often convenient to assume that $\mathcal{G}$} is sampled from certain random graph ensembles.
-A commonly used one is the random graph model ${\mathsf{G}}\left(n,p_{n},q_{n}\right)$ that generalizes the Erdős–Rényi model (Erdős and Rényi, 1960).
-Definition 2 (Random graph models)
+A commonly used one is the random graph model ${\mathsf{G}}\left(n,p_{n},q_{n}\right)$ that generalizes the Erdős–Rényi model (MR0125031).
+Definition 2 (Random graph models).
 ${\mathsf{G}}\left(n,p_{n},q_{n}\right)$ is a random graph with vertices set ${\mathcal{V}}=[n]$ where each edge $(i,j)\in{\mathcal{V}}\times{\mathcal{V}},i\neq j$ is formed independently with deterministic probability $p_{ij,n}\in\left[p_{n},q_{n}\right]$.
-Remark 3
+Remark 1.
 By definition, for fixed $p_{n}$ and $q_{n}$, different specifications of $\{p_{ij,n}\}_{1\leq i<j\leq n}$ may belong to the same ${\mathsf{G}}\left(n,p_{n},q_{n}\right)$.
 Such ambiguity is irrelevant to our discussion as only the upper and lower bounds are used in the analysis.
-In the following section, we assume that the comparison graph ${\mathcal{G}}({\mathcal{V}},{\mathcal{E}})$ is sampled from ${\mathsf{G}}\left(n,p_{n},q_{n}\right)$.
+In the following section, we assume the comparison graph ${\mathcal{G}}({\mathcal{V}},{\mathcal{E}})$ is sampled from ${\mathsf{G}}\left(n,p_{n},q_{n}\right)$.
 Meanwhile, for every $(i,j)\in{\mathcal{E}}$, we assume that a single comparison outcome $X_{ij}\sim f(x;u^{*}_{i}-u^{*}_{j})$ is observed.
@@
 $l({\bm{u}})=\sum_{(i,j)\in{\mathcal{E}}}\log f(X_{ij};u_{i}-u_{j}).$
-Note that $l({\bm{u}})$ is invariant if all components of ${\bm{u}}$ are translated by a constant.
+Note that $l({\bm{u}})$ is invariant if all components of $\bm{u}$} are shifted by a constant.
 To ensure model identifiability, we let ${\bm{1}}_{n}^{\top}{\bm{u}}=0$, where ${\bm{1}}_{n}$ is the all-ones vector with the same dimension as $\bm{u}$}.
-The MLE ${\widehat{\bm{u}}}$ for ${\bm{u}}^{*}$ satisfies
-${\widehat{\bm{u}}}=\operatorname*{arg\,max}_{{\bm{u}}\in{\mathbb{R}}^{n}:{\bm{1}}_{n}^{\top}{\bm{u}}=0}l({\bm{u}}).$ (1)
-To state the asymptotic normality result for $\widehat{{\bm{u}}}$, we need the following assumptions.
-Let $M_{n}=\max_{i,j\in[n]}|u^{*}_{i}-u^{*}_{j}|$ denote the dynamic range of ${\bm{u}}^{*}$.
+The MLE $^{\widehat{\bm{u}}}$ for $\bm{u}$$^{*}$ satisfies
+${\widehat{\bm{u}}}=\operatorname*{arg\,max}_{{\bm{u}}\in{\mathbb{R}}^{n}:{\bm{1}}_{n}^{\top}{\bm{u}}=0}l({\bm{u}}).$ (3.1)
+To state the asymptotic normality result for $\widehat{{\bm{u}}}$, we need the following assumptions. Let
+$M_{n}=\max_{i,j\in[n]}|u^{*}_{i}-u^{*}_{j}|$
+denote the dynamic range of $\bm{u}$$^{*}$.
 The global discrepancy of the model, which measures the probability of a subject with the highest score winning over another with the lowest score, is defined as
@@
 The first assumption, which ensures the unique existence of the MLE, states that $c_{n,1}$ if converging to one, must have a controlled rate.
-Assumption 4
+Assumption 1.
 The global discrepancy $c_{n,1}$ satisfies
-$\frac{\log n}{np_{n}|\log c_{n,1}|}\to 0\qquad n\to\infty.$ (2)
+$\frac{\log n}{np_{n}|\log c_{n,1}|}\to 0\qquad n\to\infty.$ (3.2)
 The next assumption is concerned with the tail decay of $g(X;y)$ with $X\sim f(x;y)$ for all $|y|\leq M_{n}$, where $g(X;y)$ is centered, that is, ${\mathbb{E}}[g(X;y)]=0$.
-Assumption 5
+Assumption 2.
 The sequence of random variables $\{g(X_{y};y)\}_{|y|\leq M_{n}}$ is uniformly subgaussian, where $X_{y}\sim f(x;y)$. That is,
 $c_{n,2}=\max_{|y|\leq M_{n}}\|g(X_{y};y)\|_{\psi_{2}}<\infty,$
-where $\|\cdot\|_{\psi_{2}}$ is the subgaussian norm (Vershynin, 2018, Definition 2.5.6).
+where $\|\cdot\|_{\psi_{2}}$ is the subgaussian norm (vershynin2018high, Definition 2.5.6).
 In addition, we further requires a boundedness condition on the partial derivative of $g(x;y)$ in $y$.
-Assumption 6
+Assumption 3.
 The partial derivative of the score function $g$ for $y$ is uniformly bounded from both above and below:
 $0<c_{n,3}=\inf_{x\in{\mathbb{A}},|y|\leq M_{n}+1}|\partial_{2}g(x;y)|\leq\sup_{x\in{\mathbb{A}},|y|\leq M_{n}+1}|\partial_{2}g(x;y)|=c_{n,4}<\infty.$
-Assumptions 4-6 are sufficient to guarantee the uniform consistency of the maximum likelihood estimator (Han et al., 2023).
+Assumptions 1-3 are sufficient to guarantee the uniform consistency of the maximum likelihood estimator (han2022general).
 The common pathway from consistency to asymptotic normality is via Taylor’s expansion.
 As a result, one often require more stringent regularity conditions on higher-order derivatives.
-Assumption 7
+Assumption 4.
 The second-order partial derivative of the score function $g(x,y)$ for $y$ is uniformly bounded for all $|y|\leq M_{n}+1$:
 $c_{n,5}=\sup_{x\in{\mathbb{A}},|y|\leq M_{n}+1}|\partial_{22}g(x;y)|<\infty.$
-Theorem 8 (Asymptotic normality)
+Theorem 1 (Asymptotic normality).
 Let ${\mathcal{G}}\sim{\mathsf{G}}(n,p_{n},q_{n})$. Under Assumptions 14, if
-$\beta_{n}=\max\left\{\frac{c_{n,2}^{2}c_{n,4}^{5/2}c_{n,5}}{c_{n,3}^{5}},\frac{c_{n,2}c_{n,4}^{11/2}}{c_{n,3}^{6}}\right\}\left\{\frac{q_{n}^{10}(\log n)^{8}}{np_{n}^{11}}\right\}^{1/2}\to 0\qquad n\to\infty,$ (3)
-then for each $i\in[n]$, $\{\rho_{i}({\bm{u}}^{*})\}^{-1/2}(\widehat{u}_{i}-u_{i}^{*})\to N(0,1)$, where $N(0,1)$ is the standard normal distribution and the asymptotic variance $\rho_{i}({\bm{u}}^{*})$ is given by
-$\rho_{i}({\bm{u}}^{*})$ $=\left[\sum_{j\in\delta\{i\}}\int_{{\mathbb{A}}}\frac{\{\partial_{2}f(x;u^{*}_{i}-u^{*}_{j})\}^{2}}{f(x;u^{*}_{i}-u^{*}_{j})}\operatorname{d}\!{x}\right]^{-1}.$ (4)
+$\beta_{n}=\max\left\{\frac{c_{n,2}^{2}c_{n,4}^{5/2}c_{n,5}}{c_{n,3}^{5}},\frac{c_{n,2}c_{n,4}^{11/2}}{c_{n,3}^{6}}\right\}\left\{\frac{q_{n}^{10}(\log n)^{8}}{np_{n}^{11}}\right\}^{1/2}\to 0\qquad n\to\infty,$ (3.3)
+then for each $i\in[n]$, $\{\rho_{i}({\bm{u}}^{*})\}^{-1/2}(\widehat{u}_{i}-u_{i}^{*})\to\mathcal{N}(0,1)$, where $\mathcal{N}(0,1)$ is the standard normal distribution and the asymptotic variance $\rho_{i}({\bm{u}}^{*})$ is given by
+$\rho_{i}({\bm{u}}^{*})$ $=\left[\sum_{j\in\delta\{i\}}\int_{{\mathbb{A}}}\frac{\{\partial_{2}f(x;u^{*}_{i}-u^{*}_{j})\}^{2}}{f(x;u^{*}_{i}-u^{*}_{j})}\operatorname{d}\!{x}\right]^{-1}.$ (3.4)
 Moreover, for any finite $S=\{i_{1},\ldots,i_{s}\}\subset{\mathbb{N}}$ where $S$ is independent of $n$, $\{\rho_{i_{1}}({\bm{u}}^{*})\}^{-1/2}(\widehat{u}_{i_{1}}-u_{i_{1}}^{*}),\ldots,\{\rho_{i_{s}}({\bm{u}}^{*})\}^{-1/2}(\widehat{u}_{i_{s}}-u_{i_{s}}^{*})$ are asymptotically independent.
-Remark 9
+Remark 2.
 Subject to additional conditions, such as a uniform upper bound on the number of edges between two subjects that is independent of $\bm{u}$$^{*}$, Theorem 1 still holds when two subjects are compared multiple times.
-Under such circumstances, the asymptotic variance formula (4) remains valid if one counts multiple edges for each adjacent vertex.
+Under such circumstances, the asymptotic variance formula (3.4) remains valid if one counts multiple edges for each adjacent vertex.
 A proof sketch for Theorem 1 is provided in Section 6.
 To better understand Theorem 1, note that although the multiplicative factor involving $c_{n,i}$ may seem daunting, it is $O(1)$ under additional regularity and boundedness conditions.
-Corollary 10
+Corollary 1.
 Assume that $M^{*}=\sup_{n}M_{n}<\infty$ and $\mathbb{A}$} is finite.
@@
 In general, Corollary 1 allows $p_{n}$ and $q_{n}$ to have different asymptotic orders.
-For instance, if $p_{n}=n^{-b}$ for $10\leq b<$ 1, then choosing $q_{n}=n^{-b^{\prime}}$ for any $b^{\prime}\geq 0$ with $b-(1-b)/10<b^{\prime}\leq b$ satisfies the required conditions.
+For instance, if $p_{n}=n^{-b}$ for $0\leq b<1$, then choosing $q_{n}=n^{-b^{\prime}}$ for any $b^{\prime}\geq 0$ with $b-(1-b)/10<b^{\prime}\leq b$ satisfies the required conditions.
 To the best of our knowledge, this is the first result that works for general comparison models with imbalanced data.
-Theorem 8 provides the limit distribution of the MLE, which lays the foundation for further statistical inference, such as the construction of interval-based estimates, testing hypothesis on the difference of merits, and testing the hypothesis on ranking position (Fan et al., 2024).
+Theorem 1 provides the limiting distribution of the MLE, which lays the foundation for further statistical inference, such as the construction of interval-based estimates, testing hypothesis on the difference of merits, and testing the hypothesis on ranking position (doi:10.1080/01621459.2024.2316364).
 Additionally, we can derive a non-asymptotic error rate for each parameter as well as a practical construction of interval-valued estimators, which are summarized in Theorem 2 and Theorem 3, respectively.
-Theorem 11 (Individual estimation error)
+Theorem 2 (Individual estimation error).
 Let ${\mathcal{G}}\sim{\mathsf{G}}(n,p_{n},q_{n})$.
-If both Assumptions 47 and (3) hold, then there exists an absolute constant $C>0$ such that for each $i\in[n]$, the following holds for all sufficiently large $n$ with probability tending to one:
-$|\widehat{u}_{i}-u_{i}^{*}|\leq C\cdot\frac{c_{n,2}}{c_{n,3}}\left(\frac{\log n}{|\delta\{i\}|}\right)^{1/2}.$ (5)
-We call the result in (5) the individual estimation error bound.
-Compared with the existing works (Han et al., 2020; Chen et al., 2022), this result presents new findings on the individual estimation error.
-Specifically, previous works considered the overall estimation error $\max_{i\in[n]}|\widehat{u}_{i}-u_{i}^{*}|$, while our result in (5) further illustrates the error for each individual.
+If both Assumptions 14 and (3.3) hold, then there exists an absolute constant $C>0$ such that for each $i\in[n]$, the following holds for all sufficiently large $n$ with probability tending to one:
+$|\widehat{u}_{i}-u_{i}^{*}|\leq C\cdot\frac{c_{n,2}}{c_{n,3}}\left(\frac{\log n}{|\delta\{i\}|}\right)^{1/2}.$ (3.5)
+We call the result in (3.5) the individual estimation error bound.
+Compared with the existing works (han2020asymptotic; chen2022optimal), this result presents new findings on the individual estimation error.
+Specifically, previous works considered the overall estimation error $\max_{i\in[n]}|\widehat{u}_{i}-u_{i}^{*}|$, while our result in (3.5) further illustrates the error for each individual.
 As expected, this error depends on the degree $|\delta\{i\}|$, which serves as a local normalization factor for each individual.
-When the comparison graph is homogeneous, individual estimation errors converge at the same rate, matching the results in Han et al.
-(2020); Chen et al.
-(2022).
+When the comparison graph is homogeneous, individual estimation errors converge at the same rate, matching the results in han2020asymptotic and chen2022optimal.
 For heterogeneous graphs, however, individuals will have different convergence rates, which cannot be reflected in the worst-case error $\max_{i\in[n]}|\widehat{u}_{i}-u_{i}^{*}|$.
 Next, we construct the confidence interval estimators. Since the asymptotic variance $\rho_{i}({\bm{u}}^{*})$ is unknown, it is natural to consider the following plug-in variance estimator:
-$\rho_{i}({\widehat{\bm{u}}})$ $=\left[\sum_{j\in\delta\{i\}}\int_{{\mathbb{A}}}\frac{\{\partial_{2}f(x;\widehat{u}_{i}-\widehat{u}_{i})\}^{2}}{f(x;\widehat{u}_{i}-\widehat{u}_{i})}\operatorname{d}\!{x}\right]^{-1}.$ (6)
+[…]
 Then for any fixed $i\in[n]$ and $\alpha\in(0,1)$, the $(1-\alpha)$-confidence interval of $u_{i}^{*}$ is
@@
 In our case, the increasing dimension of the parameter vector makes such a result less obvious, though it remains true, as justified by the following theorem.
-Theorem 12 (Confidence interval)
-Let ${\mathcal{G}}\sim{\mathsf{G}}(n,p_{n},q_{n})$. Under Assumptions 47, if (3) holds, then for any fixed $i\in[n]$ and $\alpha\in(0,1)$,
+Theorem 3 (Confidence interval).
+Let ${\mathcal{G}}\sim{\mathsf{G}}(n,p_{n},q_{n})$. Under Assumptions 14, if (3.3) holds, then for any fixed $i\in[n]$ and $\alpha\in(0,1)$,
 $\mathbb{P}\left(u^{*}_{i}\in\mathcal{CI}_{i}\right)\to 1-\alpha\ \text{ as }n\to\infty.$
-The computation of $\mathcal{CI}_{i}$ is straightforward after obtaining ${\widehat{\bm{u}}}$.
+The computation of $\mathcal{CI}_{i}$ is straightforward after obtaining $^{\widehat{\bm{u}}}$.
 For example, if one wants to construct a $95\%$-confidence interval for the $i$th subject, one can take $\mathcal{CI}_{i}=\big(\widehat{u}_{i}-1.96\{\rho_{i}({\widehat{\bm{u}}})\}^{-1/2},\widehat{u}_{i}+1.96\{\rho_{i}({\widehat{\bm{u}}})\}^{-1/2}\big)$.
@@
 For conveniences, we let $\Delta_{ij}=u^{*}_{i}-u^{*}_{j}$ for $i,j\in[n]$.
-Example 1 (BT model)
+Example 1 (BT model).
 The BT model assumes a binary comparison outcome between two subjects and parameterizes using a logistic link function:
 $f(1;y)=\frac{e^{y}}{1+e^{y}};\quad f(-1;y)=\frac{1}{1+e^{y}}.$
-It can be verified that both conditions in (2) and (3) are satisfied if
+It can be verified that both conditions in (3.2) and (3.3) are satisfied if
 $e^{6M_{n}}\left\{\frac{q_{n}^{10}(\log n)^{8}}{np_{n}^{11}}\right\}^{1/2}\to 0\qquad n\to\infty.$
@@
 $\rho_{i}({\bm{u}}^{*})=\left\{\sum_{j\in\delta\{i\}}\frac{e^{\Delta_{ij}}}{\left(1+e^{\Delta_{ij}}\right)^{2}}\right\}^{-1}.$
-That is, $\{{\rho_{i}({\bm{u}}^{*})}\}^{-1/2}(\widehat{u}_{i}-u^{*}_{i})\to N(0,1)$ as $n\to\infty$.
-Example 2 (Thurstone–Mosteller model)
+That is, $\{{\rho_{i}({\bm{u}}^{*})}\}^{-1/2}(\widehat{u}_{i}-u^{*}_{i})\to\mathcal{N}(0,1)$ as $n\to\infty$.
+Example 2 (Thurstone–Mosteller model).
 The Thurstone–Mosteller model considers an alternative link function than the BT model:
 $f(1;y)=\Phi(y);\quad f(-1;y)=1-\Phi(y),$
-where $\Phi(y)=\int_{-\infty}^{y}\varphi(z)\operatorname{d}\!{z}$ with $\varphi(z)=(2\pi)^{-1/2}e^{-z^{2}/2}$. It can be verified that both conditions in (2) and (3) are satisfied if
+where $\Phi(y)=\int_{-\infty}^{y}\varphi(z)\operatorname{d}\!{z}$ with $\varphi(z)=(2\pi)^{-1/2}e^{-z^{2}/2}$. It can be verified that both conditions in (3.2) and (3.3) are satisfied if
 $M_{n}^{6}e^{3M^{2}_{n}}\left\{\frac{q_{n}^{10}(\log n)^{8}}{np_{n}^{11}}\right\}^{1/2}\to 0\qquad n\to\infty.$
@@
 $\rho_{i}({\bm{u}}^{*})=\left\{\sum_{j\in\delta\{i\}}\frac{\varphi^{2}(\Delta_{ij})}{\Phi(\Delta_{ij})(1-\Phi(\Delta_{ij}))}\right\}^{-1}.$
-That is, $\{{\rho_{i}({\bm{u}}^{*})}\}^{-1/2}(\widehat{u}_{i}-u^{*}_{i})\to N(0,1)$ as $n\to\infty$.
-Example 3 (Rao–Kupper model)
+That is, $\{{\rho_{i}({\bm{u}}^{*})}\}^{-1/2}(\widehat{u}_{i}-u^{*}_{i})\to\mathcal{N}(0,1)$ as $n\to\infty$.
+Remark 3.
+Both the BT and Thurstone–Mosteller model models can be viewed as special cases of the Random Utility Model framework, which was formally developed for discrete choice analysis by mcfadden1973conditional.
+Specifically, the BT model arises from the assumption of i.i.d.
+Gumbel-distributed random utilities, while the Thurstone–Mosteller model assumes they are normally distributed.
+Our central findings can be generalized to the other types random utility models, which satisfies (A1)–(A5) in Definition 1.
+Example 3 (Rao–Kupper model).
 The Rao–Kupper model extends the BT model by incorporating ties. The link function is given by
 $f(1;y)$ $=\frac{e^{y}}{e^{y}+\theta};\quad f(0;y)=\frac{\left(\theta^{2}-1\right)e^{y}}{\left(e^{y}+\theta\right)\left(\theta e^{y}+1\right)};\quad f(-1;y)=\frac{1}{\theta e^{y}+1},$
-where $\theta>1$ is the threshold parameter which is predetermined. It can be verified that both conditions in (2) and (3) are satisfied if
+where $\theta>1$ is the threshold parameter which is predetermined. It can be verified that both conditions in (3.2) and (3.3) are satisfied if
 $e^{6M_{n}}\left\{\frac{q_{n}^{10}(\log n)^{8}}{np_{n}^{11}}\right\}^{1/2}\to 0\qquad n\to\infty.$
@@
 $\rho_{i}({\bm{u}}^{*})=\left\{\sum_{j\in\delta\{i\}}\left[\frac{\theta^{2}e^{\Delta_{ij}}}{(\theta+e^{\Delta_{ij}})^{3}}+\frac{\theta^{2}(\theta^{2}-1)e^{\Delta_{ij}}(1-e^{2\Delta_{ij}})^{2}}{(e^{\Delta_{ij}}+\theta)^{3}(\theta e^{\Delta_{ij}}+1)^{3}}+\frac{\theta^{2}e^{2\Delta_{ij}}}{(\theta e^{\Delta_{ij}}+1)^{3}}\right]\right\}^{-1}.$
-That is, $\{{\rho_{i}({\bm{u}}^{*})}\}^{-1/2}(\widehat{u}_{i}-u^{*}_{i})\to N(0,1)$ as $n\to\infty$.
-Example 4 (Davidson model)
+That is, $\{{\rho_{i}({\bm{u}}^{*})}\}^{-1/2}(\widehat{u}_{i}-u^{*}_{i})\to\mathcal{N}(0,1)$ as $n\to\infty$.
+Example 4 (Davidson model).
 As opposed to the Rao–Kupper model, the Davidson model considers an alternative parameterization of being tied:
 $f(1;y)=\frac{e^{y}}{e^{y}+\theta e^{\frac{y}{2}}+1};\quad f(0;y)=\frac{\theta e^{\frac{y}{2}}}{e^{y}+\theta e^{\frac{y}{2}}+1};\quad f(-1;y)$ $=\frac{1}{e^{y}+\theta e^{\frac{y}{2}}+1},$
-where $\theta>0$ is assumed to be prefixed. It can be verified that both conditions in (2) and (3) are satisfied if
+where $\theta>0$ is assumed to be prefixed. It can be verified that both conditions in (3.2) and (3.3) are satisfied if
 $e^{{3M_{n}}}\left\{\frac{q_{n}^{10}(\log n)^{8}}{np_{n}^{11}}\right\}^{1/2}\to 0\qquad n\to\infty.$
@@
 $\rho_{i}({\bm{u}}^{*})=\left\{\sum_{j\in\delta\{i\}}\frac{e^{\Delta_{ij}}(\theta e^{\Delta_{ij}/2}+2)^{2}+\theta e^{\Delta_{ij}/2}(1-e^{\Delta_{ij}})^{2}+(2e^{\Delta_{ij}}+\theta e^{\Delta_{ij}/2})^{2}}{4(e^{\Delta_{ij}}+\theta e^{\Delta_{ij}/2}+1)^{3}}\right\}^{-1}.$
-That is, $\{{\rho_{i}({\bm{u}}^{*})}\}^{-1/2}(\widehat{u}_{i}-u^{*}_{i})\to N(0,1)$ as $n\to\infty$.
-Example 5 (Cumulative link model with four outcomes)
-The cumulative link model is defined via the ordinalization of some distribution function $F$ (Agresti, 1992).
+That is, $\{{\rho_{i}({\bm{u}}^{*})}\}^{-1/2}(\widehat{u}_{i}-u^{*}_{i})\to\mathcal{N}(0,1)$ as $n\to\infty$.
+Example 5 (Cumulative link model with four outcomes).
+The cumulative link model is defined via the ordinalization of some distribution function $F$ (agresti1992analysis).
 The BT model, the Rao–Kupper model, and the Thurstone–Mosteller model can all be considered as specific cases of the cumulative link model with appropriate choices of $F$.
@@
 In this case, $\mathbb{A}=\{-2,-1,1,2\}$ and the corresponding link function takes the form of
-$f(1;y)=\frac{(\theta-1)e^{y}}{\left(\theta+e^{y}\right)\left(1+e^{y}\right)},\ \ \ f(2;y)=\frac{e^{y}}{\theta+e^{y}},$ (7)
-where $\theta>1$ is the threshold parameter that is predetermined. It can be verified that both conditions in (2) and (3) are satisfied if
+$f(1;y)=\frac{(\theta-1)e^{y}}{\left(\theta+e^{y}\right)\left(1+e^{y}\right)},\ \ \ f(2;y)=\frac{e^{y}}{\theta+e^{y}},$ (4.7)
+where $\theta>1$ is the threshold parameter that is predetermined. It can be verified that both conditions in (3.2) and (3.3) are satisfied if
 $e^{6M_{n}}\left\{\frac{q_{n}^{10}(\log n)^{8}}{np_{n}^{11}}\right\}^{1/2}\to 0\qquad n\to\infty.$
@@
 ${\rho_{i}({\bm{u}}^{*})=\left\{\sum_{j\in\delta\{i\}}\left[\frac{\theta^{2}e^{\Delta_{ij}}}{(\theta+e^{\Delta_{ij}})^{3}}+\frac{\theta^{2}e^{2\Delta_{ij}}}{(\theta e^{\Delta_{ij}}+1)^{3}}+\frac{(\theta-1)e^{\Delta_{ij}}}{(1+e^{\Delta_{ij}})^{3}}\left(\frac{(\theta-e^{\Delta_{ij}})^{2}}{(\theta+e^{\Delta_{ij}})^{3}}+\frac{1-\theta e^{2\Delta_{ij}}}{(\theta e^{\Delta_{ij}}+1)^{3}}\right)\right]\right\}^{-1}.}$
-That is, $\{{\rho_{i}({\bm{u}}^{*})}\}^{-1/2}(\widehat{u}_{i}-u^{*}_{i})\to N(0,1)$ as $n\to\infty$.
+That is, $\{{\rho_{i}({\bm{u}}^{*})}\}^{-1/2}(\widehat{u}_{i}-u^{*}_{i})\to\mathcal{N}(0,1)$ as $n\to\infty$.
 The result above can be extended to the cumulative link model with the logistic link function and finite outcomes but involve more complicated formulas for asymptotic variances.
-Example 6 (Paired cardinal model)
+Example 6 (Paired cardinal model).
 The paired cardinal model can be viewed as a fully observed version of the Thurstone–Mosteller model, where the comparison outcome takes values in $\mathbb{A}={\mathbb{R}}$, and the corresponding link function is parameterized using the normal density:
 $f(x;y)=\left(2\pi\sigma^{2}\right)^{-\frac{1}{2}}e^{-\frac{(x-y)^{2}}{2\sigma^{2}}},$
-where $\sigma>0$ is some fixed parameter. It can be verified that both conditions in (2) and (3) are satisfied if
+where $\sigma>0$ is assumed to be prefixed. It can be verified that both conditions in (3.2) and (3.3) are satisfied if
 $\max\left\{M_{n}e^{\frac{M_{n}^{2}}{2\sigma^{2}}}\frac{\log n}{np_{n}},\left({\frac{q_{n}^{10}(\log n)^{8}}{np_{n}^{11}}}\right)^{1/2}\right\}\to 0\qquad n\to\infty.$
@@
 $\rho_{i}({\bm{u}}^{*})=\frac{\sigma^{2}}{|\delta\{i\}|},$
-which is independent of ${\bm{u}}^{*}$. That is, $\{{\rho_{i}({\bm{u}}^{*})}\}^{-1/2}(\widehat{u}_{i}-u^{*}_{i})\to N(0,1)$ as $n\to\infty$.
+which is independent of $\bm{u}$$^{*}$. That is, $\{{\rho_{i}({\bm{u}}^{*})}\}^{-1/2}(\widehat{u}_{i}-u^{*}_{i})\to\mathcal{N}(0,1)$ as $n\to\infty$.
 5 Numerical studies
@@
 The total number of subjects $n$ in all the tested models is chosen from the set $\{500,1000,2000\}$.
-To generate the utility vector ${\bm{u}}^{*}$, we uniformly select values from the range $[-M_{n},M_{n}]$, where the dynamic range $M_{n}$ is chosen from $\{1,\log(\log n)\}$.
+To generate the utility vector $\bm{u}$$^{*}$, we uniformly select values from the range $M_{n},M_{n}]$, where the dynamic range $M_{n}$ is chosen from $\{1,\log(\log n)\}$.
 For global parameters in each model, we set $\theta=1$ in the Davidson model, $\theta=2$ in the Rao–Kupper model, and $\sigma=2$ in the paired cardinal model.
@@
 We report the average standard deviation of the estimated parameters and the coverage probability of the $95\%$ confidence interval, as shown in Table 2.
-We also examine the $z$-scores of the first coordinate of ${{\widehat{\bm{u}}}}$ over 300 experiments and plot the corresponding quantile-quantile plots against normal quantiles in Figure 1.
-In Table 2, we observe a decreasing trend in the average standard deviation as $n$ increases, which is consistent with our theoretical findings on asymptotic variance in Equation (4).
+We also examine the $z$-scores of the first coordinate of $^{{\widehat{\bm{u}}}}$ over 300 experiments and plot the corresponding quantile-quantile plots against normal quantiles in Figure 1.
+In Table 2, we observe a decreasing trend in the average standard deviation as $n$ increases, which is consistent with our theoretical findings on asymptotic variance in (3.4).
 Additionally, we note variations in standard deviation relative to $M_{n}$ across different models.
@@
 Simulation parameters: For each $n$, the empirical quantile curves are based on 300 experiments.
-5.2 Real data analysis
-[…]
+5.2 Real data example
+In the realm of men’s tennis, the dominance of Novak Djokovic, Roger Federer, Rafael Nadal, and Andy Murray, collectively known as the Big Four, has undeniably dominated the first quarter of the 21st century.
+However, recent discussions have emerged, advocating for the exclusion of Andy Murray in what could potentially become the Big Three.
+In this section, we consider a statistical approach to model the ATP data and uncover the latent scores that provide evidence in favour of the Big Four.
+We consider modeling the comparison data from ATP111www.tennis-data.co.uk using the cumulative link model with four outcomes from Example 5.
+The ATP dataset collects the comparison results of each game from 1968 to 2016 by 2742 players.
+In this study, we concentrate on the best-of-3 matches, which result in four possible outcomes $\{\text{0:2, 1:2, 2:1, 2:0}\}$, represented as ${\mathbb{A}}=\{-2,-1,1,2\}.$ In particular, transferring the ordinal outcomes (4.7) into the current setting reads
 $f(X_{ij}=1,u_{i}-u_{j})=\frac{(\theta-1)e^{u_{i}-u_{j}}}{\left(\theta+e^{u_{i}-u_{j}}\right)\left(1+e^{u_{i}-u_{j}}\right)},$
 $f(X_{ij}=2,u_{i}-u_{j})=\frac{e^{u_{i}-u_{j}}}{\theta+e^{u_{i}-u_{j}}},\ \ \ \theta>1.$
-To identify whether a significant distinction exists between the Big Three in tennis and Andy Murray, we employed a multiple testing procedure on the difference in their latent scores.
-Specifically, let $u_{1},u_{2},u_{3}$ and $u_{4}$ denote the latent score of Djokovic, Nadal, Federer, and Murray respectively.
-We consider the hypothesis testing with the null hypothesis $H_{0}:u_{i}=u_{4},i\in[3]$ and alternative hypothesis $H_{A}:$ there exists $i\in[3]$ such that $u_{i}\neq u_{4}$.
-The significance level $\alpha=0.05.$ We first obtained the maximum likelihood estimates of the latent scores $\widehat{{\bm{u}}}$, and the threshold parameter $\widehat{\theta}=2.32$.
-The estimator $\widehat{u}_{i}$ and estimated standard deviation $(\widehat{\rho}_{i})^{1/2}$ of the 4 candidates are presented in Table 3.
-We construct the $z$-statistics $(\widehat{u}_{i}-\widehat{u}_{4})/{{(\widehat{\rho}_{i}+\widehat{\rho}_{4})^{1/2}}},\ i\in[3]$, which yields $p$-values $0.229$, $0.226$ and $0.334$ accordingly.
-For multiple testing, we adopt the Benjamini–Hochberg procedure (Benjamini and Hochberg, 1995).
-We order three $p$-values in ascending order and compare them with the corresponding critical values, that is $(\alpha)/3$, $(2\alpha)/3$, and $\alpha$.
-It turns out that all $p$-values exceed the corresponding critical value, implying that the difference in latent scores between Big Three and Murray is not statistically significant.
-Consequently, there is no sufficient statistical evidence under the selected threshold to reject the null hypothesis which supports the Big Four.
+The estimated latent scores $\widehat{u}_{i}$ and corresponding standard deviation $(\widehat{\rho}_{i})^{1/2}$ of the 4 candidates are presented in Table 3, and it shows that Djokovic, Nadal, and Federer exhibit comparable performance levels, while Murray’s score is slightly lower.
+Moreover, the overlap in standard deviations implies potential ranking uncertainty.
+To quantify the ranking uncertainty, we use parametric bootstrap to simulate 10,000 samples by the asymptotic normal distributions $\mathcal{N}(\widehat{u}_{i},\widehat{\rho}_{i})$.
+Then we calculate empirical ranks across Monte Carlo samples and construct histogram for each rank position, shown in Figure 2(a), where each bar indicates the probability of each player ranking from 1st through 4th.
+Figure 2(a) reveals that Djokovic has the highest probability of achieving the top rank, closely followed by Nadal and Federer, while Murray consistently ranks last, indicating the gap between him and the Big Three.
+To further investigate pairwise competitiveness among players, we construct a posterior probability matrix.
+It summarizes the likelihood that player $i$ outperforms player $j$ based on their estimates.
+Specifically, we compute:
+${\mathbb{P}}\left(u_{i}>u_{j}\right)=\mathbb{P}\left(Z_{i}-Z_{j}>0\right),$
+where $Z_{i}\sim\mathcal{N}\left(\widehat{u}_{i},\widehat{\rho}_{i}\right)$ and $Z_{j}\sim\mathcal{N}\left(\widehat{u}_{j},\widehat{\rho}_{j}\right)$ are independently drawn from the posterior distributions of players $i$ and $j$, respectively.
+The asymptotic distribution guarantees that
+$\mathbb{P}\left(u_{i}>u_{j}\right)=\Phi\left(\frac{\widehat{u}_{i}-\widehat{u}_{j}}{\sqrt{\widehat{\rho}_{i}+\widehat{\rho}_{j}}}\right),$
+where $\Phi(\cdot)$ denotes the standard normal cumulative distribution function.
+The heatmap in Figure 2(b) displays the resulting posterior win probability matrix , where each entry $(i,j)$ shows the empirical probability that player $i$ wins player $j$ in terms of latent score under 10,000 repetitions.
+We can observe that Djokovic and Nadal have approximately an equal chance of outperforming each other, and both have $60$-$65\%$ probability of outperforming Federer.
+Murray has the lowest win probabilities against all other players, consistent with his lower latent score.
+In general, Djokovic and Nadal are the top players with extremely close latent strength, Federer is a strong challenger but slightly behind and Murray most frequently ranked the lowest.
+To further assess the uncertainty in individual scores, we construct confidence intervals for latent scores.
+Figure 3(a) presents the $95\%$ confidence intervals for $\widehat{u}_{i}$ of each player under the CLM4 model.
+We use standard normal approximation to the asymptotic distribution to construct the confidence intervals, i.e., $\widehat{u}_{i}\pm z_{0.95}\sqrt{\widehat{\rho}_{i}}$.
+Figure 3(a) confirms the dominance of Djokovic and Nadal, the competitive strength of Federer, and a notable performance gap for Murray.
+However, the overlapping of all confidence intervals shows that these 4 players are not significantly different.
+To further validate the distinction with statistically significance, we construct the $95\%$ confidence interval for latent score difference.
+To construct the intervals, we apply the Bonferroni-adjusted critical value to the pairwise normal differences
+$\widehat{u}_{i}-\widehat{u}_{j}\pm z_{1-\alpha/(2K)}\cdot\sqrt{\widehat{\rho}_{i}+\widehat{\rho}_{j}}$
+where $K$ is the number of pairwise comparisons and $z_{1-\alpha/(2K)}$ is the Bonferroni-corrected quantile.
+Figure 3(b) shows that none of the intervals exclude point zero, indicating that although Murray seems rank the least from previous experiments, all four players are not significantly different from others at the $95\%$ confidence level.
+In summary, both point estimates and posterior analyses indicate a discernible ranking between Murray and the Big Three.
+However, the simultaneous confidence intervals reveal that none of the pairwise differences are statistically significant.
+This highlights the critical role of formal statistical inference in practical applications.
 Table 3: The estimated scores and their standard deviation of four players by using CLM4.
-Players Estimated scores $\widehat{{\bm{u}}}_{i}$ Estimated standard deviation $(\widehat{\rho}_{i})^{1/2}$
+Players Estimated scores $\widehat{u}_{i}$ Estimated standard deviation $(\widehat{\rho}_{i})^{1/2}$
 Novak Djokovic 3.235 $0.229$
@@
 Andy Murray 2.872 $0.196$
+(a) Posterior ranking probability histogram.
+(b) Win probability heatmap.
+Figure 2: The posterior ranking and pairwise win probability for each player.
+(a) The $95\%$ confidence interval of latent score.
+(b) The $95\%$ confidence interval of difference.
+Figure 3: The $95\%$ confidence interval of the latent score of individual players and the pairwise difference between players’ scores.
 6 Proof of Theorem 1
@@
 However, the randomness of the Hessian matrix introduces additional challenges.
-The novelty of our approach lies in identifying the Hessian matrix as the weighted Laplacian matrix of appropriate random graphs, which can be carefully analyzed using truncated spectral methods.
+The crux of our approach lies in identifying the corresponding Fisher information matrix as the weighted Laplacian matrix of appropriate random graphs, which can be carefully analyzed using truncated spectral methods.
+Although leveraging the Fisher information matrix to derive asymptotic normality in network models is well established (MR1724040; yan2013central; han2023unified), existing results require specific model structures, particularly those with logistic-type link functions.
+To the best of our knowledge, our work provides the first general framework that maintains optimality conditions while accommodating broader model specifications.
 We start by computing the first- and second-order derivatives of the log-likelihood function $l$ for $\bm{u}$}. For $i,j\in[n]$ with $i\neq j$,
@@
 $\partial_{ii}l({\bm{u}})$ $=-\sum_{j\in\delta\{i\}}\partial_{ij}l({\bm{u}})\leq 0,$
-where the first inequality follows from the log-concavity of $f$ that is assumed under the validity assumption. Then the Hession of $l({\bm{u}})$, denoted by $H({\bm{u}})$ is given as
-$\{H({\bm{u}})\}_{ij}$ =$=\left\{\begin{array}[]{ll}\partial_{ij}l({\bm{u}})&\quad i\neq j\\ -\sum_{k\in\delta\{i\}}\partial_{ik}l({\bm{u}})&\quad i=j.\end{array}\right.$
+where the first inequality follows from the log-concavity of $f$ that is assumed under the validity assumption. Then the Hessian of $l({\bm{u}})$, denoted by $H({\bm{u}})$ is given as
+$\{H({\bm{u}})\}_{ij}$ $=\left\{\begin{array}[]{ll}\partial_{ij}l({\bm{u}})&\quad i\neq j\\ -\sum_{k\in\delta\{i\}}\partial_{ik}l({\bm{u}})&\quad i=j.\end{array}\right$.
 A crucial observation is that $H({\bm{u}})$ is the negation of an unnormalized graph Laplacian associated with a weighted graph on $[n]$ with edge weights $\partial_{ij}l({\bm{u}})$.
@@
 As such, we let $H^{*}({\bm{u}})={\mathbb{E}}[H({\bm{u}})\mid{\mathcal{E}}]$.
-It is easy to check that $-H^{*}({\bm{u}})$ remains an unnormalized weighted graph Laplacian that depends only on ${\mathcal{G}}({\mathcal{V}},{\mathcal{E}})$ and ${\bm{u}}$.
-In the following, we normalize $-H^{*}({\bm{u}}^{*})$ as
-$-H^{*}({\bm{u}}^{*})={\mathcal{D}}^{1/2}{\mathcal{L}}_{sym}{\mathcal{D}}^{1/2}={\mathcal{D}}^{1/2}(I-{\mathcal{A}}){\mathcal{D}}^{1/2},$ (8)
-where ${\mathcal{D}}=\text{diag}\{-H^{*}({\bm{u}}^{*})\}$ is the diagonal of $-H^{*}({\bm{u}}^{*})$ (degree matrix), ${\mathcal{L}}_{sym}$ is the normalized Laplacian of the weighted graph associated with $-H^{*}({\bm{u}}^{*})$, ${\mathcal{A}}=I-{\mathcal{L}}_{sym}$ is the normalized adjacency matrix, and $I$ is the identity matrix.
+It is easy to check that the Fisher information matrix, $H^{*}({\bm{u}})$ remains an unnormalized weighted graph Laplacian that depends only on ${\mathcal{G}}({\mathcal{V}},{\mathcal{E}})$ and $\bm{u}$}.
+In the following, we normalize $H^{*}({\bm{u}}^{*})$ as
+$-H^{*}({\bm{u}}^{*})={\mathcal{D}}^{1/2}{\mathcal{L}}_{sym}{\mathcal{D}}^{1/2}={\mathcal{D}}^{1/2}(I-{\mathcal{A}}){\mathcal{D}}^{1/2},$ (6.8)
+where ${\mathcal{D}}=\text{diag}\{-H^{*}({\bm{u}}^{*})\}$ is the diagonal of $H^{*}({\bm{u}}^{*})$ (degree matrix), ${\mathcal{L}}_{sym}$ is the normalized Laplacian of the weighted graph associated with $H^{*}({\bm{u}}^{*})$, ${\mathcal{A}}=I-{\mathcal{L}}_{sym}$ is the normalized adjacency matrix, and $I$ is the identity matrix.
 By the mean value theorem,
@@
 $\bar{H}({\bm{W}})=\begin{pmatrix}\partial_{11}l({\bm{w}}_{1}),\ldots,\partial_{1n}l({\bm{w}}_{1})\\ \\ \partial_{n1}l({\bm{w}}_{n}),\ldots,\partial_{nn}l({\bm{w}}_{n})\end{pmatrix}\in{\mathbb{R}}^{n\times n},$
-with ${\bm{W}}=({\bm{w}}_{1},\ldots,{\bm{w}}_{n}),$ and ${\bm{w}}_{1},\ldots,{\bm{w}}_{n}$ are $n$ (possibly distinct) points on the line segment between ${\widehat{\bm{u}}}$ and ${\bm{u}}^{*}$.
+with ${\bm{W}}=({\bm{w}}_{1},\ldots,{\bm{w}}_{n}),$ and ${\bm{w}}_{1},\ldots,{\bm{w}}_{n}$ are $n$ (possibly distinct) points on the line segment between $^{\widehat{\bm{u}}}$ and $\bm{u}$$^{*}$.
 It is worth noting that $\bar{H}({\bm{W}})$ is not a Hessian matrix (it is not even symmetric).
-Heuristically, thanks to the matrix concentration and uniform consistency of ${\widehat{\bm{u}}}$, $H^{*}({\bm{u}}^{*})-H({\bm{u}}^{*})=o(H({\bm{u}}^{*}))$ and $H({\bm{u}}^{*})-\bar{H}({\bm{W}})=o(H({\bm{u}}^{*}))$, where the $o(\cdot)$ is interpreted in an appropriate sense that will be made clear in Lemma 13.
-With some rearrangement and substituting (8) into $H^{*}({\bm{u}}^{*})$, we have
-${\mathcal{D}}^{1/2}({\widehat{\bm{u}}}-{\bm{u}}^{*})$ $=-{\mathcal{L}}^{\dagger}_{sym}{\mathcal{D}}^{-1/2}\nabla l({\bm{u}}^{*})$ (9)
+Heuristically, thanks to the matrix concentration and uniform consistency of $^{\widehat{\bm{u}}}$, $H^{*}({\bm{u}}^{*})-H({\bm{u}}^{*})=o(H({\bm{u}}^{*}))$ and $H({\bm{u}}^{*})-\bar{H}({\bm{W}})=o(H({\bm{u}}^{*}))$, where the $o(\cdot)$ is interpreted in an appropriate sense that will be made clear in Lemma 1.
+With some rearrangement and substituting (6.8) into $H^{*}({\bm{u}}^{*})$, we have
+${\mathcal{D}}^{1/2}({\widehat{\bm{u}}}-{\bm{u}}^{*})$ $=-{\mathcal{L}}^{\dagger}_{sym}{\mathcal{D}}^{-1/2}\nabla l({\bm{u}}^{*})$ (6.9)
 $\ \ \ \ \ +{\mathcal{L}}^{\dagger}_{sym}{\mathcal{D}}^{-1/2}\{H^{*}({\bm{u}}^{*})-H({\bm{u}}^{*})\}({\widehat{\bm{u}}}-{\bm{u}}^{*})$
@@
 where ${\mathcal{L}}_{sym}^{\dagger}$ is the pseudoinverse of ${\mathcal{L}}_{sym}$ and can be explicitly expressed using spectral expansion:
-${\mathcal{L}}_{sym}^{\dagger}=\sum_{t=0}^{\infty}({\mathcal{A}}-\mathcal{P}_{1})^{t}-\mathcal{P}_{1},$ (10)
+${\mathcal{L}}_{sym}^{\dagger}=\sum_{t=0}^{\infty}({\mathcal{A}}-\mathcal{P}_{1})^{t}-\mathcal{P}_{1},$ (6.10)
 where $\mathcal{P}_{1}$ is the orthogonal projector to the null space of $\mathcal{L}$}.
 Particularly, letting $H_{ij}=\{H^{*}({\bm{u}}^{*})\}_{ij}$, the entry of $\mathcal{A}$} and $\mathcal{P}_{1}$ can be written as
-${\mathcal{D}}_{ii}=\sum_{j\in\delta\{i\}}H_{ij},\quad{\mathcal{A}}_{ij}=\frac{H_{ij}}{{{\mathcal{D}}_{ii}^{1/2}{\mathcal{D}}_{jj}^{1/2}}},\quad\{\mathcal{P}_{1}\}_{ij}=\frac{{{\mathcal{D}}_{ii}^{1/2}{\mathcal{D}}_{jj}^{1/2}}}{\sum_{k\in[n]}{\mathcal{D}}_{kk}}.$ (11)
-Based on the decomposition in (C.2), we will proceed to prove the desired asymptotic normality result as follows.
-We will show that the first term on the right-hand side of (C.2) appropriately normalizes $({\widehat{\bm{u}}}-{\bm{u}}^{*})$ and gives the asymptotic normality.
+${\mathcal{D}}_{ii}=\sum_{j\in\delta\{i\}}H_{ij},\quad{\mathcal{A}}_{ij}=\frac{H_{ij}}{{{\mathcal{D}}_{ii}^{1/2}{\mathcal{D}}_{jj}^{1/2}}},\quad\{\mathcal{P}_{1}\}_{ij}=\frac{{{\mathcal{D}}_{ii}^{1/2}{\mathcal{D}}_{jj}^{1/2}}}{\sum_{k\in[n]}{\mathcal{D}}_{kk}}.$ (6.11)
+Based on the decomposition in (6.9), we will proceed to prove the desired asymptotic normality result as follows.
+We will show that the first term on the right-hand side of (6.9) appropriately normalizes $({\widehat{\bm{u}}}-{\bm{u}}^{*})$ and gives the asymptotic normality.
 Meanwhile, through a meticulous analysis (Lemma 1), it can be established that the remaining terms are of order $o_{p}(1)$.
 Consequently, the desired follows by applying Slutsky’s theorem.
-Lemma 13
+Lemma 1.
 Under the assumptions in Theorem 1, the following are true:
@@
 The proof of Lemma 1 is rather technical and deferred to the appendix.
-We now use the results in Lemma 13 to establish the asymptotic normality of ${\widehat{\bm{u}}}$.
-By statements (1) and (2) in Lemma 13 and Slutsky’s theorem, it remains to verify for every fixed $S\subset{\mathbb{N}}$ with $|S|=s$ ($S$ independent of $n$), $\{{\mathcal{L}}^{\dagger}_{sym}{\mathcal{D}}^{-1/2}\nabla l({\bm{u}}^{*})\}_{S}\to N(0,I_{s})$.
+We now use the results in Lemma 1 to establish the asymptotic normality of $^{\widehat{\bm{u}}}$.
+By statements (1) and (2) in Lemma 1 and Slutsky’s theorem, it remains to verify for every fixed $S\subset{\mathbb{N}}$ with $|S|=s$ ($S$ independent of $n$), $\{{\mathcal{L}}^{\dagger}_{sym}{\mathcal{D}}^{-1/2}\nabla l({\bm{u}}^{*})\}_{S}\to\mathcal{N}(0,I_{s})$.
 Since
-${\mathcal{L}}^{\dagger}_{sym}{\mathcal{D}}^{-1/2}\nabla l({\bm{u}}^{*})={\mathcal{D}}^{-1/2}\nabla l({\bm{u}}^{*})+({\mathcal{L}}^{\dagger}_{sym}-I){\mathcal{D}}^{-1/2}\nabla l({\bm{u}}^{*})$ (12)
-and $\{{\mathcal{D}}^{-1/2}\nabla l({\bm{u}}^{*})\}_{S}\to N(0,I_{s})$ by the central limit theorem, the proof is finished if $\{({\mathcal{L}}^{\dagger}_{sym}-I){\mathcal{D}}^{-1/2}\nabla l({\bm{u}}^{*})\}_{S}=o_{p}(1)$ componentwise.
+${\mathcal{L}}^{\dagger}_{sym}{\mathcal{D}}^{-1/2}\nabla l({\bm{u}}^{*})={\mathcal{D}}^{-1/2}\nabla l({\bm{u}}^{*})+({\mathcal{L}}^{\dagger}_{sym}-I){\mathcal{D}}^{-1/2}\nabla l({\bm{u}}^{*})$ (6.12)
+and $\{{\mathcal{D}}^{-1/2}\nabla l({\bm{u}}^{*})\}_{S}\to\mathcal{N}(0,I_{s})$ by the central limit theorem, the proof is finished if $\{({\mathcal{L}}^{\dagger}_{sym}-I){\mathcal{D}}^{-1/2}\nabla l({\bm{u}}^{*})\}_{S}=o_{p}(1)$ componentwise.
 To verify this, we apply Chebyshev’s inequality.
@@
 $=$ $\ ({\mathcal{L}}^{\dagger}_{sym}+\mathcal{P}_{1}-I)_{kk}+({\mathcal{L}}_{sym}+\mathcal{P}_{1}-I)_{kk}$
-${{=}}$ $\ \sum_{t=2}^{\infty}\{({\mathcal{A}}-\mathcal{P}_{1})^{t}\}_{kk}=o_{p}(1),$
+[…]
 where the second equality uses the fact that ${\mathbb{E}}[\nabla l({\bm{u}}^{*})\{\nabla l({\bm{u}}^{*})\}^{\top}]=-H^{*}({\bm{u}}^{*})$ and the last equality follows from statement (3) in Lemma 1.
@@
 Addressing these considerations may involve incorporating additional covariate information.
-We intend to pursue these directions in our future work.
-Appendices
-Appendix A Proof of Lemma 13
+In addition, one theoretical research direction is to apply our technique for establishing asymptotic normality, that is, the representation of the Fisher information matrix as a weighted graph Laplacian, to a broader class of network models.
+In particular, we plan to adapt it to the $\beta$-model (chatterjee2011random), which shares structural similarities with the BT model, and its generalized variants in econometric networks (graham2017econometric; chen2021analysis).
+These investigations will form part of our future work.
+Appendix A Appendix: Proof of Lemma 1
 We start by making two observations.
@@
 Let $\kappa_{n}=c_{n,4}/c_{n,3}\in[1,\infty)$.
-Lemma 14
-If $\lim_{n\to\infty}(\kappa^{7}_{n}q^{6}_{n}\log n)/(np^{7}_{n})=0$, then there exists an absolute constant $C$ such that $\|{\mathcal{A}}-\mathcal{P}_{1}\|_{2}\leq 1-C\{p_{n}/(\kappa_{n}q_{n})\}^{2}$.
-Lemma 15
+Lemma 2.
+If $\lim_{n\to\infty}(\kappa^{7}_{n}q^{6}_{n}\log n)/(np^{7}_{n})=0$, then there exists an absolute constant $C>0$ such that $\|{\mathcal{A}}-\mathcal{P}_{1}\|_{2}\leq 1-C\{p_{n}/(\kappa_{n}q_{n})\}^{2}$ holds with probability tending to one.
+Lemma 3.
 If Assumptions 14 hold, then
@@
 Recall that
-${\mathcal{D}}_{ii}=\sum_{j\in\delta\{i\}}H_{ij},\quad{\mathcal{A}}_{ij}=\frac{H_{ij}}{{{\mathcal{D}}_{ii}^{1/2}{\mathcal{D}}_{jj}^{1/2}}},\quad\{\mathcal{P}_{1}\}_{ij}=\frac{{{\mathcal{D}}_{ii}^{1/2}{\mathcal{D}}_{jj}^{1/2}}}{\sum_{k\in[n]}{\mathcal{D}}_{kk}}.$ (13)
+${\mathcal{D}}_{ii}=\sum_{j\in\delta\{i\}}H_{ij},\quad{\mathcal{A}}_{ij}=\frac{H_{ij}}{{{\mathcal{D}}_{ii}^{1/2}{\mathcal{D}}_{jj}^{1/2}}},\quad\{\mathcal{P}_{1}\}_{ij}=\frac{{{\mathcal{D}}_{ii}^{1/2}{\mathcal{D}}_{jj}^{1/2}}}{\sum_{k\in[n]}{\mathcal{D}}_{kk}}.$ (A.1)
 Note the all-ones vector $\bm{1}_{n}$ is in the zero-eigenspace of $H({\bm{u}}^{*})$ or $H^{*}({\bm{u}}^{*})$ (in both the left- and right-eigenspace since $H({\bm{u}}^{*})$ and $H^{*}({\bm{u}}^{*})$ are symmetric).
@@
 On the other hand, since $\nabla l({\bm{u}}^{*})-\nabla l({\widehat{\bm{u}}})=-\bar{H}({\bm{W}})({\widehat{\bm{u}}}-{\bm{u}}^{*})$,
-$-\mathcal{P}_{1}{\mathcal{D}}^{-1/2}\bar{H}({\bm{W}})({\widehat{\bm{u}}}-{\bm{u}}^{*})=\frac{1}{\sum_{i\in[n]}{\mathcal{D}}_{ii}}({\mathcal{D}}_{11}^{1/2},\ldots,{\mathcal{D}}_{nn}^{1/2})^{\top}\bm{1}^{\top}_{n}(\nabla l({\bm{u}}^{*})-\nabla l({\widehat{\bm{u}}}))$
+$)\displaystyle-\mathcal{P}_{1}{\mathcal{D}}^{-1/2}\bar{H}({\bm{W}})({\widehat{\bm{u}}}-{\bm{u}}^{*})=\frac{1}{\sum_{i\in[n]}{\mathcal{D}}_{ii}}({\mathcal{D}}_{11}^{1/2},\ldots,{\mathcal{D}}_{nn}^{1/2})^{\top}\bm{1}^{\top}_{n}(\nabla l({\bm{u}}^{*})-\nabla l({\widehat{\bm{u}}}$))
 $=$ $\ \frac{1}{\sum_{i\in[n]}{\mathcal{D}}_{ii}}\left\{\sum_{i\in[n]}\sum_{j\in\delta\{i\}}g(X_{ij},\widehat{u}_{i}-\widehat{u}_{j})-g(X_{ij},u_{i}^{*}-u_{j}^{*})\right\}({\mathcal{D}}_{11}^{1/2},\ldots,{\mathcal{D}}_{nn}^{1/2})^{\top}=0,$
-which follows as a result of the validity assumption on $f$ (that implies $g(x;y)+g(-x;-y)=0$).
+which follows as a result of the validity assumption on $f$ (that implies g(x;y)+g(x;y)=0g(x;y)+g(-x;-y)=0).
 Combining both pieces, we obtain $\mathcal{P}_{1}\bm{\zeta}=0$.
 Meanwhile, since ${\mathcal{L}}_{sym}^{\dagger}=\sum_{t=0}^{\infty}({\mathcal{A}}-\mathcal{P}_{1})^{t}-\mathcal{P}_{1}$, it follows from the Cauchy–Schwarz inequality that
-$\|{\mathcal{L}}^{\dagger}_{sym}\bm{\zeta}\|_{\infty}=$ $\left\|\sum_{t=0}^{\infty}({\mathcal{A}}-\mathcal{P}_{1})^{t}\bm{\zeta}\right\|_{\infty}$ (14)
+$\|{\mathcal{L}}^{\dagger}_{sym}\bm{\zeta}\|_{\infty}=\$$\left\|\sum_{t=0}^{\infty}({\mathcal{A}}-\mathcal{P}_{1})^{t}\bm{\zeta}\right\|_{\infty}$ (A.2)
 […]
@@
 […]
-If choosing $L=C(\kappa_{n}q_{n}/p_{n})^{2}\log n$ for some sufficiently large constant $C>0$, then the second term in (15) is $o_{p}(1)$ as a result of Lemma 14.
-For the first term in (15), it follows from the direct computation using (13) that ${\mathcal{A}}^{t}:{\mathbb{R}}^{n}\to{\mathbb{R}}^{n}$, viewed as an operator in the $\ell_{\infty}$ norm, can be bounded as
-$\|{\mathcal{A}}^{t}\|_{\ell_{\infty}\to\ell_{\infty}}$ $=\max_{i\in[n]}\sum_{j\in[n]}|\{{\mathcal{A}}^{t}\}_{ij}|$ (16)
-(13)${{\leq}}\max_{i\in[n]}\sum_{j_{1}\in[n]}\cdots\sum_{j_{t-1}\in[n]}\sum_{j\in[n]}\frac{H_{ij_{1}}\cdots H_{j_{t-1}j}}{{{\mathcal{D}}_{ii}^{1/2}}{\mathcal{D}}_{j_{1}j_{1}}\cdots{\mathcal{D}}_{j_{t-1}j_{t-1}}{{\mathcal{D}}_{jj}^{1/2}}}$
+If choosing $L=C(\kappa_{n}q_{n}/p_{n})^{2}\log n$ for some sufficiently large constant $C>0$, then the second term in (A.3) is $o_{p}(1)$ as a result of Lemma 2.
+For the first term in (A.3), it follows from the direct computation using (A.1) that ${\mathcal{A}}^{t}:{\mathbb{R}}^{n}\to{\mathbb{R}}^{n}$, viewed as an operator in the $\ell_{\infty}$ norm, can be bounded as
+$\|{\mathcal{A}}^{t}\|_{\ell_{\infty}\to\ell_{\infty}}$ $=\max_{i\in[n]}\sum_{j\in[n]}|\{{\mathcal{A}}^{t}\}_{ij}|$ (A.4)
+[…]
 $\leq\max_{i\in[n]}\frac{1}{\min_{k\in[n]}{{\mathcal{D}}_{kk}^{1/2}}}\sum_{j_{1}\in[n]}\cdots\sum_{j_{t-1}\in[n]}\frac{H_{ij_{1}}\cdots H_{j_{t-1}j}}{{{\mathcal{D}}_{ii}^{1/2}}{\mathcal{D}}_{j_{1}j_{1}}\cdots{\mathcal{D}}_{j_{t-1}j_{t-1}}}$
-$=\frac{\max_{i\in[n]}{{\mathcal{D}}_{ii}^{1/2}}}{\min_{k\in[n]}{{\mathcal{D}}_{kk}^{1/2}}}\lesssim\left(\frac{\kappa_{n}q_{n}}{p_{n}}\right)^{1/2}.$
-Consequently,
+$=\frac{\max_{i\in[n]}{{\mathcal{D}}_{ii}^{1/2}}}{\min_{k\in[n]}{{\mathcal{D}}_{kk}^{1/2}}}\lesssim\left(\frac{\kappa_{n}q_{n}}{p_{n}}\right)^{1/2},$
+where the last inequality follows from the standard degree concentration bounds (B.6). Consequently,
 $\left\|\sum_{t=0}^{L}{\mathcal{A}}^{t}\bm{\zeta}\right\|_{\infty}\lesssim L\left({\frac{\kappa_{n}q_{n}}{p_{n}}}\right)^{1/2}\|\bm{\zeta}\|_{\infty}\lesssim(\log n)\cdot\left(\frac{\kappa_{n}q_{n}}{p_{n}}\right)^{5/2}\|\bm{\zeta}\|_{\infty}=o_{p}(1),$
 where the last step follows from Lemma 3 and the assumptions in the lemma.
-To prove the last statement in Lemma 13, note that since $({\mathcal{A}}-\mathcal{P}_{1})^{t}$ is semi-positive definite, $\{({\mathcal{A}}-\mathcal{P}_{1})^{t}\}_{ii}\leq\|({\mathcal{A}}-\mathcal{P}_{1})^{t}\|_{2}$ for every $t$.
-By a similar truncation argument as above, we only need to prove for summation of $t$ from $2$ to $L$, and this can be verified by a direct computation as in (16):
-$\max_{i\in[n]}\sum_{t=2}^{L}\{({\mathcal{A}}-\mathcal{P}_{1})^{t}\}_{ii}\lesssim\frac{L}{({c_{n,3}np_{n}})^{1/2}}\lesssim\frac{c_{n,4}^{2}}{c_{n,3}^{5/2}}\left\{\frac{q_{n}^{4}(\log n)^{2}}{np_{n}^{5}}\right\}^{1/2}=o_{p}(1),$
-where the details are omitted.
-Appendix B Proofs of Lemmas 1415
-Before showing Lemmas 1415, we first cite the uniform consistency result in Han et al. (2023):
-Lemma 16
-Let ${\mathcal{G}}\sim{\mathsf{G}}(n,p_{n},q_{n})$. Under Assumptions 4-6, if
+To prove the last statement in Lemma 1, note that $\{({\mathcal{A}}-\mathcal{P}_{1})^{t}\}_{ii}\leq\|({\mathcal{A}}-\mathcal{P}_{1})^{t}\|_{2}$ for every $t$.
+By a similar truncation argument as above, we only need to prove the asymptotic negligibility of the partial sum of $t$ from $2$ to $L$:
+$\max_{i\in[n]}\sum_{t=2}^{L}\{({\mathcal{A}}-\mathcal{P}_{1})^{t}\}_{ii}\leq\max_{i\in[n]}\sum_{t=2}^{L}\{{\mathcal{A}}^{t}\}_{ii}\lesssim\frac{L\kappa_{n}^{1/2}}{(np_{n})^{1/2}}\lesssim{\kappa_{n}^{5/2}}\left\{\frac{q_{n}^{4}(\log n)^{2}}{np_{n}^{5}}\right\}^{1/2}=o_{p}(1),$
+where the first inequality follows from the fact that $({\mathcal{A}}-\mathcal{P}_{1})^{t}\preceq{\mathcal{A}}^{t}$ for $t\geq 1$, and the second inequality follows from similar computation as in (A.4), that is,
+$\{{\mathcal{A}}^{t}\}_{ii}$ ${\leq}\max_{i\in[n]}\sum_{j_{1}\in[n]}\cdots\sum_{j_{t-1}\in[n]}\frac{H_{ij_{1}}\cdots H_{j_{t-1}i}}{{{\mathcal{D}}_{ii}^{1/2}}{\mathcal{D}}_{j_{1}j_{1}}\cdots{\mathcal{D}}_{j_{t-1}j_{t-1}}{{\mathcal{D}}_{ii}^{1/2}}}$
+$\leq\max_{i\in[n]}(\max_{j_{t-1}\in[n]}\frac{H_{j_{t-1}i}}{{{\mathcal{D}}_{j_{t-1}j_{t-1}}}})\sum_{j_{1}\in[n]}\cdots\sum_{j_{t-1}\in[n]}\frac{H_{ij_{1}}\cdots H_{j_{t-2}j_{t-1}}}{{{\mathcal{D}}_{ii}}{\mathcal{D}}_{j_{1}j_{1}}\cdots{\mathcal{D}}_{j_{t-2}j_{t-2}}}$
+[…]
+Appendix B Proofs of Lemmas 2-3
+Before showing Lemmas 2-3, we first cite the uniform consistency result in han2022general:
+Lemma 4.
+Let ${\mathcal{G}}\sim{\mathsf{G}}(n,p_{n},q_{n})$. Under Assumptions 1-3, if
 $\alpha_{n}=\frac{c_{n,2}}{c_{n,3}}\left\{\frac{q_{n}^{2}(\log n)^{3}}{np_{n}^{3}}\right\}^{1/2}\to 0\qquad n\to\infty,$
@@
 B.1 Proof of Lemma 2
-Denote eigenvalues of $\mathcal{L}_{\text{sym }}=I-\mathcal{A}=I-\mathcal{D}^{-1/2}\mathcal{W}\mathcal{D}^{-1/2}$ in increasing order as $0=\lambda_{1}\leq\cdots\leq\lambda_{n}\leq 2$, where ${\mathcal{D}}$ and $\mathcal{W}$ are all random matrices (where randomness is from the comparison graph ${\mathcal{G}}$ only).
+The proof of Lemma 2 follows a similar approach as (han2023unified, Lemma 5.3).
+Denote the eigenvalues of $\mathcal{L}_{\text{sym }}=I-\mathcal{A}=I-\mathcal{D}^{-1/2}\mathcal{W}\mathcal{D}^{-1/2}$ in increasing order as $0=\lambda_{1}\leq\cdots\leq\lambda_{n}\leq 2$, where $\mathcal{D}$} and $\mathcal{W}$ are all random matrices (where randomness is from the comparison graph $\mathcal{G}$} only).
 In particular, $\mathcal{W}$ is an off-diagonal weight matrix with $\mathcal{W}_{ij}=\mathbb{I}_{\{(i,j)\in{\mathcal{E}}\}}z_{ij}$, where
 $z_{ij}$ $={\mathbb{E}}_{X_{ij}}\left[-\partial_{2}g(X_{ij},u^{*}_{i}-u^{*}_{j})\right]\in[c_{n,3},c_{n,4}],\quad i\neq j$
-due to Assumption 6, with ${\mathbb{E}}_{X_{ij}}[\cdot]$ denoting expectation for $X_{ij}$.
-It is clear $\left\|\mathcal{A}-\mathcal{P}_{1}\right\|_{2}=\max\left\{1-\lambda_{2},\lambda_{n}-1\right\}$.
-To analyze the spectrum of $\mathcal{L}_{\text{sym }}$, we consider its expected version first.
-Let $\overline{\mathcal{D}}$ and $\overline{\mathcal{W}}$ be the expectation of $\mathcal{D}$ and $\mathcal{W}$, respectively, and ¯$\bar{\mathcal{L}}_{\text{sym }}=I-\bar{\mathcal{D}}^{-1/2}\bar{\mathcal{W}}\bar{\mathcal{D}}^{-1/2}$.
-By definition, ¯$\bar{\mathcal{L}}_{\text{sym }}$ is also a normalized graph Laplacian matrix; we denote the eigenvalues of ¯$\overline{\mathcal{L}}_{\text{sym }}$ by $0=\bar{\lambda}_{1}\leq\cdots\leq\bar{\lambda}_{n}\leq 2$.
-A computation based on Cheeger’s inequalities yields a lower bound for the spectral gap of the expected Laplacian (Chung, 1997; Bauer and Jost, 2013; Han and Xu, 2023): There exists an absolute constant $c>0$ such that
-$\max\left\{1-\bar{\lambda}_{2},\bar{\lambda}_{n}-1\right\}\leq 1-c\left(\frac{p_{n}}{\kappa_{n}q_{n}}\right)^{2}.$ (17)
-We next apply a concentration argument to show that the spectra of $\mathcal{L}_{\text{sym }}$ and ¯$\bar{\mathcal{L}}_{\text{sym }}$ are close. To this end, we first apply Weyl’s inequality to obtain
-$\max\left\{|(1-\lambda_{2})-(1-\bar{\lambda}_{2})|,|(\lambda_{n}-1)-(\bar{\lambda}_{n}-1)|\right\}$
-$\leq$ $\left\|\mathcal{L}_{\text{sym }}-\overline{\mathcal{L}}_{\text{sym }}\right\|_{2}=\left\|\mathcal{D}^{-1}\mathcal{W}-\overline{\mathcal{D}}^{-1}\overline{\mathcal{W}}\right\|_{2}$
-$\leq$ $\left\|\left(\mathcal{D}^{-1}-\overline{\mathcal{D}}^{-1}\right)\overline{\mathcal{W}}\right\|_{2}+\left\|\mathcal{D}^{-1}(\mathcal{W}-\overline{\mathcal{W}})\right\|_{2}$
-$\leq$ $\left\|\mathcal{D}^{-1}-\overline{\mathcal{D}}^{-1}\right\|_{2}\|\overline{\mathcal{W}}\|_{2}+\left\|\mathcal{D}^{-1}\right\|_{2}\|\mathcal{W}-\overline{\mathcal{W}}\|_{2}.$
+due to Assumption 3, with ${\mathbb{E}}_{X_{ij}}[\cdot]$ denoting expectation for $X_{ij}$. It is clear $\left\|\mathcal{A}-\mathcal{P}_{1}\right\|_{2}=\max\left\{1-\lambda_{2},\lambda_{n}-1\right\}$.
+[…]
+$\max\left\{1-\bar{\lambda}_{2},\bar{\lambda}_{n}-1\right\}\leq 1-c\left(\frac{p_{n}}{\kappa_{n}q_{n}}\right)^{2}.$ (B.5)
+[…]
+$\max\left\{\left|(1-\lambda_{2})-(1-\bar{\lambda}_{2})\right|,\ \left|(\lambda_{n}-1)-(\bar{\lambda}_{n}-1)\right|\right\}$
+[…]
+[…]
+$+\left\|\overline{\mathcal{D}}^{-1/2}\right\|_{2}\left\|\overline{\mathcal{W}}\right\|_{2}\left\|\mathcal{D}^{-1/2}-\overline{\mathcal{D}}^{-1/2}\right\|_{2}$
 By the Chernoff bound, with probability at least $1-n^{-2}$, all $D_{ii}$ are concentrated around their means
-$\left|\mathcal{D}_{ii}-\overline{\mathcal{D}}_{ii}\right|\lesssim\left({c_{n,4}\overline{\mathcal{D}}_{ii}\log n}\right)^{1/2}\quad i\in[n],$
+$\left|\mathcal{D}_{ii}-\overline{\mathcal{D}}_{ii}\right|\lesssim\left({c_{n,4}\overline{\mathcal{D}}_{ii}\log n}\right)^{1/2}\quad i\in[n],$ (B.6)
 so that
-$\left\|\mathcal{D}^{-1}-\overline{\mathcal{D}}^{-1}\right\|_{2}$ $=\max_{i\in[n]}\left|\mathcal{D}_{ii}^{-1}-\overline{\mathcal{D}}_{ii}^{-1}\right|\lesssim\max_{i\in[n]}\left(\overline{\mathcal{D}}_{ii}^{-2}\left|\mathcal{D}_{ii}-\overline{\mathcal{D}}_{ii}\right|\right)\lesssim\left\{\frac{c_{n,4}\log n}{(c_{n,3}np_{n})^{3}}\right\}^{1/2}$
-$\left\|\mathcal{D}^{-1}\right\|_{2}$ $\lesssim\left\|\overline{\mathcal{D}}^{-1}\right\|_{2}\lesssim\frac{1}{c_{n,3}np_{n}}.$
+$\left\|\mathcal{D}^{-1/2}-\overline{\mathcal{D}}^{-1/2}\right\|_{2}$ $=\max_{i\in[n]}\left|\mathcal{D}_{ii}^{-1/2}-\overline{\mathcal{D}}_{ii}^{-1/2}\right|\lesssim\max_{i\in[n]}\left(\overline{\mathcal{D}}_{ii}^{-3/2}\left|\mathcal{D}_{ii}-\overline{\mathcal{D}}_{ii}\right|\right)\lesssim\left\{\frac{c_{n,4}\log n}{(c_{n,3}np_{n})^{2}}\right\}^{1/2}$
+$\left\|\mathcal{D}^{-1/2}\right\|_{2}$ $\lesssim\left\|\overline{\mathcal{D}}^{-1/2}\right\|_{2}\lesssim\sqrt{\frac{1}{c_{n,3}np_{n}}}.$
 On the other hand,
-$\|\overline{\mathcal{W}}\|_{2}=\left\|\overline{\mathcal{D}}\overline{\mathcal{D}}^{-1}\overline{\mathcal{W}}\right\|_{2}\leq\|\overline{\mathcal{D}}\|_{2}\left\|\overline{\mathcal{D}}^{-1}\overline{\mathcal{W}}\right\|_{2}\leq 2\|\overline{\mathcal{D}}\|_{2}=2\max_{i\in[n]}\overline{\mathcal{D}}_{ii}\lesssim c_{n,4}nq_{n}.$
-By the matrix Bernstein inequality (Tropp, 2012), with probability at least $1-n^{-2}$,
+$\|\overline{\mathcal{W}}\|_{2}=\left\|\overline{\mathcal{D}}^{1/2}(I-\bar{\mathcal{L}}_{\text{sym}})\overline{\mathcal{D}}^{1/2}\right\|_{2}\leq\|\overline{\mathcal{D}}\|_{2}\left\|I-\bar{\mathcal{L}}_{\text{sym}}\right\|_{2}\leq\|\overline{\mathcal{D}}\|_{2}=\max_{i\in[n]}\overline{\mathcal{D}}_{ii}\lesssim c_{n,4}nq_{n}.$
+By the matrix Bernstein inequality (tropp2012user), with probability at least $1-n^{-2}$,
 $\|\mathcal{W}-\overline{\mathcal{W}}\|_{2}\lesssim c_{n,4}\left({nq_{n}\log n}\right)^{1/2}.$
 In conclusion, the upper bound for $\left\|\mathcal{L}_{\text{sym}}-\overline{\mathcal{L}}_{\text{sym}}\right\|_{2}$ is
-$\left\|\mathcal{L}_{\text{sym }}-\overline{\mathcal{L}}_{\text{sym}}\right\|_{2}\lesssim\left(\frac{\kappa_{n}^{3}q_{n}^{2}\log n}{np_{n}^{3}}\right)^{1/2}+\left(\frac{\kappa_{n}^{2}q_{n}\log n}{np_{n}^{2}}\right)^{1/2}.$ (18)
-The desired result follows if the expected spectral gap in (17) dominates the concentration deviations (18):
+$\left\|\mathcal{L}_{\text{sym }}-\overline{\mathcal{L}}_{\text{sym}}\right\|_{2}\lesssim\left(\frac{\kappa_{n}^{3}q_{n}^{2}\log n}{np_{n}^{3}}\right)^{1/2}+\left(\frac{\kappa_{n}^{2}q_{n}\log n}{np_{n}^{2}}\right)^{1/2}.$ (B.7)
+The desired result follows if the expected spectral gap in (B.5) dominates the concentration deviations (B.7):
 $\frac{\left(\frac{\kappa_{n}^{3}q_{n}^{2}\log n}{np_{n}^{3}}\right)^{1/2}+\left(\frac{\kappa_{n}^{2}q_{n}\log n}{np_{n}^{2}}\right)^{1/2}}{\left(\frac{p_{n}}{\kappa_{n}q_{n}}\right)^{2}}\to 0\quad n\to\infty,$
@@
 Recall the formula for the Hessian matrix $H({\bm{u}})$ and $\bar{H}({\bm{W}})$:
-$\{H({\bm{u}})\}_{ij}$ =$=\left\{\begin{array}[]{ll}\partial_{ij}l({\bm{u}})=-\partial_{2}g(X_{ij},u_{i}-u_{j})\times\mathbb{I}_{\{(i,j)\in{\mathcal{E}}\}}&\quad i\neq j\\ -\sum_{k\in\delta\{i\}}\partial_{ik}l({\bm{u}})&\quad i=j\end{array}\right.$ (21)
-$\{\bar{H}({\bm{W}})\}_{ij}$ =$=\left\{\begin{array}[]{ll}\partial_{ij}l(\bm{w}_{i})=-\partial_{2}g(X_{ij},w_{ii}-w_{ij})\times\mathbb{I}_{\{(i,j)\in{\mathcal{E}}\}}&\quad i\neq j\\ -\sum_{k\in\delta\{i\}}\partial_{ik}l(\bm{w}_{i})&\quad i=j\end{array}\right.$ (24)
-where ${\bm{w}}_{1},\ldots,{\bm{w}}_{n}$ are $n$ points on the line segment between ${\widehat{\bm{u}}}$ and ${\bm{u}}^{*}$, and ${\bm{w}}_{i}=(w_{i1},\ldots,w_{in})\in{\mathbb{R}}^{n}$.
+$\{H({\bm{u}})\}_{ij}$ $=\left\{\begin{array}[]{ll}\partial_{ij}l({\bm{u}})=-\partial_{2}g(X_{ij},u_{i}-u_{j})\times\mathbb{I}_{\{(i,j)\in{\mathcal{E}}\}}&\quad i\neq j\\ -\sum_{k\in\delta\{i\}}\partial_{ik}l({\bm{u}})&\quad i=j\end{array}\right$.
+(B.10)
+$\{\bar{H}({\bm{W}})\}_{ij}$ $=\left\{\begin{array}[]{ll}\partial_{ij}l(\bm{w}_{i})=-\partial_{2}g(X_{ij},w_{ii}-w_{ij})\times\mathbb{I}_{\{(i,j)\in{\mathcal{E}}\}}&\quad i\neq j\\ -\sum_{k\in\delta\{i\}}\partial_{ik}l(\bm{w}_{i})&\quad i=j\end{array}\right$.
+(B.13)
+where ${\bm{w}}_{1},\ldots,{\bm{w}}_{n}$ are $n$ points on the line segment between $^{\widehat{\bm{u}}}$ and $\bm{u}$$^{*}$, and ${\bm{w}}_{i}=(w_{i1},\ldots,w_{in})\in{\mathbb{R}}^{n}$.
 Write the $i$th element of $\{\bar{H}({\bm{W}})-H\left({\bm{u}}^{*}\right)\}\left(\widehat{{\bm{u}}}-{\bm{u}}^{*}\right)$ as
-|[$\left|\left[\{\bar{H}({\bm{W}})-H\left({\bm{u}}^{*}\right)\}\left(\widehat{{\bm{u}}}-{\bm{u}}^{*}\right)\right]_{i}\right|$
+$\left|\left[\{\bar{H}({\bm{W}})-H\left({\bm{u}}^{*}\right)\}\left(\widehat{{\bm{u}}}-{\bm{u}}^{*}\right)\right]_{i}\right|$
 $=$ $\left|\sum_{j\neq i}\left\{\bar{H}({\bm{W}})-H\left({\bm{u}}^{*}\right)\right\}_{ij}(\widehat{u}_{j}-u^{*}_{j})+\left\{\bar{H}({\bm{W}})-H\left({\bm{u}}^{*}\right)\right\}_{ii}(\widehat{u}_{i}-u^{*}_{i})\right|$
@@
 By the mean value theorem, for every $i,j$ and all sufficiently large $n$, there exists $\xi_{ij}\in[-M_{n}-1,M_{n}+1]$ depending on $i,j$ such that
-|$\left|\left\{\bar{H}({\bm{W}})-H\left({\bm{u}}^{*}\right)\right\}_{ij}\right|$ =(21)|{${{=}}\left|\left\{\partial_{2}g(X_{ij},u^{*}_{i}-u^{*}_{j})-\partial_{2}g(X_{ij},w_{ii}-w_{ij})\right\}\times\mathbb{I}_{\{(i,j)\in{\mathcal{E}}\}}\right|$
+[…]
 $=\left|\partial_{22}g(X_{ij},\xi_{ij})\left\{\left(u_{i}^{*}-w_{ii}\right)-\left(u_{j}^{*}-w_{ij}\right)\right\}\times\mathbb{I}_{\{(i,j)\in{\mathcal{E}}\}}\right|$
@@
 where the penultimate inequality follows from Assumption 4. Summing over $j\in[n]$,
-|[$\left|\left[\left\{\bar{H}({\bm{W}})-H\left({\bm{u}}^{*}\right)\right\}\left(\widehat{{\bm{u}}}-{\bm{u}}^{*}\right)\right]_{i}\right|\leq 4c_{n,5}\|\widehat{{\bm{u}}}-{\bm{u}}^{*}\|^{2}_{\infty}|\delta\{i\}|.$
+$\left|\left[\left\{\bar{H}({\bm{W}})-H\left({\bm{u}}^{*}\right)\right\}\left(\widehat{{\bm{u}}}-{\bm{u}}^{*}\right)\right]_{i}\right|\leq 4c_{n,5}\|\widehat{{\bm{u}}}-{\bm{u}}^{*}\|^{2}_{\infty}|\delta\{i\}|.$
 Combining the results with Lemma 4, we can have
@@
 Next, we prove the second statement in Lemma 3.
-Similar to the previous proof, we can decompose |[$\left\lvert[\{H({\bm{u}}^{*})-H^{\ast}({\bm{u}}^{\ast})\}({\widehat{\bm{u}}}-{\bm{u}}^{*})]_{i}\right\lvert$ into two parts and obtain
-|[$\left\lvert[\{H({\bm{u}}^{*})-H^{\ast}({\bm{u}}^{\ast})\}({\widehat{\bm{u}}}-{\bm{u}}^{*})]_{i}\right\lvert$
+Similar to the previous proof, we can decompose $\left\lvert[\{H({\bm{u}}^{*})-H^{\ast}({\bm{u}}^{\ast})\}({\widehat{\bm{u}}}-{\bm{u}}^{*})]_{i}\right\lvert$ into two parts and obtain
+$\left\lvert[\{H({\bm{u}}^{*})-H^{\ast}({\bm{u}}^{\ast})\}({\widehat{\bm{u}}}-{\bm{u}}^{*})]_{i}\right\lvert$
 $\leq$ $\left|\sum_{j\neq i}\left\{H({\bm{u}}^{*})-H^{\ast}({\bm{u}}^{\ast})\right\}_{ij}(\widehat{u}_{j}-u^{*}_{j})\right|+\left|\sum_{j\neq i}\left\{H({\bm{u}}^{*})-H^{\ast}({\bm{u}}^{\ast})\right\}_{ij}(\widehat{u}_{i}-u^{*}_{i})\right|$
 $\leq$ $\underbrace{\left|\sum_{j\neq i}\left\{H({\bm{u}}^{*})-H^{\ast}({\bm{u}}^{\ast})\right\}_{ij}(\widehat{u}_{j}-u^{*}_{j})\right|}_{(A.1)}+\underbrace{\left|\sum_{j\neq i}\left\{H({\bm{u}}^{*})-H^{\ast}({\bm{u}}^{\ast})\right\}_{ij}\right|\|{\widehat{\bm{u}}}-{\bm{u}}^{*}\|_{\infty}}_{(A.2)}.$
-[…]
+To bound $(A.2)$, note $\sum_{j\neq i}\left\{H({\bm{u}}^{*})-H^{\ast}({\bm{u}}^{\ast})\right\}_{ij}$ is a sum of independent centered random variables bounded by $2c_{n,4}$.
+By Hoeffding’s inequality, with probability at least $1-n^{-2}$,
 $\left|\sum_{j\neq i}\left\{H({\bm{u}}^{*})-H^{\ast}({\bm{u}}^{\ast})\right\}_{ij}\right|\lesssim c_{n,4}\left(|\delta\{i\}|\log n\right)^{1/2}$
 Therefore,
-$(A.2)\lesssim c_{n,4}\left(|\delta\{i\}|\log n\right)^{1/2}\times\|{\widehat{\bm{u}}}-{\bm{u}}^{\ast}\|_{\infty}\lesssim\left({|\delta\{i\}|}\right)^{1/2}\times\frac{c_{n,2}c_{n,4}q_{n}(\log n)^{2}}{c_{n,3}np_{n}^{3/2}}.$ (25)
-[…]
+$(A.2)\lesssim c_{n,4}\left(|\delta\{i\}|\log n\right)^{1/2}\times\|{\widehat{\bm{u}}}-{\bm{u}}^{\ast}\|_{\infty}\lesssim\left({|\delta\{i\}|}\right)^{1/2}\times\frac{c_{n,2}c_{n,4}q_{n}(\log n)^{2}}{c_{n,3}np_{n}^{3/2}}.$ (B.14)
+Unfortunately, $(A.1)$ cannot be treated similarly as a result of dependence between
 $\left\{H({\bm{u}}^{*})-H^{\ast}({\bm{u}}^{\ast})\right\}_{ij}$ and $\widehat{u}_{j}-u^{*}_{j}$.
-To address this issue, we apply a leave-one-out approach based on a modification of Gao et al.
-(2023).
+To address this issue, we apply a leave-one-out approach based on a modification of gao2023uncertainty.
 For $i\in[n]$, let
-$l^{(-i)}({\bm{u}})=\sum_{j\neq i,k\neq i,(j,k)\in{\mathcal{E}}}\log f(X_{jk},u_{j}-u_{k}).$ (26)
+$l^{(-i)}({\bm{u}})=\sum_{j\neq i,k\neq i,(j,k)\in{\mathcal{E}}}\log f(X_{jk},u_{j}-u_{k}).$ (B.15)
 We define $\widehat{{\bm{u}}}^{(-i)}\in{\mathbb{R}}^{n-1}$ as leave-one-out estimator by maximizing the likelihood
@@
 $\|{\widehat{\bm{u}}}^{(-i)}-{\bm{u}}_{-i}^{\ast}\|_{\infty}\lesssim\frac{c_{n,2}}{c_{n,3}}\left\{\frac{q_{n}^{2}(\log n)^{3}}{np_{n}^{3}}\right\}^{1/2}.$
-[…]
-[…]
-[…]
-$(J.2)\lesssim c_{n,4}\left(|\delta\{i\}|\log n\right)^{1/2}\times\|{\widehat{\bm{u}}}^{(-i)}-{\bm{u}}_{-i}^{\ast}\|_{\infty}.$ (28)
-[…]
-[…]
-$\lesssim c_{n,4}\left({|\delta\{i\}|}\right)^{1/2}\times\|{\widehat{\bm{u}}}_{-i}-{\widehat{\bm{u}}}^{(-i)}\|_{2},$ (29)
+Now the term $(A.1)$ can be bounded using the triangle inequality as
+$(A.1)\leq$ $\underbrace{\left|\sum_{j\neq i}\left\{H({\bm{u}}^{*})-H^{\ast}({\bm{u}}^{\ast})\right\}_{ij}\left(\widehat{u}_{j}-\widehat{u}_{j}^{(i)}\right)\right|}_{(J.1)}+\underbrace{\left|\sum_{j\neq i}\left\{H({\bm{u}}^{*})-H^{\ast}({\bm{u}}^{\ast})\right\}_{ij}\left(\widehat{u}^{(i)}_{j}-u_{j}^{\ast}\right)\right|}_{(J.2)}.$ (B.16)
+Since ${\widehat{\bm{u}}}^{(-i)}$ is independent of $\{X_{ij}\}_{j\in\delta\{i\}}$, we can bound ($J.2$) similar to the steps in bounding $(A.2)$ by first conditioning on ${\widehat{\bm{u}}}^{(-i)}$.
+With probability at least $1-n^{-2}$,
+$(J.2)\lesssim c_{n,4}\left(|\delta\{i\}|\log n\right)^{1/2}\times\|{\widehat{\bm{u}}}^{(-i)}-{\bm{u}}_{-i}^{\ast}\|_{\infty}.$ (B.17)
+To bound ($J.1$), we apply the Cauchy–Schwarz inequality to obtain
+$(J.1)$ $\leq\left\|\left[\left\{H({\bm{u}}^{*})-H^{\ast}({\bm{u}}^{\ast})\right\}_{i,:}\right]_{-i}\right\|_{2}\|{\widehat{\bm{u}}}_{-i}-{\widehat{\bm{u}}}^{(-i)}\|_{2}$
+$\lesssim c_{n,4}\left({|\delta\{i\}|}\right)^{1/2}\times\|{\widehat{\bm{u}}}_{-i}-{\widehat{\bm{u}}}^{(-i)}\|_{2},$ (B.18)
 where $\left\{H({\bm{u}}^{*})-H^{\ast}({\bm{u}}^{\ast})\right\}_{i,:}$ denotes the $i$th row of $H({\bm{u}}^{*})-H^{\ast}({\bm{u}}^{\ast})$.
@@
 where $\widetilde{{\bm{u}}}$ lies on the line segment between $\widehat{{\bm{u}}}^{(-i)}$ and $\widehat{{\bm{u}}}_{-i}$. Applying the Cauchy–Schwarz inequality,
-$\|\widehat{{\bm{u}}}^{(-i)}-\widehat{{\bm{u}}}_{-i}\|_{2}\leq\frac{\lVert\nabla l^{(-i)}(\widehat{{\bm{u}}}_{-i})\lVert_{2}}{\lambda_{2}\{-\nabla^{2}l^{(-i)}(\widetilde{{\bm{u}}})\}},$ (30)
+$\|\widehat{{\bm{u}}}^{(-i)}-\widehat{{\bm{u}}}_{-i}\|_{2}\leq\frac{\lVert\nabla l^{(-i)}(\widehat{{\bm{u}}}_{-i})\lVert_{2}}{\lambda_{2}\{-\nabla^{2}l^{(-i)}(\widetilde{{\bm{u}}})\}}$, (B.19)
 where $\lambda_{2}\{\cdot\}$ (with some abuse of notation) denotes the second smallest eigenvalue.
@@
 Under Assumption 2, $\{g(X_{ij},u^{*}_{i}-u^{*}_{j})\}_{j:(i,j)\in{\mathcal{E}}}$ are subgaussian random variables with subgaussian norms uniformly bounded by $c_{n,2}$.
-According to (Vershynin, 2018, Lemma 2.7.6), $\{g(X_{ij},u^{*}_{i}-u^{*}_{j})^{2}\}_{j:(i,j)\in{\mathcal{E}}}$ are sub-exponential random variables with sub-exponential norms uniformly bounded by $c_{n,2}^{2}$.
-It follows from (Vershynin, 2018, Proposition 2.7.1, Exercise 2.7.10) that
+According to (vershynin2018high, Lemma 2.7.6), $\{g(X_{ij},u^{*}_{i}-u^{*}_{j})^{2}\}_{j:(i,j)\in{\mathcal{E}}}$ are sub-exponential random variables with sub-exponential norms uniformly bounded by $c_{n,2}^{2}$.
+It follows from (vershynin2018high, Proposition 2.7.1, Exercise 2.7.10) that
 $\max_{j:(i,j)\in{\mathcal{E}}}{\mathbb{E}}[g(X_{ij},u^{*}_{i}-u^{*}_{j})^{2}]\lesssim c_{n,2}^{2}$
@@
 holds with probability at least $1-n^{-3}$. Consequently, with probability at least $1-n^{-3}$,
-$\|\nabla l^{(-i)}(\widehat{{\bm{u}}}_{-i})\|_{2}$$\lesssim\left\{(c_{n,4}\|{\widehat{\bm{u}}}-{\bm{u}}^{\ast}\|_{\infty}^{2}+c_{n,2}^{2})nq_{n}\right\}^{1/2}\lesssim c_{n,2}(nq_{n})^{1/2},$ (31)
+$\|\nabla l^{(-i)}(\widehat{{\bm{u}}}_{-i})\|_{2}$ $\lesssim\left\{(c_{n,4}\|{\widehat{\bm{u}}}-{\bm{u}}^{\ast}\|_{\infty}^{2}+c_{n,2}^{2})nq_{n}\right\}^{1/2}\lesssim c_{n,2}(nq_{n})^{1/2},$ (B.20)
 where the last step follows from the observation
@@
 On the other hand, by a similar calculation as in the proof of Lemma 2 (for normalized Laplacian), with probability at least $1-n^{-2}$,
-$\lambda_{2}\{\nabla^{2}l^{(-i)}(\widetilde{{\bm{u}}})\}\gtrsim c_{n,3}np_{n}\left(\frac{p_{n}}{\kappa_{n}q_{n}}\right)^{2},$ (32)
-where the extra term $c_{n,3}np_{n}$ stems from a lower bound for the singular value of the normalization matrix. Combining (30)–(32), we conclude
-$\|\widehat{{\bm{u}}}^{(-i)}-\widehat{{\bm{u}}}_{-i}\|_{2}\lesssim\frac{c_{n,2}c_{n,4}^{2}q_{n}^{5/2}}{c_{n,3}^{3}p_{n}^{3}n^{1/2}}.$ (33)
-Putting (27), (28), (29), and (33) together,
-[…]
-$\lesssim\left({|\delta\{i\}|}\right)^{1/2}\times\frac{c_{n,2}c_{n,4}^{3}q_{n}^{5/2}(\log n)^{2}}{c_{n,3}^{3}p_{n}^{3}n^{1/2}}.$ (34)
-Consequently, combining (25) and (34), we have
+$\lambda_{2}\{\nabla^{2}l^{(-i)}(\widetilde{{\bm{u}}})\}\gtrsim c_{n,3}np_{n}\left(\frac{p_{n}}{\kappa_{n}q_{n}}\right)^{2},$ (B.21)
+where the extra term $c_{n,3}np_{n}$ stems from a lower bound for the singular value of the normalization matrix. Combining (B.19)–(B.21), we conclude
+$\|\widehat{{\bm{u}}}^{(-i)}-\widehat{{\bm{u}}}_{-i}\|_{2}\lesssim\frac{c_{n,2}c_{n,4}^{2}q_{n}^{5/2}}{c_{n,3}^{3}p_{n}^{3}n^{1/2}}.$ (B.22)
+Putting (B.16), (B.17), (B.18), and (B.22) together,
+$(A.1)$ $\lesssim c_{n,4}\left(|\delta\{i\}|\log n\right)^{1/2}\left(\|{\widehat{\bm{u}}}^{(-i)}-{\bm{u}}_{-i}^{\ast}\|_{\infty}+\|\widehat{{\bm{u}}}^{(-i)}-\widehat{{\bm{u}}}_{-i}\|_{2}\right)$
+$\lesssim\left({|\delta\{i\}|}\right)^{1/2}\times\frac{c_{n,2}c_{n,4}^{3}q_{n}^{5/2}(\log n)^{2}}{c_{n,3}^{3}p_{n}^{3}n^{1/2}}.$ (B.23)
+Consequently, combining (B.14) and (B.23), we have
 $|D^{-1/2}\{H({\bm{u}}^{*})-H^{\ast}({\bm{u}}^{\ast})\}({\widehat{\bm{u}}}-{\bm{u}}^{*})|_{i}$ $=D_{ii}^{-1/2}\left\lvert[\{H({\bm{u}}^{*})-H^{\ast}({\bm{u}}^{\ast})\}({\widehat{\bm{u}}}-{\bm{u}}^{*})]_{i}\right\lvert$
-[…]
+$\leq D_{ii}^{-1/2}[(A.1)+(A.2)]$
 $\lesssim\frac{c_{n,2}c_{n,4}^{3}q_{n}^{5/2}(\log n)^{2}}{c_{n,3}^{7/2}p_{n}^{3}n^{1/2}}.$
 The second part of Lemma 3 is proved.
-Appendix C Proofs of Corollary 10 and Theorems 11-12
+Appendix C Proofs of Corollary 1 and Theorems 2-3
 C.1 Proof of Corollary 1
-Proof  If $M^{*}<\infty$, then $\sup_{n}c_{n,1}<1$ so that Assumption 4 holds under condition $(\log n)^{8}/(np_{n})\to 0$.
+Proof.
+If $M^{*}<\infty$, then $\sup_{n}c_{n,1}<1$ so that Assumption 1 holds under condition $(\log n)^{8}/(np_{n})\to 0$.
 Since $\mathbb{A}$} is finite and $g(x;y)$ is continuous in $y$, $g(x;y)$ is uniformly bounded over ${\mathbb{A}}\times[-M^{*},M^{*}]$, implying $\sup_{n}c_{n,2}<\infty$.
@@
 A similar argument shows $\sup_{n}c_{n,5}<\infty$.
-The proof is finished by appealing to (3).
+The proof is finished by appealing to (3.3).
+$\Box$
 C.2 Proof of Theorem 2
-According to (C.2) and (12), we have
+According to (6.9) and (6.12), we have
 ${\mathcal{D}}^{1/2}({\widehat{\bm{u}}}-{\bm{u}}^{*})$ $=-{\mathcal{D}}^{-1/2}\nabla l({\bm{u}}^{*})-({\mathcal{L}}^{\dagger}_{sym}-I){\mathcal{D}}^{-1/2}\nabla l({\bm{u}}^{*})$
@@
 Then, for any $k\in[n]$, we have
-$\{{\mathcal{D}}^{1/2}({\widehat{\bm{u}}}-{\bm{u}}^{*})\}_{k}$ =$=-\{{\mathcal{D}}^{-1/2}\nabla l({\bm{u}}^{*})\}_{k}+o_{p}(1),$
+$\{{\mathcal{D}}^{1/2}({\widehat{\bm{u}}}-{\bm{u}}^{*})\}_{k}$ $=-\{{\mathcal{D}}^{-1/2}\nabla l({\bm{u}}^{*})\}_{k}+o_{p}(1),$
 via Lemma 1 and Chebyshev’s inequality. Since $\mathcal{D}$} is diagonal matrix,
@@
 C.3 Proof of Theorem 3
-According to Slutsky’s Theorem and Thereom 8, it suffices to show that for any fixed $i\in[n]$ that
+According to Slutsky’s theorem and Theorem 1, it suffices to show that for any fixed $i\in[n]$ that
 $\frac{\{\rho_{i}({\bm{u}}^{*})\}^{-1}-\{\rho_{i}({\widehat{\bm{u}}})\}^{-1}}{\{\rho_{i}({\bm{u}}^{*})\}^{-1}}=o_{p}(1).$
@@
 $\left|\frac{g_{ij}^{*}(x)-\widehat{g}_{ij}(x)}{g_{ij}^{*}(x)}\right|\lesssim\frac{c_{n,2}c_{n,4}}{c_{n,3}}\left\{\frac{q_{n}^{2}(\log n)^{3}}{np_{n}^{3}}\right\}^{1/2}=o(1),$
-where the last step follows assuming (3) holds. Meanwhile,
+where the last step follows assuming (3.3) holds. Meanwhile,
 $\left|\frac{g_{ij}^{*}(x)+\widehat{g}_{ij}(x)}{g_{ij}^{*}(x)}\right|=\left|2+\frac{\widehat{g}_{ij}(x)-g_{ij}^{*}(x)}{g_{ij}^{*}(x)}\right|=O_{p}(1).$
@@
 We finish the proof of Theorem 3.
+Funding
+R. Han was supported by the Hong Kong Research Grants Council (No. 14301821) and the Hong Kong Polytechnic University (P0044617, P0045351).
+Relevant code
+The code to reproduce the main results in this paper is available at https://github.com/RJ-HAN-STAT/Ranking_Code.